libbfd

The Binary File Descriptor Library

First Edition—BFD version < 3.0 April 1991

Steve Chamberlain Cygnus Support

Cygnus Support sac@cygnus.com

BFD, Revision: 1.1.1.1 TEXinfo 1.1

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1991 Free Software Foundation, Inc. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, subject to the terms of the GNU General Public License, which includes the provision that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions.

1 Introduction

BFD is a package which allows applications to use the same routines to operate on object files whatever the object file format. A new object file format can be supported simply by creating a new BFD back end and adding it to the library. BFD is split into two parts: the front end, and the back ends (one for each object file format).

• The front end of BFD provides the interface to the user. It manages memory and various canonical data structures. The front end also decides which back end to use and when to call back end routines.

• The back ends provide BFD its view of the real world. Each back end provides a set of calls which the BFD front end can use to maintain its canonical form. The back ends also may keep around information for their own use, for greater efficiency.

1.1 History

One spur behind BFD was the desire, on the part of the GNU 960 team at Intel Oregon, for interoperability of applications on their COFF and b.out file formats. Cygnus was providing GNU support for the team, and was contracted to provide the required functionality. The name came from a conversation David Wallace was having with Richard Stallman about the library: RMS said that it would be quite hard—David said “BFD”. Stallman was right, but the name stuck. At the same time, Ready Systems wanted much the same thing, but for different object file formats: IEEE-695, Oasys, Srecords, a.out and 68k coff. BFD was first implemented by members of Cygnus Support; Steve Chamberlain (sac@cygnus.com), John Gilmore (gnu@cygnus.com), K. Richard Pixley (rich@cygnus.com) and David Henkel-

Wallace (gumby@cygnus.com).

1.2 How To Use BFD

To use the library, include ‘bfd.h’ and link with ‘libbfd.a’. BFD provides a common interface to the parts of an object file for a calling application. When an application sucessfully opens a target file (object, archive, or whatever), a pointer to an internal structure is returned. This pointer points to a structure called bfd, described in ‘bfd.h’. Our convention is to call this pointer a BFD, and instances of it within code abfd. All operations on the target object file are applied as methods to the BFD. The mapping is defined within bfd.h in a set of macros, all beginning with ‘bfd_’ to reduce namespace pollution. For example, this sequence does what you would probably expect: return the number of sections in an object file attached to a BFD abfd.

     #include "bfd.h"
     unsigned int number_of_sections(abfd)
     bfd *abfd;
     {

       return bfd_count_sections(abfd);
     }

The abstraction used within BFD is that an object file has:

• a header,

• a number of sections containing raw data (see Section 2.6 [Sections], page 15),

• a set of relocations (see Section 2.10 [Relocations], page 37), and

• some symbol information (see Section 2.7 [Symbols], page 27).

Also, BFDs opened for archives have the additional attribute of an index and contain subordinate BFDs. This approach is fine for a.out and coff, but loses efficiency when applied to formats such as S-records and IEEE-695.

1.3 What BFD Version 2 Can Do

When an object file is opened, BFD subroutines automatically determine the format of the input object file. They then build a descriptor in memory with pointers to routines that will be used to access elements of the object file’s data structures. As different information from the the object files is required, BFD reads from different sections of the file and processes them. For example, a very common operation for the linker is processing symbol tables. Each BFD back end provides a routine for converting between the object file’s representation of symbols and an internal canonical format. When the linker asks for the symbol table of an object file, it calls through a memory pointer to the routine from the relevant BFD back end which reads and converts the table into a canonical form. The linker then operates upon the canonical form. When the link is finished and the linker writes the output file’s symbol table, another BFD back end routine is called to take the newly created symbol table and convert it into the chosen output format.

1.3.1 Information Loss

Information can be lost during output. The output formats supported by BFD do not provide identical facilities, and information which can be described in one form has nowhere to go in another format. One example of this is alignment information in b.out. There is nowhere in an a.out format file to store alignment information on the contained data, so when a file is linked from b.out and an a.out image is produced, alignment information will not propagate to the output file. (The linker will still use the alignment information internally, so the link is performed correctly). Another example is COFF section names. COFF files may contain an unlimited number of sections, each one with a textual section name. If the target of the link is a format which does not have many sections (e.g., a.out) or has sections without names (e.g., the Oasys format), the link cannot be done simply. You can circumvent this problem by describing the desired input-to-output section mapping with the linker command language. Information can be lost during canonicalization. The BFD internal canonical form of the external formats is not exhaustive; there are structures in input formats for which there is no direct representation internally. This means that the BFD back ends cannot maintain all possible data richness through the transformation between external to internal and back to external formats.

This limitation is only a problem when an application reads one format and writes another. Each BFD back end is responsible for maintaining as much data as possible, and the internal BFD canonical form has structures which are opaque to the BFD core, and exported only to the back ends. When a file is read in one format, the canonical form is generated for BFD and the application. At the same time, the back end saves away any information which may otherwise be lost. If the data is then written back in the same format, the back end routine will be able to use the canonical form provided by the BFD core as well as the information it prepared earlier. Since there is a great deal of commonality between back ends, there is no information lost when linking or copying big endian COFF to little endian COFF, or a.out to b.out. When a mixture of formats is linked, the information is only lost from the files whose format differs from the destination.

1.3.2 The BFD canonical object-file format

The greatest potential for loss of information occurs when there is the least overlap between the information provided by the source format, that stored by the canonical format, and that needed by the destination format. A brief description of the canonical form may help you understand which kinds of data you can count on preserving across conversions.

files Information stored on a per-file basis includes target machine architecture, particular implementation format type, a demand pageable bit, and a write protected bit. Information like Unix magic numbers is not stored here—only the magic numbers’ meaning, so a ZMAGIC file would have both the demand pageable bit and the write protected text bit set. The byte order of the target is stored on a per-file basis, so that big- and little-endian object files may be used with one another.

sections Each section in the input file contains the name of the section, the section’s original address in the object file, size and alignment information, various flags, and pointers into other BFD data structures.

symbols Each symbol contains a pointer to the information for the object file which originally defined it, its name, its value, and various flag bits. When a BFD back end reads in a symbol table, it relocates all symbols to make them relative to the base of the section where they were defined. Doing this ensures that each symbol points to its containing section. Each symbol also has a varying amount of hidden private data for the BFD back end. Since the symbol points to the original file, the private data format for that symbol is accessible. ld can operate on a collection of symbols of wildly different formats without problems. Normal global and simple local symbols are maintained on output, so an output file (no matter its format) will retain symbols pointing to functions and to global, static, and common variables. Some symbol information is not worth retaining; in a.out, type information is stored in the symbol table as long symbol names. This information would be useless to most COFF debuggers; the linker has command line switches to allow users to throw it away. There is one word of type information within the symbol, so if the format supports symbol type information within symbols (for example, COFF, IEEE, Oasys) and the type is simple enough to fit within one word (nearly everything but aggregates), the information will be preserved. relocation level

Each canonical BFD relocation record contains a pointer to the symbol to relocate to, the offset of the data to relocate, the section the data is in, and a pointer to a relocation type descriptor. Relocation is performed by passing messages through the relocation type descriptor and the symbol pointer. Therefore, relocations can be performed on output data using a relocation method that is only available in one of the input formats. For instance, Oasys provides a byte relocation format. A relocation record requesting this relocation type would point indirectly to a routine to perform this, so the relocation may be performed on a byte being written to a 68k COFF file, even though 68k COFF has no such relocation type.

line numbers Object formats can contain, for debugging purposes, some form of mapping between symbols, source line numbers, and addresses in the output file. These addresses have to be relocated along with the symbol information. Each symbol with an associated list of line number records points to the first record of the list. The head of a line number list consists of a pointer to the symbol, which allows finding out the address of the function whose line number is being described. The rest of the list is made up of pairs: offsets into the section and line numbers. Any format which can simply derive this information can pass it successfully between formats (COFF, IEEE and Oasys).

2 BFD front end

2.1 typedef bfd

A BFD has type bfd; objects of this type are the cornerstone of any application using BFD. Using BFD consists of making references though the BFD and to data in the BFD. Here is the structure that defines the type bfd. It contains the major data about the file and pointers to the rest of the data.

struct _bfd
{
/* The filename the application opened the BFD with.  */
CONST char *filename;
/* A pointer to the target jump table.             */
const struct bfd_target *xvec;
/* To avoid dragging too many header files into every file that
   includes `bfd.h', IOSTREAM has been declared as a "char
   *", and MTIME as a "long".  Their correct types, to which they
   are cast when used, are "FILE *" and "time_t".    The iostream
   is the result of an fopen on the filename.  However, if the
   BFD_IN_MEMORY flag is set, then iostream is actually a pointer
   to a bfd_in_memory struct.  */
PTR iostream;
/* Is the file descriptor being cached?  That is, can it be closed as
   needed, and re-opened when accessed later?  */
boolean cacheable;
/* Marks whether there was a default target specified when the
   BFD was opened. This is used to select which matching algorithm
   to use to choose the back end. */
boolean target_defaulted;
/* The caching routines use these to maintain a
   least-recently-used list of BFDs */
struct _bfd *lru_prev, *lru_next;
/* When a file is closed by the caching routines, BFD retains
   state information on the file here: */
file_ptr where;
/* and here: (``once'' means at least once) */

boolean opened_once;
/* Set if we have a locally maintained mtime value, rather than
   getting it from the file each time: */
boolean mtime_set;
/* File modified time, if mtime_set is true: */
long mtime;
/* Reserved for an unimplemented file locking extension.*/
int ifd;
/* The format which belongs to the BFD. (object, core, etc.) */
bfd_format format;
/* The direction the BFD was opened with*/
enum bfd_direction {no_direction = 0,
                    read_direction = 1,
                    write_direction = 2,
                    both_direction = 3} direction;
/* Format_specific flags*/
flagword flags;
/* Currently my_archive is tested before adding origin to
   anything. I believe that this can become always an add of
   origin, with origin set to 0 for non archive files.   */
file_ptr origin;
/* Remember when output has begun, to stop strange things
   from happening. */
boolean output_has_begun;
/* Pointer to linked list of sections*/
struct sec  *sections;
/* The number of sections */
unsigned int section_count;
/* Stuff only useful for object files:
   The start address. */

bfd_vma start_address;
/* Used for input and output*/
unsigned int symcount;
/* Symbol table for output BFD (with symcount entries) */
struct symbol_cache_entry  **outsymbols;
/* Pointer to structure which contains architecture information*/
const struct bfd_arch_info *arch_info;
/* Stuff only useful for archives:*/
PTR arelt_data;
struct _bfd *my_archive;     /* The containing archive BFD.  */
struct _bfd *next;           /* The next BFD in the archive.  */
struct _bfd *archive_head;   /* The first BFD in the archive.  */
boolean has_armap;
/* A chain of BFD structures involved in a link.  */
struct _bfd *link_next;
/* A field used by _bfd_generic_link_add_archive_symbols.  This will
 be used only for archive elements.  */
int archive_pass;
/* Used by the back end to hold private data. */
union
{
struct aout_data_struct *aout_data;
struct artdata *aout_ar_data;
struct _oasys_data *oasys_obj_data;
struct _oasys_ar_data *oasys_ar_data;
struct coff_tdata *coff_obj_data;
struct pe_tdata *pe_obj_data;
struct xcoff_tdata *xcoff_obj_data;
struct ecoff_tdata *ecoff_obj_data;
struct ieee_data_struct *ieee_data;
struct ieee_ar_data_struct *ieee_ar_data;
struct srec_data_struct *srec_data;
struct ihex_data_struct *ihex_data;
struct tekhex_data_struct *tekhex_data;
struct elf_obj_tdata *elf_obj_data;
struct nlm_obj_tdata *nlm_obj_data;
struct bout_data_struct *bout_data;
struct sun_core_struct *sun_core_data;
struct sco5_core_struct *sco5_core_data;
struct trad_core_struct *trad_core_data;
struct som_data_struct *som_data;

    struct hpux_core_struct *hpux_core_data;
    struct hppabsd_core_struct *hppabsd_core_data;
    struct sgi_core_struct *sgi_core_data;
    struct lynx_core_struct *lynx_core_data;
    struct osf_core_struct *osf_core_data;
    struct cisco_core_struct *cisco_core_data;
    struct versados_data_struct *versados_data;
    struct netbsd_core_struct *netbsd_core_data;
    PTR any;
    } tdata;
  /* Used by the application to hold private data*/
  PTR usrdata;
/* Where all the allocated stuff under this BFD goes.  This is a
   struct objalloc *, but we use PTR to avoid requiring the inclusion of
   objalloc.h.  */
  PTR memory;
};

2.2 Error reporting

Most BFD functions return nonzero on success (check their individual documentation for precise semantics). On an error, they call bfd_set_error to set an error condition that callers can check by calling bfd_get_error. If that returns bfd_error_system_call, then check errno. The easiest way to report a BFD error to the user is to use bfd_perror.

2.2.1 Type bfd_error_type

The values returned by bfd_get_error are defined by the enumerated type bfd_error_

type.
typedef enum bfd_error
{
bfd_error_no_error = 0,
bfd_error_system_call,
bfd_error_invalid_target,
bfd_error_wrong_format,
bfd_error_invalid_operation,
bfd_error_no_memory,
bfd_error_no_symbols,
bfd_error_no_armap,
bfd_error_no_more_archived_files,
bfd_error_malformed_archive,
bfd_error_file_not_recognized,
bfd_error_file_ambiguously_recognized,
bfd_error_no_contents,

       bfd_error_nonrepresentable_section,
       bfd_error_no_debug_section,
       bfd_error_bad_value,
       bfd_error_file_truncated,
       bfd_error_file_too_big,
       bfd_error_invalid_error_code
     } bfd_error_type;

2.2.1.1 bfd_get_error

Synopsis

     bfd_error_type bfd_get_error (void);

Description Return the current BFD error condition.

2.2.1.2 bfd_set_error

Synopsis

     void bfd_set_error (bfd_error_type error_tag);

Description Set the BFD error condition to be error tag.

2.2.1.3 bfd_errmsg

Synopsis

     CONST char *bfd_errmsg (bfd_error_type error_tag);

Description Return a string describing the error error tag, or the system error if error tag is bfd_error_

system_call.
2.2.1.4 bfd_perror

Synopsis

     void bfd_perror (CONST char *message);

Description Print to the standard error stream a string describing the last BFD error that occurred, or the last system error if the last BFD error was a system call failure. If message is non-NULL and non-empty, the error string printed is preceded by message, a colon, and a space. It is followed by a newline.

2.2.2 BFD error handler

Some BFD functions want to print messages describing the problem. They call a BFD error handler function. This function may be overriden by the program. The BFD error handler acts like printf.

     typedef void (*bfd_error_handler_type) PARAMS ((const char *, ...));

2.2.2.1 bfd_set_error_handler

Synopsis

     bfd_error_handler_type bfd_set_error_handler (bfd_error_handler_type);

Description Set the BFD error handler function. Returns the previous function.

2.2.2.2 bfd_set_error_program_name

Synopsis

     void bfd_set_error_program_name (const char *);

Description Set the program name to use when printing a BFD error. This is printed before the error message followed by a colon and space. The string must not be changed after it is passed to this function.

2.2.2.3 bfd_get_error_handler

Synopsis

     bfd_error_handler_type bfd_get_error_handler (void);

Description Return the BFD error handler function.

2.3 Symbols

2.3.0.1 bfd_get_reloc_upper_bound

Synopsis

     long bfd_get_reloc_upper_bound(bfd *abfd, asection *sect);

Description Return the number of bytes required to store the relocation information associated with section sect attached to bfd abfd. If an error occurs, return -1.

2.3.0.2 bfd_canonicalize_reloc

Synopsis

     long bfd_canonicalize_reloc
        (bfd *abfd,
         asection *sec,
         arelent **loc,
         asymbol **syms);

Description Call the back end associated with the open BFD abfd and translate the external form of the relocation information attached to sec into the internal canonical form. Place the table into memory at loc, which has been preallocated, usually by a call to bfd_get_reloc_upper_ bound. Returns the number of relocs, or -1 on error. The syms table is also needed for horrible internal magic reasons.

2.3.0.3 bfd_set_reloc

Synopsis

     void bfd_set_reloc
        (bfd *abfd, asection *sec, arelent **rel, unsigned int count)

Description Set the relocation pointer and count within section sec to the values rel and count. The argument abfd is ignored.

2.3.0.4 bfd_set_file_flags

Synopsis

     boolean bfd_set_file_flags(bfd *abfd, flagword flags);

Description Set the flag word in the BFD abfd to the value flags. Possible errors are:

• bfd_error_wrong_format - The target bfd was not of object format.

• bfd_error_invalid_operation - The target bfd was open for reading.

• bfd_error_invalid_operation - The flag word contained a bit which was not applicable to the type of file. E.g., an attempt was made to set the D_PAGED bit on a BFD format which does not support demand paging.

2.3.0.5 bfd_set_start_address

Synopsis

     boolean bfd_set_start_address(bfd *abfd, bfd_vma vma);

Description Make vma the entry point of output BFD abfd. Returns Returns true on success, false otherwise.

2.3.0.6 bfd_get_mtime

Synopsis

     long bfd_get_mtime(bfd *abfd);

Description Return the file modification time (as read from the file system, or from the archive header for archive members).

2.3.0.7 bfd_get_size

Synopsis

     long bfd_get_size(bfd *abfd);

Description Return the file size (as read from file system) for the file associated with BFD abfd. The initial motivation for, and use of, this routine is not so we can get the exact size of the object the BFD applies to, since that might not be generally possible (archive members for example). It would be ideal if someone could eventually modify it so that such results were guaranteed. Instead, we want to ask questions like "is this NNN byte sized object I’m about to try read from file offset YYY reasonable?" As as example of where we might do this, some object formats use string tables for which the first sizeof(long) bytes of the table contain the size of the table itself, including the size bytes. If an application tries to read what it thinks is one of these string tables, without some way to validate the size, and for some reason the size is wrong (byte swapping error, wrong location for the string table, etc.), the only clue is likely to be a read error when it tries to read the table, or a "virtual memory exhausted" error when it tries to allocate 15 bazillon bytes of space for the 15 bazillon byte table it is about to read. This function at least allows us to answer the quesion, "is the size reasonable?".

2.3.0.8 bfd_get_gp_size

Synopsis

     int bfd_get_gp_size(bfd *abfd);

Description Return the maximum size of objects to be optimized using the GP register under MIPS ECOFF. This is typically set by the -G argument to the compiler, assembler or linker.

2.3.0.9 bfd_set_gp_size

Synopsis

     void bfd_set_gp_size(bfd *abfd, int i);

Description Set the maximum size of objects to be optimized using the GP register under ECOFF or MIPS ELF. This is typically set by the -G argument to the compiler, assembler or linker.

2.3.0.10 bfd_scan_vma

Synopsis

     bfd_vma bfd_scan_vma(CONST char *string, CONST char **end, int base);

Description Convert, like strtoul, a numerical expression string into a bfd_vma integer, and return that integer. (Though without as many bells and whistles as strtoul.) The expression is assumed to be unsigned (i.e., positive). If given a base, it is used as the base for conversion. A base of 0 causes the function to interpret the string in hex if a leading "0x" or "0X" is found, otherwise in octal if a leading zero is found, otherwise in decimal. Overflow is not detected.

2.3.0.11 bfd_copy_private_bfd_data

Synopsis

     boolean bfd_copy_private_bfd_data(bfd *ibfd, bfd *obfd);

Description Copy private BFD information from the BFD ibfd to the the BFD obfd. Return true on success, false on error. Possible error returns are:

• bfd_error_no_memory - Not enough memory exists to create private data for obfd.

     #define bfd_copy_private_bfd_data(ibfd, obfd) \
          BFD_SEND (obfd, _bfd_copy_private_bfd_data, \
                    (ibfd, obfd))
2.3.0.12 bfd_merge_private_bfd_data

Synopsis

     boolean bfd_merge_private_bfd_data(bfd *ibfd, bfd *obfd);

Description Merge private BFD information from the BFD ibfd to the the output file BFD obfd when linking. Return true on success, false on error. Possible error returns are:

• bfd_error_no_memory - Not enough memory exists to create private data for obfd.

     #define bfd_merge_private_bfd_data(ibfd, obfd) \
          BFD_SEND (obfd, _bfd_merge_private_bfd_data, \
                    (ibfd, obfd))
2.3.0.13 bfd_set_private_flags

Synopsis

     boolean bfd_set_private_flags(bfd *abfd, flagword flags);

Description Set private BFD flag information in the BFD abfd. Return true on success, false on error. Possible error returns are:

• bfd_error_no_memory - Not enough memory exists to create private data for obfd.

     #define bfd_set_private_flags(abfd, flags) \
          BFD_SEND (abfd, _bfd_set_private_flags, \
                    (abfd, flags))

2.3.0.14 stuff

Description Stuff which should be documented:

     #define bfd_sizeof_headers(abfd, reloc) \
          BFD_SEND (abfd, _bfd_sizeof_headers, (abfd, reloc))
     #define bfd_find_nearest_line(abfd, sec, syms, off, file, func, line) \
          BFD_SEND (abfd, _bfd_find_nearest_line,  (abfd, sec, syms, off, file, func, line))
             /* Do these three do anything useful at all, for any back end?  */
     #define bfd_debug_info_start(abfd) \
             BFD_SEND (abfd, _bfd_debug_info_start, (abfd))
     #define bfd_debug_info_end(abfd) \
             BFD_SEND (abfd, _bfd_debug_info_end, (abfd))
     #define bfd_debug_info_accumulate(abfd, section) \
             BFD_SEND (abfd, _bfd_debug_info_accumulate, (abfd, section))

#define bfd_stat_arch_elt(abfd, stat) \
        BFD_SEND (abfd, _bfd_stat_arch_elt,(abfd, stat))
#define bfd_update_armap_timestamp(abfd) \
        BFD_SEND (abfd, _bfd_update_armap_timestamp, (abfd))
#define bfd_set_arch_mach(abfd, arch, mach)\
        BFD_SEND ( abfd, _bfd_set_arch_mach, (abfd, arch, mach))
#define bfd_relax_section(abfd, section, link_info, again) \
       BFD_SEND (abfd, _bfd_relax_section, (abfd, section, link_info, again))
#define bfd_gc_sections(abfd, link_info) \
       BFD_SEND (abfd, _bfd_gc_sections, (abfd, link_info))
#define bfd_link_hash_table_create(abfd) \
       BFD_SEND (abfd, _bfd_link_hash_table_create, (abfd))
#define bfd_link_add_symbols(abfd, info) \
       BFD_SEND (abfd, _bfd_link_add_symbols, (abfd, info))
#define bfd_final_link(abfd, info) \
       BFD_SEND (abfd, _bfd_final_link, (abfd, info))
#define bfd_free_cached_info(abfd) \
       BFD_SEND (abfd, _bfd_free_cached_info, (abfd))
#define bfd_get_dynamic_symtab_upper_bound(abfd) \
       BFD_SEND (abfd, _bfd_get_dynamic_symtab_upper_bound, (abfd))
#define bfd_print_private_bfd_data(abfd, file)\
       BFD_SEND (abfd, _bfd_print_private_bfd_data, (abfd, file))
#define bfd_canonicalize_dynamic_symtab(abfd, asymbols) \
       BFD_SEND (abfd, _bfd_canonicalize_dynamic_symtab, (abfd, asymbols))
#define bfd_get_dynamic_reloc_upper_bound(abfd) \
       BFD_SEND (abfd, _bfd_get_dynamic_reloc_upper_bound, (abfd))
#define bfd_canonicalize_dynamic_reloc(abfd, arels, asyms) \
       BFD_SEND (abfd, _bfd_canonicalize_dynamic_reloc, (abfd, arels, asyms))
extern bfd_byte *bfd_get_relocated_section_contents
       PARAMS ((bfd *, struct bfd_link_info *,
                 struct bfd_link_order *, bfd_byte *,
                 boolean, asymbol **));

2.4 Memory usage

BFD keeps all of its internal structures in obstacks. There is one obstack per open BFD file, into which the current state is stored. When a BFD is closed, the obstack is deleted, and so everything which has been allocated by BFD for the closing file is thrown away. BFD does not free anything created by an application, but pointers into bfd structures become invalid on a bfd_close; for example, after a bfd_close the vector passed to bfd_ canonicalize_symtab is still around, since it has been allocated by the application, but the data that it pointed to are lost. The general rule is to not close a BFD until all operations dependent upon data from the BFD have been completed, or all the data from within the file has been copied. To help with the management of memory, there is a function (bfd_alloc_size) which returns the number of bytes in obstacks associated with the supplied BFD. This could be used to select the greediest open BFD, close it to reclaim the memory, perform some operation and reopen the BFD again, to get a fresh copy of the data structures.

2.5 Initialization

These are the functions that handle initializing a BFD.

2.5.0.1 bfd_init

Synopsis

     void bfd_init(void);

Description This routine must be called before any other BFD function to initialize magical internal data structures.

2.6 Sections

The raw data contained within a BFD is maintained through the section abstraction. A single BFD may have any number of sections. It keeps hold of them by pointing to the first; each one points to the next in the list. Sections are supported in BFD in section.c.

2.6.1 Section input

When a BFD is opened for reading, the section structures are created and attached to the BFD. Each section has a name which describes the section in the outside world—for example, a.out would contain at least three sections, called .text, .data and .bss. Names need not be unique; for example a COFF file may have several sections named .data. Sometimes a BFD will contain more than the “natural” number of sections. A back end may attach other sections containing constructor data, or an application may add a section (using bfd_make_section) to the sections attached to an already open BFD. For example, the linker creates an extra section COMMON for each input file’s BFD to hold information about common storage. The raw data is not necessarily read in when the section descriptor is created. Some targets may leave the data in place until a bfd_get_section_contents call is made. Other back ends may read in all the data at once. For example, an S-record file has to be read once to determine the size of the data. An IEEE-695 file doesn’t contain raw data in sections, but data and relocation expressions intermixed, so the data area has to be parsed to get out the data and relocations.

2.6.2 Section output

To write a new object style BFD, the various sections to be written have to be created. They are attached to the BFD in the same way as input sections; data is written to the

sections using bfd_set_section_contents.

Any program that creates or combines sections (e.g., the assembler and linker) must use the asection fields output_section and output_offset to indicate the file sections to which each section must be written. (If the section is being created from scratch, output_section should probably point to the section itself and output_offset should probably be zero.) The data to be written comes from input sections attached (via output_section pointers) to the output sections. The output section structure can be considered a filter for the input section: the output section determines the vma of the output data and the name, but the input section determines the offset into the output section of the data to be written. E.g., to create a section "O", starting at 0x100, 0x123 long, containing two subsections, "A" at offset 0x0 (i.e., at vma 0x100) and "B" at offset 0x20 (i.e., at vma 0x120) the asection structures would look like:

        section name          "A"
          output_offset   0x00
          size            0x20
          output_section ----------->  section name    "O"
                                  |    vma             0x100
        section name          "B" |    size            0x123
          output_offset   0x20    |
          size            0x103   |
          output_section  --------|

2.6.3 Link orders

The data within a section is stored in a link order. These are much like the fixups in gas. The link order abstraction allows a section to grow and shrink within itself. A link order knows how big it is, and which is the next link order and where the raw data for it is; it also points to a list of relocations which apply to it. The link order is used by the linker to perform relaxing on final code. The compiler creates code which is as big as necessary to make it work without relaxing, and the user can select whether to relax. Sometimes relaxing takes a lot of time. The linker runs around the relocations to see if any are attached to data which can be shrunk, if so it does it on a link order by link order basis.

2.6.4 typedef asection

Here is the section structure:

/* This structure is used for a comdat section, as in PE.  A comdat
section is associated with a particular symbol.  When the linker
sees a comdat section, it keeps only one of the sections with a
given name and associated with a given symbol. */
struct bfd_comdat_info
{
/* The name of the symbol associated with a comdat section.  */
const char *name;
/* The local symbol table index of the symbol associated with a
 comdat section.  This is only meaningful to the object file format
 specific code; it is not an index into the list returned by
 bfd_canonicalize_symtab.  */
long symbol;
/* If this section is being discarded, the linker uses this field
 to point to the input section which is being kept.  */
struct sec *sec;
};
typedef struct sec
{
    /* The name of the section; the name isn't a copy, the pointer is
    the same as that passed to bfd_make_section. */
CONST char *name;
    /* Which section is it; 0..nth.      */
int index;
    /* The next section in the list belonging to the BFD, or NULL. */
struct sec *next;
    /* The field flags contains attributes of the section. Some
       flags are read in from the object file, and some are
       synthesized from other information.  */
flagword flags;
#define SEC_NO_FLAGS   0x000
    /* Tells the OS to allocate space for this section when loading.

This is clear for a section containing debug information
only. */
#define SEC_ALLOC      0x001
/* Tells the OS to load the section from the file when loading.
This is clear for a .bss section. */
#define SEC_LOAD       0x002
/* The section contains data still to be relocated, so there is
some relocation information too. */
#define SEC_RELOC      0x004
#if 0   /* Obsolete ? */
#define SEC_BALIGN     0x008
#endif
/* A signal to the OS that the section contains read only
data. */
#define SEC_READONLY   0x010
/* The section contains code only. */
#define SEC_CODE       0x020
/* The section contains data only. */
#define SEC_DATA       0x040
/* The section will reside in ROM. */
#define SEC_ROM        0x080
/* The section contains constructor information. This section
type is used by the linker to create lists of constructors and
destructors used by g++. When a back end sees a symbol
which should be used in a constructor list, it creates a new
section for the type of name (e.g., __CTOR_LIST__), attaches
the symbol to it, and builds a relocation. To build the lists
of constructors, all the linker has to do is catenate all the
sections called __CTOR_LIST__and relocate the data
contained within - exactly the operations it would peform on
standard data. */
#define SEC_CONSTRUCTOR 0x100
/* The section is a constructor, and should be placed at the
end of the text, data, or bss section(?). */
#define SEC_CONSTRUCTOR_TEXT 0x1100
#define SEC_CONSTRUCTOR_DATA 0x2100
#define SEC_CONSTRUCTOR_BSS  0x3100
/* The section has contents - a data section could be
SEC_ALLOC | SEC_HAS_CONTENTS; a debug section could be

SEC_HAS_CONTENTS */
#define SEC_HAS_CONTENTS 0x200
/* An instruction to the linker to not output the section
even if it has information which would normally be written. */
#define SEC_NEVER_LOAD 0x400
/* The section is a COFF shared library section.  This flag is
only for the linker.  If this type of section appears in
the input file, the linker must copy it to the output file
without changing the vma or size.  FIXME: Although this
was originally intended to be general, it really is COFF
specific (and the flag was renamed to indicate this).  It
might be cleaner to have some more general mechanism to
allow the back end to control what the linker does with
sections. */
#define SEC_COFF_SHARED_LIBRARY 0x800
/* The section contains common symbols (symbols may be defined
multiple times, the value of a symbol is the amount of
space it requires, and the largest symbol value is the one
used).  Most targets have exactly one of these (which we
translate to bfd_com_section_ptr), but ECOFF has two. */
#define SEC_IS_COMMON 0x8000
/* The section contains only debugging information.  For
example, this is set for ELF .debug and .stab sections.
strip tests this flag to see if a section can be
discarded. */
#define SEC_DEBUGGING 0x10000
/* The contents of this section are held in memory pointed to
by the contents field.  This is checked by
bfd_get_section_contents, and the data is retrieved from
memory if appropriate.  */
#define SEC_IN_MEMORY 0x20000
/* The contents of this section are to be excluded by the
linker for executable and shared objects unless those
objects are to be further relocated.  */
#define SEC_EXCLUDE 0x40000
/* The contents of this section are to be sorted by the
based on the address specified in the associated symbol
table.  */
#define SEC_SORT_ENTRIES 0x80000
/* When linking, duplicate sections of the same name should be
discarded, rather than being combined into a single section as

is usually done.  This is similar to how common symbols are
handled.  See SEC_LINK_DUPLICATES below.  */
#define SEC_LINK_ONCE 0x100000
/* If SEC_LINK_ONCE is set, this bitfield describes how the linker
should handle duplicate sections.  */
#define SEC_LINK_DUPLICATES 0x600000
/* This value for SEC_LINK_DUPLICATES means that duplicate
sections with the same name should simply be discarded. */
#define SEC_LINK_DUPLICATES_DISCARD 0x0
/* This value for SEC_LINK_DUPLICATES means that the linker
should warn if there are any duplicate sections, although
it should still only link one copy.  */
#define SEC_LINK_DUPLICATES_ONE_ONLY 0x200000
/* This value for SEC_LINK_DUPLICATES means that the linker
should warn if any duplicate sections are a different size.  */
#define SEC_LINK_DUPLICATES_SAME_SIZE 0x400000
/* This value for SEC_LINK_DUPLICATES means that the linker
should warn if any duplicate sections contain different
contents.  */
#define SEC_LINK_DUPLICATES_SAME_CONTENTS 0x600000
/* This section was created by the linker as part of dynamic
relocation or other arcane processing.  It is skipped when
going through the first-pass output, trusting that someone
else up the line will take care of it later.  */
#define SEC_LINKER_CREATED 0x800000
/* This section should not be subject to garbage collection.  */
#define SEC_KEEP 0x1000000
/* This section contains "short" data, and should be placed
"near" the GP.  */
#define SEC_SMALL_DATA 0x2000000
/* This section contains data which may be shared with other
executables or shared objects.  */
#define SEC_SHARED 0x4000000
/* When a section with this flag is being linked, then if the size of
the input section is less than a page, it should not cross a page
boundary.  If the size of the input section is one page or more, it
should be aligned on a page boundary.  */
#define SEC_BLOCK 0x8000000

/* Conditionally link this section; do not link if there are no
   references found to any symbol in the section.  */
#define SEC_CLINK 0x10000000
/*  End of section flags.  */
/* Some internal packed boolean fields.  */
/* See the vma field.  */
unsigned int user_set_vma : 1;
/* Whether relocations have been processed.  */
unsigned int reloc_done : 1;
/* A mark flag used by some of the linker backends.  */
unsigned int linker_mark : 1;
/* A mark flag used by some linker backends for garbage collection.  */
unsigned int gc_mark : 1;
/* End of internal packed boolean fields.  */
/*  The virtual memory address of the section - where it will be
    at run time.  The symbols are relocated against this.  The
    user_set_vma flag is maintained by bfd; if it's not set, the
    backend can assign addresses (for example, in a.out, where
    the default address for .data is dependent on the specific
    target and various flags).  */
bfd_vma vma;
/*  The load address of the section - where it would be in a
    rom image; really only used for writing section header
    information. */
bfd_vma lma;
 /* The size of the section in octets, as it will be output.
    Contains a value even if the section has no contents (e.g., the
    size of .bss).  This will be filled in after relocation.  */
bfd_size_type _cooked_size;
 /* The original size on disk of the section, in octets.  Normally this
    value is the same as the size, but if some relaxing has
    been done, then this value will be bigger.  */
bfd_size_type _raw_size;

/* If this section is going to be output, then this value is the
   offset in *bytes* into the output section of the first byte in the
   input section (byte ==> smallest addressable unit on the
   target).  In most cases, if this was going to start at the
   100th octet (8-bit quantity) in the output section, this value
   would be 100.  However, if the target byte size is 16 bits
   (bfd_octets_per_byte is "2"), this value would be 50. */
bfd_vma output_offset;
/* The output section through which to map on output. */
struct sec *output_section;
/* The alignment requirement of the section, as an exponent of 2 -
   e.g., 3 aligns to 2^3 (or 8). */
unsigned int alignment_power;
/* If an input section, a pointer to a vector of relocation
   records for the data in this section. */
struct reloc_cache_entry *relocation;
/* If an output section, a pointer to a vector of pointers to
   relocation records for the data in this section. */
struct reloc_cache_entry **orelocation;
/* The number of relocation records in one of the above  */
unsigned reloc_count;
/* Information below is back end specific - and not always used
   or updated.  */
/* File position of section data    */
file_ptr filepos;
/* File position of relocation info */
file_ptr rel_filepos;
/* File position of line data       */
file_ptr line_filepos;

     /* Pointer to data for applications */
PTR userdata;
     /* If the SEC_IN_MEMORY flag is set, this points to the actual
        contents.  */
unsigned char *contents;
     /* Attached line number information */
alent *lineno;
     /* Number of line number records   */
unsigned int lineno_count;
     /* Optional information about a COMDAT entry; NULL if not COMDAT */
struct bfd_comdat_info *comdat;
     /* When a section is being output, this value changes as more
        linenumbers are written out */
file_ptr moving_line_filepos;
     /* What the section number is in the target world  */
int target_index;
PTR used_by_bfd;
     /* If this is a constructor section then here is a list of the
        relocations created to relocate items within it. */
struct relent_chain *constructor_chain;
     /* The BFD which owns the section. */
bfd *owner;
     /* A symbol which points at this section only */
struct symbol_cache_entry *symbol;
struct symbol_cache_entry **symbol_ptr_ptr;
struct bfd_link_order *link_order_head;
struct bfd_link_order *link_order_tail;
} asection ;
 /* These sections are global, and are managed by BFD.  The application

       and target back end are not permitted to change the values in
       these sections.  New code should use the section_ptr macros rather
       than referring directly to the const sections.  The const sections
       may eventually vanish.  */
#define BFD_ABS_SECTION_NAME "*ABS*"
#define BFD_UND_SECTION_NAME "*UND*"
#define BFD_COM_SECTION_NAME "*COM*"
#define BFD_IND_SECTION_NAME "*IND*"
    /* the absolute section */
extern const asection bfd_abs_section;
#define bfd_abs_section_ptr ((asection *) &bfd_abs_section)
#define bfd_is_abs_section(sec) ((sec) == bfd_abs_section_ptr)
    /* Pointer to the undefined section */
extern const asection bfd_und_section;
#define bfd_und_section_ptr ((asection *) &bfd_und_section)
#define bfd_is_und_section(sec) ((sec) == bfd_und_section_ptr)
    /* Pointer to the common section */
extern const asection bfd_com_section;
#define bfd_com_section_ptr ((asection *) &bfd_com_section)
    /* Pointer to the indirect section */
extern const asection bfd_ind_section;
#define bfd_ind_section_ptr ((asection *) &bfd_ind_section)
#define bfd_is_ind_section(sec) ((sec) == bfd_ind_section_ptr)
extern const struct symbol_cache_entry * const bfd_abs_symbol;
extern const struct symbol_cache_entry * const bfd_com_symbol;
extern const struct symbol_cache_entry * const bfd_und_symbol;
extern const struct symbol_cache_entry * const bfd_ind_symbol;
#define bfd_get_section_size_before_reloc(section) \
     ((section)->reloc_done ? (abort (), (bfd_size_type) 1) \
                            : (section)->_raw_size)
#define bfd_get_section_size_after_reloc(section) \
     ((section)->reloc_done ? (section)->_cooked_size \
                            : (abort (), (bfd_size_type) 1))

2.6.5 Section prototypes

These are the functions exported by the section handling part of BFD.

2.6.5.1 bfd_get_section_by_name

Synopsis

asection *bfd_get_section_by_name(bfd *abfd, CONST char *name);

Description Run through abfd and return the one of the asections whose name matches name, otherwise NULL. See Section 2.6 [Sections], page 15, for more information.

This should only be used in special cases; the normal way to process all sections of a given name is to use bfd_map_over_sections and strcmp on the name (or better yet, base it on the section flags or something else) for each section.

2.6.5.2 bfd_make_section_old_way

Synopsis

     asection *bfd_make_section_old_way(bfd *abfd, CONST char *name);

Description Create a new empty section called name and attach it to the end of the chain of sections for the BFD abfd. An attempt to create a section with a name which is already in use returns its pointer without changing the section chain. It has the funny name since this is the way it used to be before it was rewritten.... Possible errors are:

• bfd_error_invalid_operation - If output has already started for this BFD.

• bfd_error_no_memory - If memory allocation fails.

2.6.5.3 bfd_make_section_anyway

Synopsis

     asection *bfd_make_section_anyway(bfd *abfd, CONST char *name);

Description Create a new empty section called name and attach it to the end of the chain of sections for abfd. Create a new section even if there is already a section with that name. Return NULL and set bfd_error on error; possible errors are:

• bfd_error_invalid_operation - If output has already started for abfd.

• bfd_error_no_memory - If memory allocation fails.

2.6.5.4 bfd_make_section

Synopsis

     asection *bfd_make_section(bfd *, CONST char *name);

Description Like bfd_make_section_anyway, but return NULL (without calling bfd set error ()) without changing the section chain if there is already a section named name. If there is an error,

return NULL and set bfd_error.
2.6.5.5 bfd_set_section_flags

Synopsis

     boolean bfd_set_section_flags(bfd *abfd, asection *sec, flagword flags);

Description Set the attributes of the section sec in the BFD abfd to the value flags. Return true on success, false on error. Possible error returns are:

• bfd_error_invalid_operation - The section cannot have one or more of the attributes requested. For example, a .bss section in a.out may not have the SEC_HAS_

    CONTENTS field set.

2.6.5.6 bfd_map_over_sections

Synopsis

     void bfd_map_over_sections(bfd *abfd,
         void (*func)(bfd *abfd,
         asection *sect,
         PTR obj),
         PTR obj);

Description Call the provided function func for each section attached to the BFD abfd, passing obj as an argument. The function will be called as if by

            func(abfd, the_section, obj);

This is the prefered method for iterating over sections; an alternative would be to use a loop:

               section *p;
               for (p = abfd->sections; p != NULL; p = p->next)
                  func(abfd, p, ...)
2.6.5.7 bfd_set_section_size

Synopsis

     boolean bfd_set_section_size(bfd *abfd, asection *sec, bfd_size_type val);

Description Set sec to the size val. If the operation is ok, then true is returned, else false.

Possible error returns:

• bfd_error_invalid_operation - Writing has started to the BFD, so setting the size is invalid.

2.6.5.8 bfd_set_section_contents

Synopsis

     boolean bfd_set_section_contents
        (bfd *abfd,
         asection *section,
         PTR data,
         file_ptr offset,
         bfd_size_type count);

Description Sets the contents of the section section in BFD abfd to the data starting in memory at data. The data is written to the output section starting at offset offset for count octets.

Normally true is returned, else false. Possible error returns are:

  • bfd_error_no_contents - The output section does not have the SEC_HAS_CONTENTS

attribute, so nothing can be written to it.

• and some more too

This routine is front end to the back end function _bfd_set_section_contents.

2.6.5.9 bfd_get_section_contents

Synopsis

     boolean bfd_get_section_contents
        (bfd *abfd, asection *section, PTR location,
         file_ptr offset, bfd_size_type count);

Description Read data from section in BFD abfd into memory starting at location. The data is read at an offset of offset from the start of the input section, and is read for count bytes. If the contents of a constructor with the SEC_CONSTRUCTOR flag set are requested or if the section does not have the SEC_HAS_CONTENTS flag set, then the location is filled with zeroes. If no errors occur, true is returned, else false.

2.6.5.10 bfd_copy_private_section_data

Synopsis

     boolean bfd_copy_private_section_data(bfd *ibfd, asection *isec, bfd *obfd, asection *osec);

Description Copy private section information from isec in the BFD ibfd to the section osec in the BFD obfd. Return true on success, false on error. Possible error returns are:

• bfd_error_no_memory - Not enough memory exists to create private data for osec.

     #define bfd_copy_private_section_data(ibfd, isection, obfd, osection) \
          BFD_SEND (obfd, _bfd_copy_private_section_data, \
                    (ibfd, isection, obfd, osection))
2.6.5.11 _bfd_strip_section_from_output

Synopsis

     void _bfd_strip_section_from_output
        (struct bfd_link_info *info, asection *section);

Description Remove section from the output. If the output section becomes empty, remove it from the output bfd. info may be NULL; if it is not, it is used to decide whether the output section is empty.

2.7 Symbols

BFD tries to maintain as much symbol information as it can when it moves information from file to file. BFD passes information to applications though the asymbol structure. When the application requests the symbol table, BFD reads the table in the native form and translates parts of it into the internal format. To maintain more than the information passed to applications, some targets keep some information “behind the scenes” in a structure only the particular back end knows about. For example, the coff back end keeps the original symbol table structure as well as the canonical structure when a BFD is read in. On output, the coff back end can reconstruct the output symbol table so that no information is lost, even information unique to coff which BFD doesn’t know or understand. If a coff symbol table were read, but were written through an a.out back end, all the coff specific information would be lost. The symbol table of a BFD is not necessarily read in until a canonicalize request is made. Then the BFD back end fills in a table provided by the application with pointers to the canonical information. To output symbols, the application provides BFD with a table of pointers to pointers to asymbols. This allows applications like the linker to output a symbol as it was read, since the “behind the scenes” information will be still available.

2.7.1 Reading symbols

There are two stages to reading a symbol table from a BFD: allocating storage, and the actual reading process. This is an excerpt from an application which reads the symbol table:

              long storage_needed;
              asymbol **symbol_table;
              long number_of_symbols;
              long i;
              storage_needed = bfd_get_symtab_upper_bound (abfd);
              if (storage_needed < 0)
                FAIL
              if (storage_needed == 0) {
                 return ;
              }
              symbol_table = (asymbol **) xmalloc (storage_needed);
                ...
              number_of_symbols =
                 bfd_canonicalize_symtab (abfd, symbol_table);
              if (number_of_symbols < 0)
                FAIL
              for (i = 0; i < number_of_symbols; i++) {
                 process_symbol (symbol_table[i]);
              }

All storage for the symbols themselves is in an objalloc connected to the BFD; it is freed when the BFD is closed.

2.7.2 Writing symbols

Writing of a symbol table is automatic when a BFD open for writing is closed. The application attaches a vector of pointers to pointers to symbols to the BFD being written, and fills in the symbol count. The close and cleanup code reads through the table provided and performs all the necessary operations. The BFD output code must always be provided with an “owned” symbol: one which has come from another BFD, or one which has been created using bfd_make_empty_symbol. Here is an example showing the creation of a symbol table with only one element:

            #include "bfd.h"
            main()
            {

              bfd *abfd;
              asymbol *ptrs[2];
              asymbol *new;
              abfd = bfd_openw("foo","a.out-sunos-big");
              bfd_set_format(abfd, bfd_object);
              new = bfd_make_empty_symbol(abfd);
              new->name = "dummy_symbol";
              new->section = bfd_make_section_old_way(abfd, ".text");
              new->flags = BSF_GLOBAL;
              new->value = 0x12345;
              ptrs[0] = new;
              ptrs[1] = (asymbol *)0;
              bfd_set_symtab(abfd, ptrs, 1);
              bfd_close(abfd);
            }
            ./makesym
            nm foo
            00012345 A dummy_symbol

Many formats cannot represent arbitary symbol information; for instance, the a.out object format does not allow an arbitary number of sections. A symbol pointing to a section which is not one of .text, .data or .bss cannot be described.

2.7.3 Mini Symbols

Mini symbols provide read-only access to the symbol table. They use less memory space, but require more time to access. They can be useful for tools like nm or objdump, which may have to handle symbol tables of extremely large executables. The bfd_read_minisymbols function will read the symbols into memory in an internal form. It will return a void * pointer to a block of memory, a symbol count, and the size of each symbol. The pointer is allocated using malloc, and should be freed by the caller when it is no longer needed. The function bfd_minisymbol_to_symbol will take a pointer to a minisymbol, and a pointer to a structure returned by bfd_make_empty_symbol, and return a asymbol structure. The return value may or may not be the same as the value from bfd_make_empty_symbol which was passed in.

2.7.4 typedef asymbol

An asymbol has the form:

     typedef struct symbol_cache_entry
     {
            /* A pointer to the BFD which owns the symbol. This information
               is necessary so that a back end can work out what additional

information (invisible to the application writer) is carried
with the symbol.
This field is *almost* redundant, since you can use section->owner
instead, except that some symbols point to the global sections
bfd_{abs,com,und}_section.  This could be fixed by making
these globals be per-bfd (or per-target-flavor).  FIXME. */
struct _bfd *the_bfd; /* Use bfd_asymbol_bfd(sym) to access this field. */
/* The text of the symbol. The name is left alone, and not copied; the
application may not alter it. */
CONST char *name;
/* The value of the symbol.  This really should be a union of a
numeric value with a pointer, since some flags indicate that
a pointer to another symbol is stored here.  */
symvalue value;
/* Attributes of a symbol: */
#define BSF_NO_FLAGS    0x00
/* The symbol has local scope; static in C. The value
is the offset into the section of the data. */
#define BSF_LOCAL      0x01
/* The symbol has global scope; initialized data in C. The
value is the offset into the section of the data. */
#define BSF_GLOBAL     0x02
/* The symbol has global scope and is exported. The value is
the offset into the section of the data. */
#define BSF_EXPORT     BSF_GLOBAL /* no real difference */
/* A normal C symbol would be one of:
BSF_LOCAL, BSF_FORT_COMM,  BSF_UNDEFINED or
BSF_GLOBAL */
/* The symbol is a debugging record. The value has an arbitary
meaning, unless BSF_DEBUGGING_RELOC is also set.  */
#define BSF_DEBUGGING  0x08
/* The symbol denotes a function entry point.  Used in ELF,
perhaps others someday.  */
#define BSF_FUNCTION    0x10
/* Used by the linker. */
#define BSF_KEEP        0x20

#define BSF_KEEP_G      0x40
/* A weak global symbol, overridable without warnings by
a regular global symbol of the same name.  */
#define BSF_WEAK        0x80
/* This symbol was created to point to a section, e.g. ELF's
STT_SECTION symbols.  */
#define BSF_SECTION_SYM 0x100
/* The symbol used to be a common symbol, but now it is
allocated. */
#define BSF_OLD_COMMON  0x200
/* The default value for common data. */
#define BFD_FORT_COMM_DEFAULT_VALUE 0
/* In some files the type of a symbol sometimes alters its
location in an output file - ie in coff a ISFCN symbol
which is also C_EXT symbol appears where it was
declared and not at the end of a section.  This bit is set
by the target BFD part to convey this information. */
#define BSF_NOT_AT_END    0x400
/* Signal that the symbol is the label of constructor section. */
#define BSF_CONSTRUCTOR   0x800
/* Signal that the symbol is a warning symbol.  The name is a
warning.  The name of the next symbol is the one to warn about;
if a reference is made to a symbol with the same name as the next
symbol, a warning is issued by the linker. */
#define BSF_WARNING       0x1000
/* Signal that the symbol is indirect.  This symbol is an indirect
pointer to the symbol with the same name as the next symbol. */
#define BSF_INDIRECT      0x2000
/* BSF_FILE marks symbols that contain a file name.  This is used
for ELF STT_FILE symbols.  */
#define BSF_FILE          0x4000
/* Symbol is from dynamic linking information.  */
#define BSF_DYNAMIC       0x8000
/* The symbol denotes a data object.  Used in ELF, and perhaps
others someday.  */
#define BSF_OBJECT        0x10000

            /* This symbol is a debugging symbol.  The value is the offset
               into the section of the data.  BSF_DEBUGGING should be set
               as well.  */
     #define BSF_DEBUGGING_RELOC 0x20000
       flagword flags;
            /* A pointer to the section to which this symbol is
               relative.  This will always be non NULL, there are special
               sections for undefined and absolute symbols.  */
       struct sec *section;
            /* Back end special data.  */
       union
         {
           PTR p;
           bfd_vma i;
         } udata;
     } asymbol;

2.7.5 Symbol handling functions

2.7.5.1 bfd_get_symtab_upper_bound

Description Return the number of bytes required to store a vector of pointers to asymbols for all the symbols in the BFD abfd, including a terminal NULL pointer. If there are no symbols in the BFD, then return 0. If an error occurs, return -1.

     #define bfd_get_symtab_upper_bound(abfd) \
          BFD_SEND (abfd, _bfd_get_symtab_upper_bound, (abfd))
2.7.5.2 bfd_is_local_label

Synopsis

     boolean bfd_is_local_label(bfd *abfd, asymbol *sym);

Description Return true if the given symbol sym in the BFD abfd is a compiler generated local label, else return false.

2.7.5.3 bfd_is_local_label_name

Synopsis

     boolean bfd_is_local_label_name(bfd *abfd, const char *name);

Description Return true if a symbol with the name name in the BFD abfd is a compiler generated local label, else return false. This just checks whether the name has the form of a local label.

     #define bfd_is_local_label_name(abfd, name) \
          BFD_SEND (abfd, _bfd_is_local_label_name, (abfd, name))

2.7.5.4 bfd_canonicalize_symtab

Description Read the symbols from the BFD abfd, and fills in the vector location with pointers to the symbols and a trailing NULL. Return the actual number of symbol pointers, not including the NULL.

     #define bfd_canonicalize_symtab(abfd, location) \
          BFD_SEND (abfd, _bfd_canonicalize_symtab,\
                       (abfd, location))
2.7.5.5 bfd_set_symtab

Synopsis

     boolean bfd_set_symtab (bfd *abfd, asymbol **location, unsigned int count);

Description Arrange that when the output BFD abfd is closed, the table location of count pointers to symbols will be written.

2.7.5.6 bfd_print_symbol_vandf

Synopsis

     void bfd_print_symbol_vandf(PTR file, asymbol *symbol);

Description Print the value and flags of the symbol supplied to the stream file.

2.7.5.7 bfd_make_empty_symbol

Description Create a new asymbol structure for the BFD abfd and return a pointer to it. This routine is necessary because each back end has private information surrounding the asymbol. Building your own asymbol and pointing to it will not create the private information, and will cause problems later on.

     #define bfd_make_empty_symbol(abfd) \
          BFD_SEND (abfd, _bfd_make_empty_symbol, (abfd))
2.7.5.8 bfd_make_debug_symbol

Description Create a new asymbol structure for the BFD abfd, to be used as a debugging symbol. Further details of its use have yet to be worked out.

     #define bfd_make_debug_symbol(abfd,ptr,size) \
             BFD_SEND (abfd, _bfd_make_debug_symbol, (abfd, ptr, size))
2.7.5.9 bfd_decode_symclass

Description Return a character corresponding to the symbol class of symbol, or ’?’ for an unknown class. Synopsis

     int bfd_decode_symclass(asymbol *symbol);

2.7.5.10 bfd_is_undefined_symclass

Description Returns non-zero if the class symbol returned by bfd decode symclass represents an undefined symbol. Returns zero otherwise. Synopsis

     boolean bfd_is_undefined_symclass (int symclass);
2.7.5.11 bfd_symbol_info

Description Fill in the basic info about symbol that nm needs. Additional info may be added by the back-ends after calling this function. Synopsis

     void bfd_symbol_info(asymbol *symbol, symbol_info *ret);
2.7.5.12 bfd_copy_private_symbol_data

Synopsis

     boolean bfd_copy_private_symbol_data(bfd *ibfd, asymbol *isym, bfd *obfd, asymbol *osym);

Description Copy private symbol information from isym in the BFD ibfd to the symbol osym in the BFD obfd. Return true on success, false on error. Possible error returns are:

• bfd_error_no_memory - Not enough memory exists to create private data for osec.

     #define bfd_copy_private_symbol_data(ibfd, isymbol, obfd, osymbol) \
          BFD_SEND (obfd, _bfd_copy_private_symbol_data, \
                    (ibfd, isymbol, obfd, osymbol))

2.8 Archives

Description An archive (or library) is just another BFD. It has a symbol table, although there’s not much a user program will do with it. The big difference between an archive BFD and an ordinary BFD is that the archive doesn’t have sections. Instead it has a chain of BFDs that are considered its contents. These BFDs can be manipulated like any other. The BFDs contained in an archive opened for reading will all be opened for reading. You may put either input or output BFDs into an archive opened for output; they will be handled correctly when the archive is closed. Use bfd_openr_next_archived_file to step through the contents of an archive opened for input. You don’t have to read the entire archive if you don’t want to! Read it until you find what you want. Archive contents of output BFDs are chained through the next pointer in a BFD. The first one is findable through the archive_head slot of the archive. Set it with bfd_set_ archive_head (q.v.). A given BFD may be in only one open output archive at a time. As expected, the BFD archive code is more general than the archive code of any given environment. BFD archives may contain files of different formats (e.g., a.out and coff) and even different architectures. You may even place archives recursively into archives!

This can cause unexpected confusion, since some archive formats are more expressive than others. For instance, Intel COFF archives can preserve long filenames; SunOS a.out archives cannot. If you move a file from the first to the second format and back again, the filename may be truncated. Likewise, different a.out environments have different conventions as to how they truncate filenames, whether they preserve directory names in filenames, etc. When interoperating with native tools, be sure your files are homogeneous. Beware: most of these formats do not react well to the presence of spaces in filenames. We do the best we can, but can’t always handle this case due to restrictions in the format of archives. Many Unix utilities are braindead in regards to spaces and such in filenames anyway, so this shouldn’t be much of a restriction. Archives are supported in BFD in archive.c.

2.8.0.1 bfd_get_next_mapent

Synopsis

     symindex bfd_get_next_mapent(bfd *abfd, symindex previous, carsym **sym);

Description Step through archive abfd’s symbol table (if it has one). Successively update sym with the next symbol’s information, returning that symbol’s (internal) index into the symbol table. Supply BFD_NO_MORE_SYMBOLS as the previous entry to get the first one; returns BFD_NO_ MORE_SYMBOLS when you’ve already got the last one. A carsym is a canonical archive symbol. The only user-visible element is its name, a nullterminated string.

2.8.0.2 bfd_set_archive_head

Synopsis

     boolean bfd_set_archive_head(bfd *output, bfd *new_head);

Description Set the head of the chain of BFDs contained in the archive output to new head.

2.8.0.3 bfd_openr_next_archived_file

Synopsis

     bfd *bfd_openr_next_archived_file(bfd *archive, bfd *previous);

Description Provided a BFD, archive, containing an archive and NULL, open an input BFD on the first contained element and returns that. Subsequent calls should pass the archive and the previous return value to return a created BFD to the next contained element. NULL is returned when there are no more.

2.9 File formats

A format is a BFD concept of high level file contents type. The formats supported by BFD are:

  • bfd_object

The BFD may contain data, symbols, relocations and debug info.

  • bfd_archive

The BFD contains other BFDs and an optional index.

  • bfd_core

The BFD contains the result of an executable core dump.

2.9.0.1 bfd_check_format

Synopsis

     boolean bfd_check_format(bfd *abfd, bfd_format format);

Description Verify if the file attached to the BFD abfd is compatible with the format format (i.e., one

of bfd_object, bfd_archive or bfd_core).

If the BFD has been set to a specific target before the call, only the named target and format combination is checked. If the target has not been set, or has been set to default, then all the known target backends is interrogated to determine a match. If the default target matches, it is used. If not, exactly one target must recognize the file, or an error results. The function returns true on success, otherwise false with one of the following error codes:

  • bfd_error_invalid_operation - if format is not one of bfd_object, bfd_archive
    or bfd_core.

• bfd_error_system_call - if an error occured during a read - even some file mismatches can cause bfd error system calls.

• file_not_recognised - none of the backends recognised the file format.

  • bfd_error_file_ambiguously_recognized - more than one backend recognised the

file format.

2.9.0.2 bfd_check_format_matches

Synopsis

     boolean bfd_check_format_matches(bfd *abfd, bfd_format format, char ***matching);

Description Like bfd_check_format, except when it returns false with bfd_errno set to bfd_error_ file_ambiguously_recognized. In that case, if matching is not NULL, it will be filled in with a NULL-terminated list of the names of the formats that matched, allocated with malloc. Then the user may choose a format and try again. When done with the list that matching points to, the caller should free it.

2.9.0.3 bfd_set_format

Synopsis

     boolean bfd_set_format(bfd *abfd, bfd_format format);

Description This function sets the file format of the BFD abfd to the format format. If the target set in the BFD does not support the format requested, the format is invalid, or the BFD is not open for writing, then an error occurs.

2.9.0.4 bfd_format_string

Synopsis

     CONST char *bfd_format_string(bfd_format format);

Description Return a pointer to a const string invalid, object, archive, core, or unknown, depending upon the value of format.

2.10 Relocations

BFD maintains relocations in much the same way it maintains symbols: they are left alone until required, then read in en-mass and translated into an internal form. A common routine bfd_perform_relocation acts upon the canonical form to do the fixup. Relocations are maintained on a per section basis, while symbols are maintained on a per BFD basis. All that a back end has to do to fit the BFD interface is to create a struct reloc_cache_ entry for each relocation in a particular section, and fill in the right bits of the structures.

2.10.1 typedef arelent

This is the structure of a relocation entry:

     typedef enum bfd_reloc_status
     {
            /* No errors detected */
       bfd_reloc_ok,
            /* The relocation was performed, but there was an overflow. */
       bfd_reloc_overflow,
            /* The address to relocate was not within the section supplied. */
       bfd_reloc_outofrange,
            /* Used by special functions */
       bfd_reloc_continue,
            /* Unsupported relocation size requested. */
       bfd_reloc_notsupported,
            /* Unused */
       bfd_reloc_other,
            /* The symbol to relocate against was undefined. */
       bfd_reloc_undefined,
            /* The relocation was performed, but may not be ok - presently
               generated only when linking i960 coff files with i960 b.out
               symbols.  If this type is returned, the error_message argument

               to bfd_perform_relocation will be set.  */
       bfd_reloc_dangerous
      }
      bfd_reloc_status_type;

     typedef struct reloc_cache_entry
     {
            /* A pointer into the canonical table of pointers  */
       struct symbol_cache_entry **sym_ptr_ptr;
            /* offset in section */
       bfd_size_type address;
            /* addend for relocation value */
       bfd_vma addend;
            /* Pointer to how to perform the required relocation */
       reloc_howto_type *howto;
     } arelent;

Description Here is a description of each of the fields within an arelent:

  • sym_ptr_ptr

The symbol table pointer points to a pointer to the symbol associated with the relocation request. It is the pointer into the table returned by the back end’s get_symtab action. See Section 2.7 [Symbols], page 27. The symbol is referenced through a pointer to a pointer so that tools like the linker can fix up all the symbols of the same name by modifying only one pointer. The relocation routine looks in the symbol and uses the base of the section the symbol is attached to and the value of the symbol as the initial relocation offset. If the symbol pointer is zero, then the section provided is looked up.

  • address

The address field gives the offset in bytes from the base of the section data which owns the relocation record to the first byte of relocatable information. The actual data relocated will be relative to this point; for example, a relocation type which modifies the bottom two bytes of a four byte word would not touch the first byte pointed to in a big endian world.

  • addend

The addend is a value provided by the back end to be added (!) to the relocation offset. Its interpretation is dependent upon the howto. For example, on the 68k the code:

             char foo[];
             main()
                     {
                     return foo[0x12345678];
                     }

Could be compiled into:

             linkw fp,#-4

        moveb @#12345678,d0
        extbl d0
        unlk fp
        rts

This could create a reloc pointing to foo, but leave the offset in the data, something like:

RELOCATION RECORDS FOR [.text]:
offset   type      value
00000006 32        _foo
00000000 4e56 fffc          ; linkw fp,#-4
00000004 1039 1234 5678     ; moveb @#12345678,d0
0000000a 49c0               ; extbl d0
0000000c 4e5e               ; unlk fp
0000000e 4e75               ; rts

Using coff and an 88k, some instructions don’t have enough space in them to represent the full address range, and pointers have to be loaded in two parts. So you’d get something like:

        or.u     r13,r0,hi16(_foo+0x12345678)
        ld.b     r2,r13,lo16(_foo+0x12345678)
        jmp      r1

This should create two relocs, both pointing to _foo, and with 0x12340000 in their addend field. The data would consist of:

RELOCATION RECORDS FOR [.text]:
offset   type      value
00000002 HVRT16    _foo+0x12340000
00000006 LVRT16    _foo+0x12340000
00000000 5da05678           ; or.u r13,r0,0x5678
00000004 1c4d5678           ; ld.b r2,r13,0x5678
00000008 f400c001           ; jmp r1

The relocation routine digs out the value from the data, adds it to the addend to get the original offset, and then adds the value of _foo. Note that all 32 bits have to be kept around somewhere, to cope with carry from bit 15 to bit 16.

One further example is the sparc and the a.out format. The sparc has a similar problem to the 88k, in that some instructions don’t have room for an entire offset, but on the sparc the parts are created in odd sized lumps. The designers of the a.out format chose to not use the data within the section for storing part of the offset; all the offset is kept within the reloc. Anything in the data should be ignored.

        save %sp,-112,%sp
        sethi %hi(_foo+0x12345678),%g2
        ldsb [%g2+%lo(_foo+0x12345678)],%i0
        ret
        restore

Both relocs contain a pointer to foo, and the offsets contain junk.

RELOCATION RECORDS FOR [.text]:
offset   type      value
00000004 HI22      _foo+0x12345678
00000008 LO10      _foo+0x12345678

00000000 9de3bf90     ; save %sp,-112,%sp
00000004 05000000     ; sethi %hi(_foo+0),%g2
00000008 f048a000     ; ldsb [%g2+%lo(_foo+0)],%i0
0000000c 81c7e008     ; ret
00000010 81e80000     ; restore
• howto

The howto field can be imagined as a relocation instruction. It is a pointer to a structure which contains information on what to do with all of the other information in the reloc record and data section. A back end would normally have a relocation instruction set and turn relocations into pointers to the correct structure on input - but it would be possible to create each howto field on demand.

2.10.1.1 enum complain_overflow

Indicates what sort of overflow checking should be done when performing a relocation.

enum complain_overflow
{
       /* Do not complain on overflow. */
  complain_overflow_dont,
       /* Complain if the bitfield overflows, whether it is considered
          as signed or unsigned. */
  complain_overflow_bitfield,
       /* Complain if the value overflows when considered as signed
          number. */
  complain_overflow_signed,
       /* Complain if the value overflows when considered as an
          unsigned number. */
  complain_overflow_unsigned
};
2.10.1.2 reloc_howto_type

The reloc_howto_type is a structure which contains all the information that libbfd needs to know to tie up a back end’s data.

struct symbol_cache_entry;             /* Forward declaration */
struct reloc_howto_struct
{
       /*  The type field has mainly a documentary use - the back end can
           do what it wants with it, though normally the back end's
           external idea of what a reloc number is stored
           in this field. For example, a PC relative word relocation
           in a coff environment has the type 023 - because that's

what the outside world calls a R_PCRWORD reloc. */
unsigned int type;
/*  The value the final relocation is shifted right by. This drops
unwanted data from the relocation.  */
unsigned int rightshift;
/*  The size of the item to be relocated.  This is *not* a
power-of-two measure.  To get the number of bytes operated
on by a type of relocation, use bfd_get_reloc_size.  */
int size;
/*  The number of bits in the item to be relocated.  This is used
when doing overflow checking.  */
unsigned int bitsize;
/*  Notes that the relocation is relative to the location in the
data section of the addend. The relocation function will
subtract from the relocation value the address of the location
being relocated. */
boolean pc_relative;
/*  The bit position of the reloc value in the destination.
The relocated value is left shifted by this amount. */
unsigned int bitpos;
/* What type of overflow error should be checked for when
relocating. */
enum complain_overflow complain_on_overflow;
/* If this field is non null, then the supplied function is
called rather than the normal function. This allows really
strange relocation methods to be accomodated (e.g., i960 callj
instructions). */
bfd_reloc_status_type (*special_function)
                        PARAMS ((bfd *abfd,
                                 arelent *reloc_entry,
                                 struct symbol_cache_entry *symbol,
                                 PTR data,
                                 asection *input_section,
                                 bfd *output_bfd,
                                 char **error_message));
/* The textual name of the relocation type. */
char *name;
/* Some formats record a relocation addend in the section contents
rather than with the relocation.  For ELF formats this is the
distinction between USE_REL and USE_RELA (though the code checks

for USE_REL == 1/0).  The value of this field is TRUE if the
addend is recorded with the section contents; when performing a
partial link (ld -r) the section contents (the data) will be
modified.  The value of this field is FALSE if addends are
recorded with the relocation (in arelent.addend); when performing
a partial link the relocation will be modified.
All relocations for all ELF USE_RELA targets should set this field
to FALSE (values of TRUE should be looked on with suspicion).
However, the converse is not true: not all relocations of all ELF
USE_REL targets set this field to TRUE.  Why this is so is peculiar
to each particular target.  For relocs that aren't used in partial
links (e.g. GOT stuff) it doesn't matter what this is set to.  */
boolean partial_inplace;
/* The src_mask selects which parts of the read in data
are to be used in the relocation sum.  E.g., if this was an 8 bit
byte of data which we read and relocated, this would be
0x000000ff. When we have relocs which have an addend, such as
sun4 extended relocs, the value in the offset part of a
relocating field is garbage so we never use it. In this case
the mask would be 0x00000000. */
bfd_vma src_mask;
/* The dst_mask selects which parts of the instruction are replaced
into the instruction. In most cases src_mask == dst_mask,
except in the above special case, where dst_mask would be
0x000000ff, and src_mask would be 0x00000000.   */
bfd_vma dst_mask;
/* When some formats create PC relative instructions, they leave
the value of the pc of the place being relocated in the offset
slot of the instruction, so that a PC relative relocation can
be made just by adding in an ordinary offset (e.g., sun3 a.out).
Some formats leave the displacement part of an instruction
empty (e.g., m88k bcs); this flag signals the fact.*/
boolean pcrel_offset;
};
2.10.1.3 The HOWTO Macro

Description The HOWTO define is horrible and will go away.

#define HOWTO(C, R,S,B, P, BI, O, SF, NAME, INPLACE, MASKSRC, MASKDST, PC) \
{(unsigned)C,R,S,B, P, BI, O,SF,NAME,INPLACE,MASKSRC,MASKDST,PC}

Description And will be replaced with the totally magic way. But for the moment, we are compatible, so do it this way.

#define NEWHOWTO( FUNCTION, NAME,SIZE,REL,IN) HOWTO(0,0,SIZE,0,REL,0,complain_overflow_dont,FUNCTION, NAME,false,0,0,IN)

Description This is used to fill in an empty howto entry in an array.

     #define EMPTY_HOWTO(C) \
       HOWTO((C),0,0,0,false,0,complain_overflow_dont,NULL,NULL,false,0,0,false)

Description Helper routine to turn a symbol into a relocation value.

     #define HOWTO_PREPARE(relocation, symbol)      \
       {                                            \
       if (symbol != (asymbol *)NULL) {             \
         if (bfd_is_com_section (symbol->section)) { \
           relocation = 0;                          \
         }                                          \
         else {                                     \
           relocation = symbol->value;              \
         }                                          \
       }                                            \
     }
2.10.1.4 bfd_get_reloc_size

Synopsis

     unsigned int bfd_get_reloc_size (reloc_howto_type *);

Description For a reloc howto type that operates on a fixed number of bytes, this returns the number of bytes operated on.

2.10.1.5 arelent_chain

Description How relocs are tied together in an asection:

     typedef struct relent_chain {
       arelent relent;
       struct   relent_chain *next;
     } arelent_chain;
2.10.1.6 bfd_check_overflow

Synopsis

     bfd_reloc_status_type
     bfd_check_overflow
        (enum complain_overflow how,
         unsigned int bitsize,
         unsigned int rightshift,
         unsigned int addrsize,
         bfd_vma relocation);

Description Perform overflow checking on relocation which has bitsize significant bits and will be shifted right by rightshift bits, on a machine with addresses containing addrsize significant bits.

The result is either of bfd_reloc_ok or bfd_reloc_overflow.

2.10.1.7 bfd_perform_relocation

Synopsis

     bfd_reloc_status_type
     bfd_perform_relocation
        (bfd *abfd,
         arelent *reloc_entry,
         PTR data,
         asection *input_section,
         bfd *output_bfd,
         char **error_message);

Description If output bfd is supplied to this function, the generated image will be relocatable; the relocations are copied to the output file after they have been changed to reflect the new state of the world. There are two ways of reflecting the results of partial linkage in an output file: by modifying the output data in place, and by modifying the relocation record. Some native formats (e.g., basic a.out and basic coff) have no way of specifying an addend in the relocation type, so the addend has to go in the output data. This is no big deal since in these formats the output data slot will always be big enough for the addend. Complex reloc types with addends were invented to solve just this problem. The error message argument is set to an error message if this return bfd_reloc_dangerous.

2.10.1.8 bfd_install_relocation

Synopsis

     bfd_reloc_status_type
     bfd_install_relocation
        (bfd *abfd,
         arelent *reloc_entry,
         PTR data, bfd_vma data_start,
         asection *input_section,
         char **error_message);

Description This looks remarkably like bfd_perform_relocation, except it does not expect that the section contents have been filled in. I.e., it’s suitable for use when creating, rather than applying a relocation.

For now, this function should be considered reserved for the assembler.

2.11 The howto manager

When an application wants to create a relocation, but doesn’t know what the target machine might call it, it can find out by using this bit of code.

2.11.0.1 bfd_reloc_code_type

Description The insides of a reloc code. The idea is that, eventually, there will be one enumerator for every type of relocation we ever do. Pass one of these values to bfd_reloc_type_lookup, and it’ll return a howto pointer.

This does mean that the application must determine the correct enumerator value; you can’t get a howto pointer from a random set of attributes.

Here are the possible values for enum bfd_reloc_code_real:

BFD RELOC 64 BFD RELOC 32 BFD RELOC 26 BFD RELOC 24 BFD RELOC 16 BFD RELOC 14 BFD RELOC 8

Basic absolute relocations of N bits.

BFD RELOC 64 PCREL BFD RELOC 32 PCREL BFD RELOC 24 PCREL BFD RELOC 16 PCREL BFD RELOC 12 PCREL BFD RELOC 8 PCREL

PC-relative relocations. Sometimes these are relative to the address of the relocation itself; sometimes they are relative to the start of the section containing the relocation. It depends on the specific target.

The 24-bit relocation is used in some Intel 960 configurations.

BFD RELOC 32 GOT PCREL BFD RELOC 16 GOT PCREL BFD RELOC 8 GOT PCREL BFD RELOC 32 GOTOFF BFD RELOC 16 GOTOFF BFD RELOC LO16 GOTOFF BFD RELOC HI16 GOTOFF BFD RELOC HI16 S GOTOFF BFD RELOC 8 GOTOFF BFD RELOC 32 PLT PCREL BFD RELOC 24 PLT PCREL BFD RELOC 16 PLT PCREL BFD RELOC 8 PLT PCREL BFD RELOC 32 PLTOFF BFD RELOC 16 PLTOFF BFD RELOC LO16 PLTOFF BFD RELOC HI16 PLTOFF BFD RELOC HI16 S PLTOFF BFD RELOC 8 PLTOFF

For ELF.

BFD RELOC 68K GLOB DAT BFD RELOC 68K JMP SLOT BFD RELOC 68K RELATIVE

Relocations used by 68K ELF.

BFD RELOC 32 BASEREL BFD RELOC 16 BASEREL BFD RELOC LO16 BASEREL BFD RELOC HI16 BASEREL BFD RELOC HI16 S BASEREL BFD RELOC 8 BASEREL BFD RELOC RVA

Linkage-table relative.

BFD RELOC 8 FFnn

Absolute 8-bit relocation, but used to form an address like 0xFFnn.

BFD RELOC 32 PCREL S2 BFD RELOC 16 PCREL S2 BFD RELOC 23 PCREL S2

These PC-relative relocations are stored as word displacements – i.e., byte displacements shifted right two bits. The 30-bit word displacement (<<32 PCREL S2>> – 32 bits, shifted 2) is used on the SPARC. (SPARC tools generally refer to this as <<WDISP30>>.) The signed 16-bit displacement is used on the MIPS, and the 23-bit displacement is used on the Alpha.

BFD RELOC HI22 BFD RELOC LO10

High 22 bits and low 10 bits of 32-bit value, placed into lower bits of the target word. These are used on the SPARC.

BFD RELOC GPREL16 BFD RELOC GPREL32

For systems that allocate a Global Pointer register, these are displacements off that register. These relocation types are handled specially, because the value the register will have is decided relatively late.

BFD RELOC I960 CALLJ

Reloc types used for i960/b.out.

BFD RELOC NONE BFD RELOC SPARC WDISP22 BFD RELOC SPARC22 BFD RELOC SPARC13 BFD RELOC SPARC GOT10 BFD RELOC SPARC GOT13 BFD RELOC SPARC GOT22 BFD RELOC SPARC PC10 BFD RELOC SPARC PC22 BFD RELOC SPARC WPLT30 BFD RELOC SPARC COPY BFD RELOC SPARC GLOB DAT BFD RELOC SPARC JMP SLOT BFD RELOC SPARC RELATIVE BFD RELOC SPARC UA32

SPARC ELF relocations. There is probably some overlap with other relocation types already defined.

BFD RELOC SPARC BASE13 BFD RELOC SPARC BASE22

I think these are specific to SPARC a.out (e.g., Sun 4).

BFD RELOC SPARC 64 BFD RELOC SPARC 10 BFD RELOC SPARC 11 BFD RELOC SPARC OLO10 BFD RELOC SPARC HH22 BFD RELOC SPARC HM10 BFD RELOC SPARC LM22 BFD RELOC SPARC PC HH22 BFD RELOC SPARC PC HM10 BFD RELOC SPARC PC LM22 BFD RELOC SPARC WDISP16 BFD RELOC SPARC WDISP19 BFD RELOC SPARC 7 BFD RELOC SPARC 6 BFD RELOC SPARC 5 BFD RELOC SPARC DISP64 BFD RELOC SPARC PLT64 BFD RELOC SPARC HIX22 BFD RELOC SPARC LOX10 BFD RELOC SPARC H44 BFD RELOC SPARC M44 BFD RELOC SPARC L44 BFD RELOC SPARC REGISTER

SPARC64 relocations

BFD RELOC SPARC REV32

SPARC little endian relocation

BFD RELOC ALPHA GPDISP HI16

Alpha ECOFF and ELF relocations. Some of these treat the symbol or "addend" in some special way. For GPDISP HI16 ("gpdisp") relocations, the symbol is ignored when writing; when reading, it will be the absolute section symbol. The addend is the displacement in bytes of the "lda" instruction from the "ldah" instruction (which is at the address of this reloc).

BFD RELOC ALPHA GPDISP LO16

For GPDISP LO16 ("ignore") relocations, the symbol is handled as with GPDISP HI16 relocs. The addend is ignored when writing the relocations out, and is filled in with the file’s GP value on reading, for convenience.

BFD RELOC ALPHA GPDISP

The ELF GPDISP relocation is exactly the same as the GPDISP HI16 relocation except that there is no accompanying GPDISP LO16 relocation.

BFD RELOC ALPHA LITERAL BFD RELOC ALPHA ELF LITERAL BFD RELOC ALPHA LITUSE

The Alpha LITERAL/LITUSE relocs are produced by a symbol reference; the assembler turns it into a LDQ instruction to load the address of the symbol, and then fills in a register in the real instruction. The LITERAL reloc, at the LDQ instruction, refers to the .lita section symbol. The addend is ignored when writing, but is filled in with the file’s GP value on reading, for convenience, as with the GPDISP LO16 reloc. The ELF LITERAL reloc is somewhere between 16 GOTOFF and GPDISP LO16. It should refer to the symbol to be referenced, as with 16 GOTOFF, but it generates output not based on the position within the .got section, but relative to the GP value chosen for the file during the final link stage. The LITUSE reloc, on the instruction using the loaded address, gives information to the linker that it might be able to use to optimize away some literal section references. The symbol is ignored (read as the absolute section symbol), and the "addend" indicates the type of instruction using the register: 1 - "memory" fmt insn 2 - bytemanipulation (byte offset reg) 3 - jsr (target of branch) The GNU linker currently doesn’t do any of this optimizing.

BFD RELOC ALPHA USER LITERAL BFD RELOC ALPHA USER LITUSE BASE BFD RELOC ALPHA USER LITUSE BYTOFF BFD RELOC ALPHA USER LITUSE JSR BFD RELOC ALPHA USER GPDISP BFD RELOC ALPHA USER GPRELHIGH BFD RELOC ALPHA USER GPRELLOW

The BFD RELOC ALPHA USER * relocations are used by the assembler to process the explicit !<reloc>!sequence relocations, and are mapped into the normal relocations at the end of processing.

BFD RELOC ALPHA HINT

The HINT relocation indicates a value that should be filled into the "hint" field of a jmp/jsr/ret instruction, for possible branch- prediction logic which may be provided on some processors.

BFD RELOC ALPHA LINKAGE

The LINKAGE relocation outputs a linkage pair in the object file, which is filled by the linker.

BFD RELOC ALPHA CODEADDR

The CODEADDR relocation outputs a STO CA in the object file, which is filled by the linker.

BFD RELOC MIPS JMP

Bits 27..2 of the relocation address shifted right 2 bits; simple reloc otherwise.

BFD RELOC MIPS16 JMP

The MIPS16 jump instruction.

BFD RELOC MIPS16 GPREL

MIPS16 GP relative reloc.

BFD RELOC HI16

High 16 bits of 32-bit value; simple reloc.

BFD RELOC HI16 S

High 16 bits of 32-bit value but the low 16 bits will be sign extended and added to form the final result. If the low 16 bits form a negative number, we need to add one to the high value to compensate for the borrow when the low bits are added.

BFD RELOC LO16

Low 16 bits.

BFD RELOC PCREL HI16 S

Like BFD RELOC HI16 S, but PC relative.

BFD RELOC PCREL LO16

Like BFD RELOC LO16, but PC relative.

BFD RELOC MIPS GPREL

Relocation relative to the global pointer.

BFD RELOC MIPS LITERAL

Relocation against a MIPS literal section.

BFD RELOC MIPS GOT16 BFD RELOC MIPS CALL16 BFD RELOC MIPS GPREL32 BFD RELOC MIPS GOT HI16 BFD RELOC MIPS GOT LO16 BFD RELOC MIPS CALL HI16 BFD RELOC MIPS CALL LO16 BFD RELOC MIPS SUB BFD RELOC MIPS GOT PAGE BFD RELOC MIPS GOT OFST BFD RELOC MIPS GOT DISP

MIPS ELF relocations.

BFD RELOC 386 GOT32 BFD RELOC 386 PLT32 BFD RELOC 386 COPY BFD RELOC 386 GLOB DAT BFD RELOC 386 JUMP SLOT BFD RELOC 386 RELATIVE BFD RELOC 386 GOTOFF BFD RELOC 386 GOTPC

i386/elf relocations

BFD RELOC NS32K IMM 8 BFD RELOC NS32K IMM 16 BFD RELOC NS32K IMM 32 BFD RELOC NS32K IMM 8 PCREL BFD RELOC NS32K IMM 16 PCREL BFD RELOC NS32K IMM 32 PCREL BFD RELOC NS32K DISP 8 BFD RELOC NS32K DISP 16 BFD RELOC NS32K DISP 32 BFD RELOC NS32K DISP 8 PCREL BFD RELOC NS32K DISP 16 PCREL BFD RELOC NS32K DISP 32 PCREL

ns32k relocations

BFD RELOC PJ CODE HI16 BFD RELOC PJ CODE LO16 BFD RELOC PJ CODE DIR16 BFD RELOC PJ CODE DIR32 BFD RELOC PJ CODE REL16 BFD RELOC PJ CODE REL32

Picojava relocs. Not all of these appear in object files.

BFD RELOC PPC B26 BFD RELOC PPC BA26 BFD RELOC PPC TOC16 BFD RELOC PPC B16 BFD RELOC PPC B16 BRTAKEN BFD RELOC PPC B16 BRNTAKEN BFD RELOC PPC BA16 BFD RELOC PPC BA16 BRTAKEN BFD RELOC PPC BA16 BRNTAKEN BFD RELOC PPC COPY BFD RELOC PPC GLOB DAT BFD RELOC PPC JMP SLOT BFD RELOC PPC RELATIVE BFD RELOC PPC LOCAL24PC BFD RELOC PPC EMB NADDR32 BFD RELOC PPC EMB NADDR16 BFD RELOC PPC EMB NADDR16 LO BFD RELOC PPC EMB NADDR16 HI BFD RELOC PPC EMB NADDR16 HA BFD RELOC PPC EMB SDAI16 BFD RELOC PPC EMB SDA2I16 BFD RELOC PPC EMB SDA2REL BFD RELOC PPC EMB SDA21 BFD RELOC PPC EMB MRKREF BFD RELOC PPC EMB RELSEC16 BFD RELOC PPC EMB RELST LO BFD RELOC PPC EMB RELST HI BFD RELOC PPC EMB RELST HA BFD RELOC PPC EMB BIT FLD BFD RELOC PPC EMB RELSDA

Power(rs6000) and PowerPC relocations.

BFD RELOC I370 D12

IBM 370/390 relocations

BFD RELOC CTOR

The type of reloc used to build a contructor table - at the moment probably a 32 bit wide absolute relocation, but the target can choose. It generally does map to one of the other relocation types.

BFD RELOC ARM PCREL BRANCH

ARM 26 bit pc-relative branch. The lowest two bits must be zero and are not stored in the instruction.

BFD RELOC ARM PCREL BLX

ARM 26 bit pc-relative branch. The lowest bit must be zero and is not stored in the instruction. The 2nd lowest bit comes from a 1 bit field in the instruction.

BFD RELOC THUMB PCREL BLX

Thumb 22 bit pc-relative branch. The lowest bit must be zero and is not stored in the instruction. The 2nd lowest bit comes from a 1 bit field in the instruction.

BFD RELOC ARM IMMEDIATE BFD RELOC ARM ADRL IMMEDIATE BFD RELOC ARM OFFSET IMM BFD RELOC ARM SHIFT IMM BFD RELOC ARM SWI BFD RELOC ARM MULTI BFD RELOC ARM CP OFF IMM BFD RELOC ARM ADR IMM BFD RELOC ARM LDR IMM BFD RELOC ARM LITERAL BFD RELOC ARM IN POOL BFD RELOC ARM OFFSET IMM8 BFD RELOC ARM HWLITERAL BFD RELOC ARM THUMB ADD BFD RELOC ARM THUMB IMM BFD RELOC ARM THUMB SHIFT BFD RELOC ARM THUMB OFFSET BFD RELOC ARM GOT12 BFD RELOC ARM GOT32 BFD RELOC ARM JUMP SLOT BFD RELOC ARM COPY BFD RELOC ARM GLOB DAT BFD RELOC ARM PLT32 BFD RELOC ARM RELATIVE BFD RELOC ARM GOTOFF BFD RELOC ARM GOTPC

These relocs are only used within the ARM assembler. They are not (at present) written to any object files.

BFD RELOC SH PCDISP8BY2 BFD RELOC SH PCDISP12BY2 BFD RELOC SH IMM4 BFD RELOC SH IMM4BY2 BFD RELOC SH IMM4BY4 BFD RELOC SH IMM8 BFD RELOC SH IMM8BY2 BFD RELOC SH IMM8BY4 BFD RELOC SH PCRELIMM8BY2 BFD RELOC SH PCRELIMM8BY4 BFD RELOC SH SWITCH16 BFD RELOC SH SWITCH32 BFD RELOC SH USES BFD RELOC SH COUNT BFD RELOC SH ALIGN BFD RELOC SH CODE BFD RELOC SH DATA BFD RELOC SH LABEL BFD RELOC SH LOOP START BFD RELOC SH LOOP END

Hitachi SH relocs. Not all of these appear in object files.

BFD RELOC THUMB PCREL BRANCH9 BFD RELOC THUMB PCREL BRANCH12 BFD RELOC THUMB PCREL BRANCH23

Thumb 23-, 12- and 9-bit pc-relative branches. The lowest bit must be zero and is not stored in the instruction.

BFD RELOC ARC B22 PCREL

Argonaut RISC Core (ARC) relocs. ARC 22 bit pc-relative branch. The lowest two bits must be zero and are not stored in the instruction. The high 20 bits are installed in bits 26 through 7 of the instruction.

BFD RELOC ARC B26

ARC 26 bit absolute branch. The lowest two bits must be zero and are not stored in the instruction. The high 24 bits are installed in bits 23 through 0.

BFD RELOC D10V 10 PCREL R

Mitsubishi D10V relocs. This is a 10-bit reloc with the right 2 bits assumed to be 0.

BFD RELOC D10V 10 PCREL L

Mitsubishi D10V relocs. This is a 10-bit reloc with the right 2 bits assumed to be 0. This is the same as the previous reloc except it is in the left container, i.e., shifted left 15 bits.

BFD RELOC D10V 18

This is an 18-bit reloc with the right 2 bits assumed to be 0.

BFD RELOC D10V 18 PCREL

This is an 18-bit reloc with the right 2 bits assumed to be 0.

BFD RELOC D30V 6

Mitsubishi D30V relocs. This is a 6-bit absolute reloc.

BFD RELOC D30V 9 PCREL

This is a 6-bit pc-relative reloc with the right 3 bits assumed to be 0.

BFD RELOC D30V 9 PCREL R

This is a 6-bit pc-relative reloc with the right 3 bits assumed to be 0. Same as the previous reloc but on the right side of the container.

BFD RELOC D30V 15

This is a 12-bit absolute reloc with the right 3 bitsassumed to be 0.

BFD RELOC D30V 15 PCREL

This is a 12-bit pc-relative reloc with the right 3 bits assumed to be 0.

BFD RELOC D30V 15 PCREL R

This is a 12-bit pc-relative reloc with the right 3 bits assumed to be 0. Same as the previous reloc but on the right side of the container.

BFD RELOC D30V 21

This is an 18-bit absolute reloc with the right 3 bits assumed to be 0.

BFD RELOC D30V 21 PCREL

This is an 18-bit pc-relative reloc with the right 3 bits assumed to be 0.

BFD RELOC D30V 21 PCREL R

This is an 18-bit pc-relative reloc with the right 3 bits assumed to be 0. Same as the previous reloc but on the right side of the container.

BFD RELOC D30V 32

This is a 32-bit absolute reloc.

BFD RELOC D30V 32 PCREL

This is a 32-bit pc-relative reloc.

BFD RELOC M32R 24

Mitsubishi M32R relocs. This is a 24 bit absolute address.

BFD RELOC M32R 10 PCREL

This is a 10-bit pc-relative reloc with the right 2 bits assumed to be 0.

BFD RELOC M32R 18 PCREL

This is an 18-bit reloc with the right 2 bits assumed to be 0.

BFD RELOC M32R 26 PCREL

This is a 26-bit reloc with the right 2 bits assumed to be 0.

BFD RELOC M32R HI16 ULO

This is a 16-bit reloc containing the high 16 bits of an address used when the lower 16 bits are treated as unsigned.

BFD RELOC M32R HI16 SLO

This is a 16-bit reloc containing the high 16 bits of an address used when the lower 16 bits are treated as signed.

BFD RELOC M32R LO16

This is a 16-bit reloc containing the lower 16 bits of an address.

BFD RELOC M32R SDA16

This is a 16-bit reloc containing the small data area offset for use in add3, load, and store instructions.

BFD RELOC V850 9 PCREL

This is a 9-bit reloc

BFD RELOC V850 22 PCREL

This is a 22-bit reloc

BFD RELOC V850 SDA 16 16 OFFSET

This is a 16 bit offset from the short data area pointer.

BFD RELOC V850 SDA 15 16 OFFSET

This is a 16 bit offset (of which only 15 bits are used) from the short data area pointer.

BFD RELOC V850 ZDA 16 16 OFFSET

This is a 16 bit offset from the zero data area pointer.

BFD RELOC V850 ZDA 15 16 OFFSET

This is a 16 bit offset (of which only 15 bits are used) from the zero data area pointer.

BFD RELOC V850 TDA 6 8 OFFSET

This is an 8 bit offset (of which only 6 bits are used) from the tiny data area pointer.

BFD RELOC V850 TDA 7 8 OFFSET

This is an 8bit offset (of which only 7 bits are used) from the tiny data area pointer.

BFD RELOC V850 TDA 7 7 OFFSET

This is a 7 bit offset from the tiny data area pointer.

BFD RELOC V850 TDA 16 16 OFFSET

This is a 16 bit offset from the tiny data area pointer.

BFD RELOC V850 TDA 4 5 OFFSET

This is a 5 bit offset (of which only 4 bits are used) from the tiny data area pointer.

BFD RELOC V850 TDA 4 4 OFFSET

This is a 4 bit offset from the tiny data area pointer.

BFD RELOC V850 SDA 16 16 SPLIT OFFSET

This is a 16 bit offset from the short data area pointer, with the bits placed noncontigously in the instruction.

BFD RELOC V850 ZDA 16 16 SPLIT OFFSET

This is a 16 bit offset from the zero data area pointer, with the bits placed noncontigously in the instruction.

BFD RELOC V850 CALLT 6 7 OFFSET

This is a 6 bit offset from the call table base pointer.

BFD RELOC V850 CALLT 16 16 OFFSET

This is a 16 bit offset from the call table base pointer.

BFD RELOC MN10300 32 PCREL

This is a 32bit pcrel reloc for the mn10300, offset by two bytes in the instruction.

BFD RELOC MN10300 16 PCREL

This is a 16bit pcrel reloc for the mn10300, offset by two bytes in the instruction.

BFD RELOC TIC30 LDP

This is a 8bit DP reloc for the tms320c30, where the most significant 8 bits of a 24 bit word are placed into the least significant 8 bits of the opcode.

BFD RELOC TIC54X PARTLS7

This is a 7bit reloc for the tms320c54x, where the least significant 7 bits of a 16 bit word are placed into the least significant 7 bits of the opcode.

BFD RELOC TIC54X PARTMS9

This is a 9bit DP reloc for the tms320c54x, where the most significant 9 bits of a 16 bit word are placed into the least significant 9 bits of the opcode.

BFD RELOC TIC54X 23

This is an extended address 23-bit reloc for the tms320c54x.

BFD RELOC TIC54X 16 OF 23

This is a 16-bit reloc for the tms320c54x, where the least significant 16 bits of a 23-bit extended address are placed into the opcode.

BFD RELOC TIC54X MS7 OF 23

This is a reloc for the tms320c54x, where the most significant 7 bits of a 23-bit extended address are placed into the opcode.

BFD RELOC FR30 48

This is a 48 bit reloc for the FR30 that stores 32 bits.

BFD RELOC FR30 20

This is a 32 bit reloc for the FR30 that stores 20 bits split up into two sections.

BFD RELOC FR30 6 IN 4

This is a 16 bit reloc for the FR30 that stores a 6 bit word offset in 4 bits.

BFD RELOC FR30 8 IN 8

This is a 16 bit reloc for the FR30 that stores an 8 bit byte offset into 8 bits.

BFD RELOC FR30 9 IN 8

This is a 16 bit reloc for the FR30 that stores a 9 bit short offset into 8 bits.

BFD RELOC FR30 10 IN 8

This is a 16 bit reloc for the FR30 that stores a 10 bit word offset into 8 bits.

BFD RELOC FR30 9 PCREL

This is a 16 bit reloc for the FR30 that stores a 9 bit pc relative short offset into 8 bits.

BFD RELOC FR30 12 PCREL

This is a 16 bit reloc for the FR30 that stores a 12 bit pc relative short offset into 11 bits.

BFD RELOC MCORE PCREL IMM8BY4 BFD RELOC MCORE PCREL IMM11BY2 BFD RELOC MCORE PCREL IMM4BY2 BFD RELOC MCORE PCREL 32 BFD RELOC MCORE PCREL JSR IMM11BY2 BFD RELOC MCORE RVA

Motorola Mcore relocations.

BFD RELOC AVR 7 PCREL

This is a 16 bit reloc for the AVR that stores 8 bit pc relative short offset into 7 bits.

BFD RELOC AVR 13 PCREL

This is a 16 bit reloc for the AVR that stores 13 bit pc relative short offset into 12 bits.

BFD RELOC AVR 16 PM

This is a 16 bit reloc for the AVR that stores 17 bit value (usually program memory address) into 16 bits.

BFD RELOC AVR LO8 LDI

This is a 16 bit reloc for the AVR that stores 8 bit value (usually data memory address) into 8 bit immediate value of LDI insn.

BFD RELOC AVR HI8 LDI

This is a 16 bit reloc for the AVR that stores 8 bit value (high 8 bit of data memory address) into 8 bit immediate value of LDI insn.

BFD RELOC AVR HH8 LDI

This is a 16 bit reloc for the AVR that stores 8 bit value (most high 8 bit of program memory address) into 8 bit immediate value of LDI insn.

BFD RELOC AVR LO8 LDI NEG

This is a 16 bit reloc for the AVR that stores negated 8 bit value (usually data memory address) into 8 bit immediate value of SUBI insn.

BFD RELOC AVR HI8 LDI NEG

This is a 16 bit reloc for the AVR that stores negated 8 bit value (high 8 bit of data memory address) into 8 bit immediate value of SUBI insn.

BFD RELOC AVR HH8 LDI NEG

This is a 16 bit reloc for the AVR that stores negated 8 bit value (most high 8 bit of program memory address) into 8 bit immediate value of LDI or SUBI insn.

BFD RELOC AVR LO8 LDI PM

This is a 16 bit reloc for the AVR that stores 8 bit value (usually command address) into 8 bit immediate value of LDI insn.

BFD RELOC AVR HI8 LDI PM

This is a 16 bit reloc for the AVR that stores 8 bit value (high 8 bit of command address) into 8 bit immediate value of LDI insn.

BFD RELOC AVR HH8 LDI PM

This is a 16 bit reloc for the AVR that stores 8 bit value (most high 8 bit of command address) into 8 bit immediate value of LDI insn.

BFD RELOC AVR LO8 LDI PM NEG

This is a 16 bit reloc for the AVR that stores negated 8 bit value (usually command address) into 8 bit immediate value of SUBI insn.

BFD RELOC AVR HI8 LDI PM NEG

This is a 16 bit reloc for the AVR that stores negated 8 bit value (high 8 bit of 16 bit command address) into 8 bit immediate value of SUBI insn.

BFD RELOC AVR HH8 LDI PM NEG

This is a 16 bit reloc for the AVR that stores negated 8 bit value (high 6 bit of 22 bit command address) into 8 bit immediate value of SUBI insn.

BFD RELOC AVR CALL

This is a 32 bit reloc for the AVR that stores 23 bit value into 22 bits.

BFD RELOC VTABLE INHERIT BFD RELOC VTABLE ENTRY

These two relocations are used by the linker to determine which of the entries in a C++ virtual function table are actually used. When the –gc-sections option is given, the linker will zero out the entries that are not used, so that the code for those functions need not be included in the output.

VTABLE INHERIT is a zero-space relocation used to describe to the linker the inheritence tree of a C++ virtual function table. The relocation’s symbol should be the parent class’ vtable, and the relocation should be located at the child vtable.

VTABLE ENTRY is a zero-space relocation that describes the use of a virtual function table entry. The reloc’s symbol should refer to the table of the class mentioned in the code. Off of that base, an offset describes the entry that is being used. For Rela hosts, this offset is stored in the reloc’s addend. For Rel hosts, we are forced to put this offset in the reloc’s section offset.

typedef enum bfd_reloc_code_real bfd_reloc_code_real_type;
2.11.0.2 bfd_reloc_type_lookup

Synopsis

reloc_howto_type *
bfd_reloc_type_lookup (bfd *abfd, bfd_reloc_code_real_type code);

Description Return a pointer to a howto structure which, when invoked, will perform the relocation code on data from the architecture noted.

2.11.0.3 bfd_default_reloc_type_lookup

Synopsis

reloc_howto_type *bfd_default_reloc_type_lookup
   (bfd *abfd, bfd_reloc_code_real_type  code);

Description Provides a default relocation lookup routine for any architecture.

2.11.0.4 bfd_get_reloc_code_name

Synopsis

     const char *bfd_get_reloc_code_name (bfd_reloc_code_real_type code);

Description Provides a printable name for the supplied relocation code. Useful mainly for printing error messages.

2.11.0.5 bfd_generic_relax_section

Synopsis

     boolean bfd_generic_relax_section
        (bfd *abfd,
         asection *section,
         struct bfd_link_info *,
         boolean *);

Description Provides default handling for relaxing for back ends which don’t do relaxing – i.e., does nothing.

2.11.0.6 bfd_generic_gc_sections

Synopsis

     boolean bfd_generic_gc_sections
        (bfd *, struct bfd_link_info *);

Description Provides default handling for relaxing for back ends which don’t do section gc – i.e., does nothing.

2.11.0.7 bfd_generic_get_relocated_section_contents

Synopsis

     bfd_byte *
     bfd_generic_get_relocated_section_contents (bfd *abfd,
         struct bfd_link_info *link_info,
         struct bfd_link_order *link_order,
         bfd_byte *data,
         boolean relocateable,
         asymbol **symbols);

Description Provides default handling of relocation effort for back ends which can’t be bothered to do it efficiently.

2.12 Core files

Description These are functions pertaining to core files.

2.12.0.1 bfd_core_file_failing_command

Synopsis

     CONST char *bfd_core_file_failing_command(bfd *abfd);

Description Return a read-only string explaining which program was running when it failed and produced the core file abfd.

2.12.0.2 bfd_core_file_failing_signal

Synopsis

     int bfd_core_file_failing_signal(bfd *abfd);

Description Returns the signal number which caused the core dump which generated the file the BFD abfd is attached to.

2.12.0.3 core_file_matches_executable_p

Synopsis

     boolean core_file_matches_executable_p
        (bfd *core_bfd, bfd *exec_bfd);

Description Return true if the core file attached to core bfd was generated by a run of the executable file attached to exec bfd, false otherwise.

2.13 Targets

Description Each port of BFD to a different machine requries the creation of a target back end. All the back end provides to the root part of BFD is a structure containing pointers to functions which perform certain low level operations on files. BFD translates the applications’s requests through a pointer into calls to the back end routines. When a file is opened with bfd_openr, its format and target are unknown. BFD uses various mechanisms to determine how to interpret the file. The operations performed are:

• Create a BFD by calling the internal routine _bfd_new_bfd, then call bfd_find_ target with the target string supplied to bfd_openr and the new BFD pointer.

• If a null target string was provided to bfd_find_target, look up the environment variable GNUTARGET and use that as the target string.

• If the target string is still NULL, or the target string is default, then use the first item in the target vector as the target type, and set target_defaulted in the BFD to cause bfd_check_format to loop through all the targets. See Section 2.13.1 [bfd target], page 63. See Section 2.9 [Formats], page 35.

• Otherwise, inspect the elements in the target vector one by one, until a match on target name is found. When found, use it.

  • Otherwise return the error bfd_error_invalid_target to bfd_openr.

• bfd_openr attempts to open the file using bfd_open_file, and returns the BFD.

Once the BFD has been opened and the target selected, the file format may be determined. This is done by calling bfd_check_format on the BFD with a suggested format. If target_ defaulted has been set, each possible target type is tried to see if it recognizes the specified format. bfd_check_format returns true when the caller guesses right.

2.13.1 bfd target

Description This structure contains everything that BFD knows about a target. It includes things like its byte order, name, and which routines to call to do various operations.

Every BFD points to a target structure with its xvec member.

The macros below are used to dispatch to functions through the bfd_target vector. They are used in a number of macros further down in ‘bfd.h’, and are also used when calling various routines by hand inside the BFD implementation. The arglist argument must be parenthesized; it contains all the arguments to the called function.

They make the documentation (more) unpleasant to read, so if someone wants to fix this and not break the above, please do.

     #define BFD_SEND(bfd, message, arglist) \
                    ((*((bfd)->xvec->message)) arglist)
     #ifdef DEBUG_BFD_SEND
     #undef BFD_SEND
     #define BFD_SEND(bfd, message, arglist) \
       (((bfd) && (bfd)->xvec && (bfd)->xvec->message) ? \
         ((*((bfd)->xvec->message)) arglist) : \
         (bfd_assert (__FILE__,__LINE__), NULL))
     #endif

For operations which index on the BFD format:

     #define BFD_SEND_FMT(bfd, message, arglist) \
                 (((bfd)->xvec->message[(int)((bfd)->format)]) arglist)
     #ifdef DEBUG_BFD_SEND
     #undef BFD_SEND_FMT
     #define BFD_SEND_FMT(bfd, message, arglist) \
       (((bfd) && (bfd)->xvec && (bfd)->xvec->message) ? \
        (((bfd)->xvec->message[(int)((bfd)->format)]) arglist) : \
        (bfd_assert (__FILE__,__LINE__), NULL))
     #endif

This is the structure which defines the type of BFD this is. The xvec member of the struct bfd itself points here. Each module that implements access to a different target under BFD, defines one of these.

FIXME, these names should be rationalised with the names of the entry points which call them. Too bad we can’t have one macro to define them both!

     enum bfd_flavour {
       bfd_target_unknown_flavour,
       bfd_target_aout_flavour,
       bfd_target_coff_flavour,

bfd_target_ecoff_flavour,
bfd_target_elf_flavour,
bfd_target_ieee_flavour,
bfd_target_nlm_flavour,
bfd_target_oasys_flavour,
bfd_target_tekhex_flavour,
bfd_target_srec_flavour,
bfd_target_ihex_flavour,
bfd_target_som_flavour,
bfd_target_os9k_flavour,
bfd_target_versados_flavour,
bfd_target_msdos_flavour,
bfd_target_ovax_flavour,
bfd_target_evax_flavour
};
enum bfd_endian { BFD_ENDIAN_BIG, BFD_ENDIAN_LITTLE, BFD_ENDIAN_UNKNOWN };
/* Forward declaration.  */
typedef struct bfd_link_info _bfd_link_info;
typedef struct bfd_target
{

Identifies the kind of target, e.g., SunOS4, Ultrix, etc.

char *name;

The "flavour" of a back end is a general indication about the contents of a file.

enum bfd_flavour flavour;

The order of bytes within the data area of a file.

enum bfd_endian byteorder;

The order of bytes within the header parts of a file.

enum bfd_endian header_byteorder;

A mask of all the flags which an executable may have set - from the set BFD_NO_FLAGS,

HAS_RELOC, ...D_PAGED.
flagword object_flags;

A mask of all the flags which a section may have set - from the set SEC_NO_FLAGS, SEC_

ALLOC, ...SET_NEVER_LOAD.
flagword section_flags;

The character normally found at the front of a symbol (if any), perhaps ‘ ’.

char symbol_leading_char;

The pad character for file names within an archive header.

char ar_pad_char;

The maximum number of characters in an archive header.

unsigned short ar_max_namelen;

Entries for byte swapping for data. These are different from the other entry points, since they don’t take a BFD asthe first argument. Certain other handlers could do the same.

bfd_vma      (*bfd_getx64) PARAMS ((const bfd_byte *));
bfd_signed_vma (*bfd_getx_signed_64) PARAMS ((const bfd_byte *));
void         (*bfd_putx64) PARAMS ((bfd_vma, bfd_byte *));

bfd_vma      (*bfd_getx32) PARAMS ((const bfd_byte *));
bfd_signed_vma (*bfd_getx_signed_32) PARAMS ((const bfd_byte *));
void         (*bfd_putx32) PARAMS ((bfd_vma, bfd_byte *));
bfd_vma      (*bfd_getx16) PARAMS ((const bfd_byte *));
bfd_signed_vma (*bfd_getx_signed_16) PARAMS ((const bfd_byte *));
void         (*bfd_putx16) PARAMS ((bfd_vma, bfd_byte *));

Byte swapping for the headers

bfd_vma      (*bfd_h_getx64) PARAMS ((const bfd_byte *));
bfd_signed_vma (*bfd_h_getx_signed_64) PARAMS ((const bfd_byte *));
void         (*bfd_h_putx64) PARAMS ((bfd_vma, bfd_byte *));
bfd_vma      (*bfd_h_getx32) PARAMS ((const bfd_byte *));
bfd_signed_vma (*bfd_h_getx_signed_32) PARAMS ((const bfd_byte *));
void         (*bfd_h_putx32) PARAMS ((bfd_vma, bfd_byte *));
bfd_vma      (*bfd_h_getx16) PARAMS ((const bfd_byte *));
bfd_signed_vma (*bfd_h_getx_signed_16) PARAMS ((const bfd_byte *));
void         (*bfd_h_putx16) PARAMS ((bfd_vma, bfd_byte *));

Format dependent routines: these are vectors of entry points within the target vector structure, one for each format to check.

Check the format of a file being read. Return a bfd_target * or zero.

const struct bfd_target *(*_bfd_check_format[bfd_type_end]) PARAMS ((bfd *));

Set the format of a file being written.

boolean             (*_bfd_set_format[bfd_type_end]) PARAMS ((bfd *));

Write cached information into a file being written, at bfd_close.

boolean             (*_bfd_write_contents[bfd_type_end]) PARAMS ((bfd *));

The general target vector. These vectors are initialized using the BFD JUMP TABLE macros.

 /* Generic entry points.  */
#define BFD_JUMP_TABLE_GENERIC(NAME)\
CAT(NAME,_close_and_cleanup),\
CAT(NAME,_bfd_free_cached_info),\
CAT(NAME,_new_section_hook),\
CAT(NAME,_get_section_contents),\
CAT(NAME,_get_section_contents_in_window)
 /* Called when the BFD is being closed to do any necessary cleanup.  */
boolean       (*_close_and_cleanup) PARAMS ((bfd *));
 /* Ask the BFD to free all cached information.  */
boolean (*_bfd_free_cached_info) PARAMS ((bfd *));
 /* Called when a new section is created.  */
boolean       (*_new_section_hook) PARAMS ((bfd *, sec_ptr));
 /* Read the contents of a section.  */
boolean       (*_bfd_get_section_contents) PARAMS ((bfd *, sec_ptr, PTR,
                                          file_ptr, bfd_size_type));
boolean       (*_bfd_get_section_contents_in_window)
                        PARAMS ((bfd *, sec_ptr, bfd_window *,
                                 file_ptr, bfd_size_type));

   /* Entry points to copy private data.  */
#define BFD_JUMP_TABLE_COPY(NAME)\
CAT(NAME,_bfd_copy_private_bfd_data),\
CAT(NAME,_bfd_merge_private_bfd_data),\
CAT(NAME,_bfd_copy_private_section_data),\
CAT(NAME,_bfd_copy_private_symbol_data),\
CAT(NAME,_bfd_set_private_flags),\
CAT(NAME,_bfd_print_private_bfd_data)\
   /* Called to copy BFD general private data from one object file
     to another.  */
  boolean       (*_bfd_copy_private_bfd_data) PARAMS ((bfd *, bfd *));
   /* Called to merge BFD general private data from one object file
     to a common output file when linking.  */
  boolean       (*_bfd_merge_private_bfd_data) PARAMS ((bfd *, bfd *));
   /* Called to copy BFD private section data from one object file
     to another.  */
  boolean       (*_bfd_copy_private_section_data) PARAMS ((bfd *, sec_ptr,
                                                       bfd *, sec_ptr));
   /* Called to copy BFD private symbol data from one symbol
     to another.  */
  boolean       (*_bfd_copy_private_symbol_data) PARAMS ((bfd *, asymbol *,
                                                          bfd *, asymbol *));
   /* Called to set private backend flags */
  boolean       (*_bfd_set_private_flags) PARAMS ((bfd *, flagword));
   /* Called to print private BFD data */
  boolean       (*_bfd_print_private_bfd_data) PARAMS ((bfd *, PTR));
   /* Core file entry points.  */
#define BFD_JUMP_TABLE_CORE(NAME)\
CAT(NAME,_core_file_failing_command),\
CAT(NAME,_core_file_failing_signal),\
CAT(NAME,_core_file_matches_executable_p)
  char *   (*_core_file_failing_command) PARAMS ((bfd *));
  int      (*_core_file_failing_signal) PARAMS ((bfd *));
  boolean  (*_core_file_matches_executable_p) PARAMS ((bfd *, bfd *));
   /* Archive entry points.  */
#define BFD_JUMP_TABLE_ARCHIVE(NAME)\
CAT(NAME,_slurp_armap),\
CAT(NAME,_slurp_extended_name_table),\
CAT(NAME,_construct_extended_name_table),\
CAT(NAME,_truncate_arname),\
CAT(NAME,_write_armap),\
CAT(NAME,_read_ar_hdr),\
CAT(NAME,_openr_next_archived_file),\
CAT(NAME,_get_elt_at_index),\
CAT(NAME,_generic_stat_arch_elt),\
CAT(NAME,_update_armap_timestamp)

boolean  (*_bfd_slurp_armap) PARAMS ((bfd *));
boolean  (*_bfd_slurp_extended_name_table) PARAMS ((bfd *));
boolean  (*_bfd_construct_extended_name_table)
           PARAMS ((bfd *, char **, bfd_size_type *, const char **));
void     (*_bfd_truncate_arname) PARAMS ((bfd *, CONST char *, char *));
boolean  (*write_armap) PARAMS ((bfd *arch,
                            unsigned int elength,
                            struct orl *map,
                            unsigned int orl_count,
                            int stridx));
PTR (*_bfd_read_ar_hdr_fn) PARAMS ((bfd *));
bfd *    (*openr_next_archived_file) PARAMS ((bfd *arch, bfd *prev));
#define bfd_get_elt_at_index(b,i) BFD_SEND(b, _bfd_get_elt_at_index, (b,i))
bfd *    (*_bfd_get_elt_at_index) PARAMS ((bfd *, symindex));
int      (*_bfd_stat_arch_elt) PARAMS ((bfd *, struct stat *));
boolean  (*_bfd_update_armap_timestamp) PARAMS ((bfd *));
 /* Entry points used for symbols.  */
#define BFD_JUMP_TABLE_SYMBOLS(NAME)\
CAT(NAME,_get_symtab_upper_bound),\
CAT(NAME,_get_symtab),\
CAT(NAME,_make_empty_symbol),\
CAT(NAME,_print_symbol),\
CAT(NAME,_get_symbol_info),\
CAT(NAME,_bfd_is_local_label_name),\
CAT(NAME,_get_lineno),\
CAT(NAME,_find_nearest_line),\
CAT(NAME,_bfd_make_debug_symbol),\
CAT(NAME,_read_minisymbols),\
CAT(NAME,_minisymbol_to_symbol)
long  (*_bfd_get_symtab_upper_bound) PARAMS ((bfd *));
long  (*_bfd_canonicalize_symtab) PARAMS ((bfd *,
                                           struct symbol_cache_entry **));
struct symbol_cache_entry  *
              (*_bfd_make_empty_symbol) PARAMS ((bfd *));
void          (*_bfd_print_symbol) PARAMS ((bfd *, PTR,
                                    struct symbol_cache_entry *,
                                    bfd_print_symbol_type));
#define bfd_print_symbol(b,p,s,e) BFD_SEND(b, _bfd_print_symbol, (b,p,s,e))
void          (*_bfd_get_symbol_info) PARAMS ((bfd *,
                                    struct symbol_cache_entry *,
                                    symbol_info *));
#define bfd_get_symbol_info(b,p,e) BFD_SEND(b, _bfd_get_symbol_info, (b,p,e))
boolean       (*_bfd_is_local_label_name) PARAMS ((bfd *, const char *));
alent *    (*_get_lineno) PARAMS ((bfd *, struct symbol_cache_entry *));
boolean    (*_bfd_find_nearest_line) PARAMS ((bfd *abfd,
                  struct sec *section, struct symbol_cache_entry **symbols,
                  bfd_vma offset, CONST char **file, CONST char **func,

                    unsigned int *line));
  /* Back-door to allow format-aware applications to create debug symbols
    while using BFD for everything else.  Currently used by the assembler
    when creating COFF files.  */
  asymbol *  (*_bfd_make_debug_symbol) PARAMS ((
       bfd *abfd,
       void *ptr,
       unsigned long size));
#define bfd_read_minisymbols(b, d, m, s) \
  BFD_SEND (b, _read_minisymbols, (b, d, m, s))
  long  (*_read_minisymbols) PARAMS ((bfd *, boolean, PTR *,
                                      unsigned int *));
#define bfd_minisymbol_to_symbol(b, d, m, f) \
  BFD_SEND (b, _minisymbol_to_symbol, (b, d, m, f))
  asymbol *(*_minisymbol_to_symbol) PARAMS ((bfd *, boolean, const PTR,
                                             asymbol *));
   /* Routines for relocs.  */
#define BFD_JUMP_TABLE_RELOCS(NAME)\
CAT(NAME,_get_reloc_upper_bound),\
CAT(NAME,_canonicalize_reloc),\
CAT(NAME,_bfd_reloc_type_lookup)
  long  (*_get_reloc_upper_bound) PARAMS ((bfd *, sec_ptr));
  long  (*_bfd_canonicalize_reloc) PARAMS ((bfd *, sec_ptr, arelent **,
                                            struct symbol_cache_entry **));
   /* See documentation on reloc types.  */
  reloc_howto_type *
       (*reloc_type_lookup) PARAMS ((bfd *abfd,
                                     bfd_reloc_code_real_type code));
   /* Routines used when writing an object file.  */
#define BFD_JUMP_TABLE_WRITE(NAME)\
CAT(NAME,_set_arch_mach),\
CAT(NAME,_set_section_contents)
  boolean    (*_bfd_set_arch_mach) PARAMS ((bfd *, enum bfd_architecture,
                    unsigned long));
  boolean       (*_bfd_set_section_contents) PARAMS ((bfd *, sec_ptr, PTR,
                                            file_ptr, bfd_size_type));
   /* Routines used by the linker.  */
#define BFD_JUMP_TABLE_LINK(NAME)\
CAT(NAME,_sizeof_headers),\
CAT(NAME,_bfd_get_relocated_section_contents),\
CAT(NAME,_bfd_relax_section),\
CAT(NAME,_bfd_link_hash_table_create),\
CAT(NAME,_bfd_link_add_symbols),\
CAT(NAME,_bfd_final_link),\
CAT(NAME,_bfd_link_split_section),\
CAT(NAME,_bfd_gc_sections)

int        (*_bfd_sizeof_headers) PARAMS ((bfd *, boolean));
bfd_byte * (*_bfd_get_relocated_section_contents) PARAMS ((bfd *,
                  struct bfd_link_info *, struct bfd_link_order *,
                  bfd_byte *data, boolean relocateable,
                  struct symbol_cache_entry **));
boolean    (*_bfd_relax_section) PARAMS ((bfd *, struct sec *,
                  struct bfd_link_info *, boolean *again));
 /* Create a hash table for the linker.  Different backends store
   different information in this table.  */
struct bfd_link_hash_table *(*_bfd_link_hash_table_create) PARAMS ((bfd *));
 /* Add symbols from this object file into the hash table.  */
boolean (*_bfd_link_add_symbols) PARAMS ((bfd *, struct bfd_link_info *));
 /* Do a link based on the link_order structures attached to each
   section of the BFD.  */
boolean (*_bfd_final_link) PARAMS ((bfd *, struct bfd_link_info *));
 /* Should this section be split up into smaller pieces during linking.  */
boolean (*_bfd_link_split_section) PARAMS ((bfd *, struct sec *));
 /* Remove sections that are not referenced from the output.  */
boolean (*_bfd_gc_sections) PARAMS ((bfd *, struct bfd_link_info *));
 /* Routines to handle dynamic symbols and relocs.  */
#define BFD_JUMP_TABLE_DYNAMIC(NAME)\
CAT(NAME,_get_dynamic_symtab_upper_bound),\
CAT(NAME,_canonicalize_dynamic_symtab),\
CAT(NAME,_get_dynamic_reloc_upper_bound),\
CAT(NAME,_canonicalize_dynamic_reloc)
 /* Get the amount of memory required to hold the dynamic symbols. */
long  (*_bfd_get_dynamic_symtab_upper_bound) PARAMS ((bfd *));
 /* Read in the dynamic symbols.  */
long  (*_bfd_canonicalize_dynamic_symtab)
  PARAMS ((bfd *, struct symbol_cache_entry **));
 /* Get the amount of memory required to hold the dynamic relocs.  */
long  (*_bfd_get_dynamic_reloc_upper_bound) PARAMS ((bfd *));
 /* Read in the dynamic relocs.  */
long  (*_bfd_canonicalize_dynamic_reloc)
  PARAMS ((bfd *, arelent **, struct symbol_cache_entry **));

A pointer to an alternative bfd target in case the current one is not satisfactory. This can happen when the target cpu supports both big and little endian code, and target chosen by the linker has the wrong endianness. The function open output() in ld/ldlang.c uses this field to find an alternative output format that is suitable.

/* Opposite endian version of this target.  */
const struct bfd_target * alternative_target;

Data for use by back-end routines, which isn’t generic enough to belong in this structure.

      PTR backend_data;
     } bfd_target;
2.13.1.1 bfd_set_default_target

Synopsis

     boolean bfd_set_default_target (const char *name);

Description Set the default target vector to use when recognizing a BFD. This takes the name of the target, which may be a BFD target name or a configuration triplet.

2.13.1.2 bfd_find_target

Synopsis

     const bfd_target *bfd_find_target(CONST char *target_name, bfd *abfd);

Description Return a pointer to the transfer vector for the object target named target name. If target name is NULL, choose the one in the environment variable GNUTARGET; if that is null or not defined, then choose the first entry in the target list. Passing in the string "default" or setting the environment variable to "default" will cause the first entry in the target list to be returned, and "target defaulted" will be set in the BFD. This causes bfd_check_format to loop over all the targets to find the one that matches the file being read.

2.13.1.3 bfd_target_list

Synopsis

     const char **bfd_target_list(void);

Description Return a freshly malloced NULL-terminated vector of the names of all the valid BFD targets. Do not modify the names.

2.13.1.4 bfd_seach_for_target

Synopsis

     const bfd_target * bfd_search_for_target (int (* search_func)(const bfd_target *, void *), void *);

Description Return a pointer to the first transfer vector in the list of transfer vectors maintained by BFD that produces a non-zero result when passed to the function search func. The parameter data is passed, unexamined, to the search function.

2.14 Architectures

BFD keeps one atom in a BFD describing the architecture of the data attached to the BFD:

a pointer to a bfd_arch_info_type.

Pointers to structures can be requested independently of a BFD so that an architecture’s information can be interrogated without access to an open BFD.

The architecture information is provided by each architecture package. The set of default architectures is selected by the macro SELECT_ARCHITECTURES. This is normally set up in the ‘config/target.mt’ file of your choice. If the name is not defined, then all the architectures supported are included. When BFD starts up, all the architectures are called with an initialize method. It is up to the architecture back end to insert as many items into the list of architectures as it wants to; generally this would be one for each machine and one for the default case (an item with a machine field of 0). BFD’s idea of an architecture is implemented in ‘archures.c’.

2.14.1 bfd architecture

Description This enum gives the object file’s CPU architecture, in a global sense—i.e., what processor family does it belong to? Another field indicates which processor within the family is in use. The machine gives a number which distinguishes different versions of the architecture, containing, for example, 2 and 3 for Intel i960 KA and i960 KB, and 68020 and 68030 for Motorola 68020 and 68030.

enum bfd_architecture
{
  bfd_arch_unknown,    /* File arch not known */
  bfd_arch_obscure,    /* Arch known, not one of these */
  bfd_arch_m68k,       /* Motorola 68xxx */
#define bfd_mach_m68000 1
#define bfd_mach_m68008 2
#define bfd_mach_m68010 3
#define bfd_mach_m68020 4
#define bfd_mach_m68030 5
#define bfd_mach_m68040 6
#define bfd_mach_m68060 7
#define bfd_mach_cpu32  8
  bfd_arch_vax,        /* DEC Vax */
  bfd_arch_i960,       /* Intel 960 */
     /* The order of the following is important.
       lower number indicates a machine type that
       only accepts a subset of the instructions
       available to machines with higher numbers.
       The exception is the "ca", which is
       incompatible with all other machines except
       "core". */
#define bfd_mach_i960_core      1
#define bfd_mach_i960_ka_sa     2
#define bfd_mach_i960_kb_sb     3
#define bfd_mach_i960_mc        4
#define bfd_mach_i960_xa        5
#define bfd_mach_i960_ca        6
#define bfd_mach_i960_jx        7

#define bfd_mach_i960_hx        8
  bfd_arch_a29k,       /* AMD 29000 */
  bfd_arch_sparc,      /* SPARC */
#define bfd_mach_sparc                 1
 /* The difference between v8plus and v9 is that v9 is a true 64 bit env.  */
#define bfd_mach_sparc_sparclet        2
#define bfd_mach_sparc_sparclite       3
#define bfd_mach_sparc_v8plus          4
#define bfd_mach_sparc_v8plusa         5  /* with ultrasparc add'ns */
#define bfd_mach_sparc_sparclite_le    6
#define bfd_mach_sparc_v9              7
#define bfd_mach_sparc_v9a             8  /* with ultrasparc add'ns */
 /* Nonzero if MACH has the v9 instruction set.  */
#define bfd_mach_sparc_v9_p(mach) \
  ((mach) >= bfd_mach_sparc_v8plus && (mach) <= bfd_mach_sparc_v9a)
  bfd_arch_mips,       /* MIPS Rxxxx */
#define bfd_mach_mips3000              3000
#define bfd_mach_mips3900              3900
#define bfd_mach_mips4000              4000
#define bfd_mach_mips4010              4010
#define bfd_mach_mips4100              4100
#define bfd_mach_mips4111              4111
#define bfd_mach_mips4300              4300
#define bfd_mach_mips4400              4400
#define bfd_mach_mips4600              4600
#define bfd_mach_mips4650              4650
#define bfd_mach_mips5000              5000
#define bfd_mach_mips6000              6000
#define bfd_mach_mips8000              8000
#define bfd_mach_mips10000             10000
#define bfd_mach_mips16                16
  bfd_arch_i386,       /* Intel 386 */
#define bfd_mach_i386_i386 0
#define bfd_mach_i386_i8086 1
#define bfd_mach_i386_i386_intel_syntax 2
  bfd_arch_we32k,      /* AT&T WE32xxx */
  bfd_arch_tahoe,      /* CCI/Harris Tahoe */
  bfd_arch_i860,       /* Intel 860 */
  bfd_arch_i370,       /* IBM 360/370 Mainframes */
  bfd_arch_romp,       /* IBM ROMP PC/RT */
  bfd_arch_alliant,    /* Alliant */
  bfd_arch_convex,     /* Convex */
  bfd_arch_m88k,       /* Motorola 88xxx */
  bfd_arch_pyramid,    /* Pyramid Technology */
  bfd_arch_h8300,      /* Hitachi H8/300 */
#define bfd_mach_h8300   1
#define bfd_mach_h8300h  2
#define bfd_mach_h8300s  3

  bfd_arch_powerpc,    /* PowerPC */
  bfd_arch_rs6000,     /* IBM RS/6000 */
  bfd_arch_hppa,       /* HP PA RISC */
  bfd_arch_d10v,       /* Mitsubishi D10V */
#define bfd_mach_d10v          0
#define bfd_mach_d10v_ts2      2
#define bfd_mach_d10v_ts3      3
  bfd_arch_d30v,       /* Mitsubishi D30V */
  bfd_arch_z8k,        /* Zilog Z8000 */
#define bfd_mach_z8001         1
#define bfd_mach_z8002         2
  bfd_arch_h8500,      /* Hitachi H8/500 */
  bfd_arch_sh,         /* Hitachi SH */
#define bfd_mach_sh            0
#define bfd_mach_sh2        0x20
#define bfd_mach_sh_dsp     0x2d
#define bfd_mach_sh3        0x30
#define bfd_mach_sh3_dsp    0x3d
#define bfd_mach_sh3e       0x3e
#define bfd_mach_sh4        0x40
  bfd_arch_alpha,      /* Dec Alpha */
#define bfd_mach_alpha_ev4  0x10
#define bfd_mach_alpha_ev5  0x20
#define bfd_mach_alpha_ev6  0x30
  bfd_arch_arm,        /* Advanced Risc Machines ARM */
#define bfd_mach_arm_2         1
#define bfd_mach_arm_2a        2
#define bfd_mach_arm_3         3
#define bfd_mach_arm_3M        4
#define bfd_mach_arm_4         5
#define bfd_mach_arm_4T        6
#define bfd_mach_arm_5         7
#define bfd_mach_arm_5T        8
  bfd_arch_ns32k,      /* National Semiconductors ns32000 */
  bfd_arch_w65,        /* WDC 65816 */
  bfd_arch_tic30,      /* Texas Instruments TMS320C30 */
  bfd_arch_tic54x,     /* Texas Instruments TMS320C54X */
  bfd_arch_tic80,      /* TI TMS320c80 (MVP) */
  bfd_arch_v850,       /* NEC V850 */
#define bfd_mach_v850          0
#define bfd_mach_v850e         'E'
#define bfd_mach_v850ea        'A'
  bfd_arch_arc,        /* Argonaut RISC Core */
#define bfd_mach_arc_base 0
  bfd_arch_m32r,       /* Mitsubishi M32R/D */
#define bfd_mach_m32r          0  /* backwards compatibility */
#define bfd_mach_m32rx         'x'
  bfd_arch_mn10200,    /* Matsushita MN10200 */
  bfd_arch_mn10300,    /* Matsushita MN10300 */

#define bfd_mach_mn10300               300
#define bfd_mach_am33          330
bfd_arch_fr30,
#define bfd_mach_fr30          0x46523330
bfd_arch_mcore,
bfd_arch_pj,
bfd_arch_avr,        /* Atmel AVR microcontrollers */
#define bfd_mach_avr1          1
#define bfd_mach_avr2          2
#define bfd_mach_avr3          3
#define bfd_mach_avr4          4
bfd_arch_last
};

2.14.2 bfd arch info

Description This structure contains information on architectures for use within BFD.

typedef struct bfd_arch_info
{
int bits_per_word;
int bits_per_address;
int bits_per_byte;
enum bfd_architecture arch;
unsigned long mach;
const char *arch_name;
const char *printable_name;
unsigned int section_align_power;
/* true if this is the default machine for the architecture */
boolean the_default;
const struct bfd_arch_info * (*compatible)
     PARAMS ((const struct bfd_arch_info *a,
              const struct bfd_arch_info *b));
boolean (*scan) PARAMS ((const struct bfd_arch_info *, const char *));
const struct bfd_arch_info *next;
} bfd_arch_info_type;
2.14.2.1 bfd_printable_name

Synopsis

const char *bfd_printable_name(bfd *abfd);

Description Return a printable string representing the architecture and machine from the pointer to the architecture info structure.

2.14.2.2 bfd_scan_arch

Synopsis

     const bfd_arch_info_type *bfd_scan_arch(const char *string);

Description Figure out if BFD supports any cpu which could be described with the name string. Return a pointer to an arch_info structure if a machine is found, otherwise NULL.

2.14.2.3 bfd_arch_list

Synopsis

     const char **bfd_arch_list(void);

Description Return a freshly malloced NULL-terminated vector of the names of all the valid BFD architectures. Do not modify the names.

2.14.2.4 bfd_arch_get_compatible

Synopsis

     const bfd_arch_info_type *bfd_arch_get_compatible(
         const bfd *abfd,
         const bfd *bbfd);

Description Determine whether two BFDs’ architectures and machine types are compatible. Calculates the lowest common denominator between the two architectures and machine types implied by the BFDs and returns a pointer to an arch_info structure describing the compatible machine.

2.14.2.5 bfd_default_arch_struct

Description

The bfd_default_arch_struct is an item of bfd_arch_info_type which has been initial-

ized to a fairly generic state. A BFD starts life by pointing to this structure, until the correct back end has determined the real architecture of the file.

     extern const bfd_arch_info_type bfd_default_arch_struct;
2.14.2.6 bfd_set_arch_info

Synopsis

     void bfd_set_arch_info(bfd *abfd, const bfd_arch_info_type *arg);

Description Set the architecture info of abfd to arg.

2.14.2.7 bfd_default_set_arch_mach

Synopsis

     boolean bfd_default_set_arch_mach(bfd *abfd,
         enum bfd_architecture arch,
         unsigned long mach);

Description Set the architecture and machine type in BFD abfd to arch and mach. Find the correct pointer to a structure and insert it into the arch_info pointer.

2.14.2.8 bfd_get_arch

Synopsis

     enum bfd_architecture bfd_get_arch(bfd *abfd);

Description Return the enumerated type which describes the BFD abfd’s architecture.

2.14.2.9 bfd_get_mach

Synopsis

     unsigned long bfd_get_mach(bfd *abfd);

Description Return the long type which describes the BFD abfd’s machine.

2.14.2.10 bfd_arch_bits_per_byte

Synopsis

     unsigned int bfd_arch_bits_per_byte(bfd *abfd);

Description Return the number of bits in one of the BFD abfd’s architecture’s bytes.

2.14.2.11 bfd_arch_bits_per_address

Synopsis

     unsigned int bfd_arch_bits_per_address(bfd *abfd);

Description Return the number of bits in one of the BFD abfd’s architecture’s addresses.

2.14.2.12 bfd_default_compatible

Synopsis

     const bfd_arch_info_type *bfd_default_compatible
        (const bfd_arch_info_type *a,
         const bfd_arch_info_type *b);

Description The default function for testing for compatibility.

2.14.2.13 bfd_default_scan

Synopsis

     boolean bfd_default_scan(const struct bfd_arch_info *info, const char *string);

Description The default function for working out whether this is an architecture hit and a machine hit.

2.14.2.14 bfd_get_arch_info

Synopsis

     const bfd_arch_info_type * bfd_get_arch_info(bfd *abfd);

Description Return the architecture info struct in abfd.

2.14.2.15 bfd_lookup_arch

Synopsis

     const bfd_arch_info_type *bfd_lookup_arch
        (enum bfd_architecture
         arch,
         unsigned long machine);

Description Look for the architecure info structure which matches the arguments arch and machine. A machine of 0 matches the machine/architecture structure which marks itself as the default.

2.14.2.16 bfd_printable_arch_mach

Synopsis

     const char *bfd_printable_arch_mach
        (enum bfd_architecture arch, unsigned long machine);

Description Return a printable string representing the architecture and machine type.

This routine is depreciated.

2.14.2.17 bfd_octets_per_byte

Synopsis

     unsigned int bfd_octets_per_byte(bfd *abfd);

Description Return the number of octets (8-bit quantities) per target byte (minimum addressable unit). In most cases, this will be one, but some DSP targets have 16, 32, or even 48 bits per byte.

2.14.2.18 bfd_arch_mach_octets_per_byte

Synopsis

     unsigned int bfd_arch_mach_octets_per_byte(enum bfd_architecture arch,
         unsigned long machine);

Description See bfd octets per byte. This routine is provided for those cases where a bfd * is not available

2.15 Opening and closing BFDs

2.15.0.1 bfd_openr

Synopsis

     bfd *bfd_openr(CONST char *filename, CONST char *target);

Description Open the file filename (using fopen) with the target target. Return a pointer to the created BFD.

Calls bfd_find_target, so target is interpreted as by that function.

If NULL is returned then an error has occured. Possible errors are bfd_error_no_memory,

bfd_error_invalid_target or system_call error.
2.15.0.2 bfd_fdopenr

Synopsis

     bfd *bfd_fdopenr(CONST char *filename, CONST char *target, int fd);

Description bfd_fdopenr is to bfd_fopenr much like fdopen is to fopen. It opens a BFD on a file already described by the fd supplied.

When the file is later bfd_closed, the file descriptor will be closed.

If the caller desires that this file descriptor be cached by BFD (opened as needed, closed as needed to free descriptors for other opens), with the supplied fd used as an initial file descriptor (but subject to closure at any time), call bfd set cacheable(bfd, 1) on the returned BFD. The default is to assume no cacheing; the file descriptor will remain open until bfd_ close, and will not be affected by BFD operations on other files.

Possible errors are bfd_error_no_memory, bfd_error_invalid_target and bfd_error_
system_call.
2.15.0.3 bfd_openstreamr

Synopsis

     bfd *bfd_openstreamr(const char *, const char *, PTR);

Description Open a BFD for read access on an existing stdio stream. When the BFD is passed to bfd_close, the stream will be closed.

2.15.0.4 bfd_openw

Synopsis

     bfd *bfd_openw(CONST char *filename, CONST char *target);

Description Create a BFD, associated with file filename, using the file format target, and return a pointer to it.

Possible errors are bfd_error_system_call, bfd_error_no_memory, bfd_error_invalid_
target.

2.15.0.5 bfd_close

Synopsis

     boolean bfd_close(bfd *abfd);

Description Close a BFD. If the BFD was open for writing, then pending operations are completed and the file written out and closed. If the created file is executable, then chmod is called to mark it as such.

All memory attached to the BFD is released.

The file descriptor associated with the BFD is closed (even if it was passed in to BFD by

bfd_fdopenr).

Returns true is returned if all is ok, otherwise false.

2.15.0.6 bfd_close_all_done

Synopsis

     boolean bfd_close_all_done(bfd *);

Description Close a BFD. Differs from bfd_close since it does not complete any pending operations. This routine would be used if the application had just used BFD for swapping and didn’t want to use any of the writing code.

If the created file is executable, then chmod is called to mark it as such.

All memory attached to the BFD is released.

Returns true is returned if all is ok, otherwise false.

2.15.0.7 bfd_create

Synopsis

     bfd *bfd_create(CONST char *filename, bfd *templ);

Description Create a new BFD in the manner of bfd_openw, but without opening a file. The new BFD takes the target from the target used by template. The format is always set to bfd_object.

2.15.0.8 bfd_make_writable

Synopsis

     boolean bfd_make_writable(bfd *abfd);

Description Takes a BFD as created by bfd_create and converts it into one like as returned by bfd_ openw. It does this by converting the BFD to BFD IN MEMORY. It’s assumed that you will call bfd_make_readable on this bfd later.

Returns true is returned if all is ok, otherwise false.

2.15.0.9 bfd_make_readable

Synopsis

     boolean bfd_make_readable(bfd *abfd);

Description Takes a BFD as created by bfd_create and bfd_make_writable and converts it into one like as returned by bfd_openr. It does this by writing the contents out to the memory buffer, then reversing the direction. Returns true is returned if all is ok, otherwise false.

2.15.0.10 bfd_alloc

Synopsis

     PTR bfd_alloc (bfd *abfd, size_t wanted);

Description Allocate a block of wanted bytes of memory attached to abfd and return a pointer to it.

2.16 Internal functions

Description These routines are used within BFD. They are not intended for export, but are documented here for completeness.

2.16.0.1 bfd_write_bigendian_4byte_int

Synopsis

     void bfd_write_bigendian_4byte_int(bfd *abfd,  int i);

Description Write a 4 byte integer i to the output BFD abfd, in big endian order regardless of what else is going on. This is useful in archives.

2.16.0.2 bfd_put_size

2.16.0.3 bfd_get_size

Description These macros as used for reading and writing raw data in sections; each access (except for bytes) is vectored through the target format of the BFD and mangled accordingly. The mangling performs any necessary endian translations and removes alignment restrictions. Note that types accepted and returned by these macros are identical so they can be swapped around in macros—for example, ‘libaout.h’ defines GET_WORD to either bfd_get_32 or

bfd_get_64.

In the put routines, val must be a bfd_vma. If we are on a system without prototypes, the caller is responsible for making sure that is true, with a cast if necessary. We don’t cast them in the macro definitions because that would prevent lint or gcc -Wall from detecting sins such as passing a pointer. To detect calling these with less than a bfd_vma, use gcc

-Wconversion on a host with 64 bit bfd_vma’s.
/* Byte swapping macros for user section data.  */
#define bfd_put_8(abfd, val, ptr) \
((void) (*((unsigned char *)(ptr)) = (unsigned char)(val)))
#define bfd_put_signed_8 \
bfd_put_8
#define bfd_get_8(abfd, ptr) \
(*(unsigned char *)(ptr))
#define bfd_get_signed_8(abfd, ptr) \
((*(unsigned char *)(ptr) ^0x80) - 0x80)
#define bfd_put_16(abfd, val, ptr) \
BFD_SEND(abfd, bfd_putx16, ((val),(ptr)))
#define bfd_put_signed_16 \
bfd_put_16
#define bfd_get_16(abfd, ptr) \
BFD_SEND(abfd, bfd_getx16, (ptr))
#define bfd_get_signed_16(abfd, ptr) \
BFD_SEND (abfd, bfd_getx_signed_16, (ptr))
#define bfd_put_32(abfd, val, ptr) \
BFD_SEND(abfd, bfd_putx32, ((val),(ptr)))
#define bfd_put_signed_32 \
bfd_put_32
#define bfd_get_32(abfd, ptr) \
BFD_SEND(abfd, bfd_getx32, (ptr))
#define bfd_get_signed_32(abfd, ptr) \
BFD_SEND(abfd, bfd_getx_signed_32, (ptr))
#define bfd_put_64(abfd, val, ptr) \
BFD_SEND(abfd, bfd_putx64, ((val), (ptr)))
#define bfd_put_signed_64 \
bfd_put_64
#define bfd_get_64(abfd, ptr) \
BFD_SEND(abfd, bfd_getx64, (ptr))
#define bfd_get_signed_64(abfd, ptr) \
BFD_SEND(abfd, bfd_getx_signed_64, (ptr))
#define bfd_get(bits, abfd, ptr)                               \
((bits) == 8 ? bfd_get_8 (abfd, ptr)           \
: (bits) == 16 ? bfd_get_16 (abfd, ptr)        \
: (bits) == 32 ? bfd_get_32 (abfd, ptr)        \
: (bits) == 64 ? bfd_get_64 (abfd, ptr)        \
: (abort (), (bfd_vma) - 1))
#define bfd_put(bits, abfd, val, ptr)                          \

((bits) == 8 ? bfd_put_8 (abfd, val, ptr)      \
: (bits) == 16 ? bfd_put_16 (abfd, val, ptr)   \
: (bits) == 32 ? bfd_put_32 (abfd, val, ptr)   \
: (bits) == 64 ? bfd_put_64 (abfd, val, ptr)   \
: (abort (), (void) 0))

2.16.0.4 bfd_h_put_size

Description These macros have the same function as their bfd_get_x bretheren, except that they are used for removing information for the header records of object files. Believe it or not, some object files keep their header records in big endian order and their data in little endian order.

/* Byte swapping macros for file header data.  */
#define bfd_h_put_8(abfd, val, ptr) \
bfd_put_8 (abfd, val, ptr)
#define bfd_h_put_signed_8(abfd, val, ptr) \
bfd_put_8 (abfd, val, ptr)
#define bfd_h_get_8(abfd, ptr) \
bfd_get_8 (abfd, ptr)
#define bfd_h_get_signed_8(abfd, ptr) \
bfd_get_signed_8 (abfd, ptr)
#define bfd_h_put_16(abfd, val, ptr) \
BFD_SEND(abfd, bfd_h_putx16,(val,ptr))
#define bfd_h_put_signed_16 \
bfd_h_put_16
#define bfd_h_get_16(abfd, ptr) \
BFD_SEND(abfd, bfd_h_getx16,(ptr))
#define bfd_h_get_signed_16(abfd, ptr) \
BFD_SEND(abfd, bfd_h_getx_signed_16, (ptr))
#define bfd_h_put_32(abfd, val, ptr) \
BFD_SEND(abfd, bfd_h_putx32,(val,ptr))
#define bfd_h_put_signed_32 \
bfd_h_put_32
#define bfd_h_get_32(abfd, ptr) \
BFD_SEND(abfd, bfd_h_getx32,(ptr))
#define bfd_h_get_signed_32(abfd, ptr) \
BFD_SEND(abfd, bfd_h_getx_signed_32, (ptr))
#define bfd_h_put_64(abfd, val, ptr) \
BFD_SEND(abfd, bfd_h_putx64,(val, ptr))
#define bfd_h_put_signed_64 \
bfd_h_put_64
#define bfd_h_get_64(abfd, ptr) \

                     BFD_SEND(abfd, bfd_h_getx64,(ptr))
     #define bfd_h_get_signed_64(abfd, ptr) \
                     BFD_SEND(abfd, bfd_h_getx_signed_64, (ptr))

2.16.0.5 bfd_log2

Synopsis

     unsigned int bfd_log2(bfd_vma x);

Description Return the log base 2 of the value supplied, rounded up. E.g., an x of 1025 returns 11.

2.17 File caching

The file caching mechanism is embedded within BFD and allows the application to open as many BFDs as it wants without regard to the underlying operating system’s file descriptor limit (often as low as 20 open files). The module in cache.c maintains a least recently used list of BFD_CACHE_MAX_OPEN files, and exports the name bfd_cache_lookup, which runs around and makes sure that the required BFD is open. If not, then it chooses a file to close, closes it and opens the one wanted, returning its file handle.

2.17.0.1 BFD_CACHE_MAX_OPEN macro

Description The maximum number of files which the cache will keep open at one time.

     #define BFD_CACHE_MAX_OPEN 10
2.17.0.2 bfd_last_cache

Synopsis

     extern bfd *bfd_last_cache;

Description Zero, or a pointer to the topmost BFD on the chain. This is used by the bfd_cache_lookup macro in ‘libbfd.h’ to determine when it can avoid a function call.

2.17.0.3 bfd_cache_lookup

Description Check to see if the required BFD is the same as the last one looked up. If so, then it can use the stream in the BFD with impunity, since it can’t have changed since the last lookup; otherwise, it has to perform the complicated lookup function.

     #define bfd_cache_lookup(x) \
         ((x)==bfd_last_cache? \
           (FILE*)(bfd_last_cache->iostream): \
            bfd_cache_lookup_worker(x))
2.17.0.4 bfd_cache_init

Synopsis

     boolean bfd_cache_init (bfd *abfd);

Description Add a newly opened BFD to the cache.

2.17.0.5 bfd_cache_close

Synopsis

     boolean bfd_cache_close (bfd *abfd);

Description Remove the BFD abfd from the cache. If the attached file is open, then close it too. Returns false is returned if closing the file fails, true is returned if all is well.

2.17.0.6 bfd_open_file

Synopsis

     FILE* bfd_open_file(bfd *abfd);

Description Call the OS to open a file for abfd. Return the FILE * (possibly NULL) that results from this operation. Set up the BFD so that future accesses know the file is open. If the FILE * returned is NULL, then it won’t have been put in the cache, so it won’t have to be removed from it.

2.17.0.7 bfd_cache_lookup_worker

Synopsis

     FILE *bfd_cache_lookup_worker(bfd *abfd);

Description Called when the macro bfd_cache_lookup fails to find a quick answer. Find a file descriptor for abfd. If necessary, it open it. If there are already more than BFD_CACHE_MAX_OPEN files open, it tries to close one first, to avoid running out of file descriptors.

2.18 Linker Functions

The linker uses three special entry points in the BFD target vector. It is not necessary to write special routines for these entry points when creating a new BFD back end, since generic versions are provided. However, writing them can speed up linking and make it use significantly less runtime memory. The first routine creates a hash table used by the other routines. The second routine adds the symbols from an object file to the hash table. The third routine takes all the object files and links them together to create the output file. These routines are designed so that the linker proper does not need to know anything about the symbols in the object files that it is linking. The linker merely arranges the sections as directed by the linker script and lets BFD handle the details of symbols and relocs. The second routine and third routines are passed a pointer to a struct bfd_link_info structure (defined in bfdlink.h) which holds information relevant to the link, including the linker hash table (which was created by the first routine) and a set of callback functions to the linker proper.

The generic linker routines are in linker.c, and use the header file genlink.h. As of this writing, the only back ends which have implemented versions of these routines are a.out (in aoutx.h) and ECOFF (in ecoff.c). The a.out routines are used as examples throughout this section.

2.18.1 Creating a linker hash table

The linker routines must create a hash table, which must be derived from struct bfd_ link_hash_table described in bfdlink.c. See Section 2.19 [Hash Tables], page 89, for information on how to create a derived hash table. This entry point is called using the target vector of the linker output file. The _bfd_link_hash_table_create entry point must allocate and initialize an instance of the desired hash table. If the back end does not require any additional information to be stored with the entries in the hash table, the entry point may simply create a struct bfd_link_hash_table. Most likely, however, some additional information will be needed. For example, with each entry in the hash table the a.out linker keeps the index the symbol has in the final output file (this index number is used so that when doing a relocateable link the symbol index used in the output file can be quickly filled in when copying over a reloc). The a.out linker code defines the required structures and functions for a hash table derived from struct bfd_link_hash_table. The a.out linker hash table is created by the function NAME(aout,link_hash_table_create); it simply allocates space for the hash table, initializes it, and returns a pointer to it. When writing the linker routines for a new back end, you will generally not know exactly which fields will be required until you have finished. You should simply create a new hash table which defines no additional fields, and then simply add fields as they become necessary.

2.18.2 Adding symbols to the hash table

The linker proper will call the _bfd_link_add_symbols entry point for each object file or archive which is to be linked (typically these are the files named on the command line, but some may also come from the linker script). The entry point is responsible for examining the file. For an object file, BFD must add any relevant symbol information to the hash table. For an archive, BFD must determine which elements of the archive should be used and adding them to the link. The a.out version of this entry point is NAME(aout,link_add_symbols).

2.18.2.1 Differing file formats

Normally all the files involved in a link will be of the same format, but it is also possible to link together different format object files, and the back end must support that. The _bfd_link_add_symbols entry point is called via the target vector of the file to be added. This has an important consequence: the function may not assume that the hash table is the type created by the corresponding _bfd_link_hash_table_create vector. All the _ bfd_link_add_symbols function can assume about the hash table is that it is derived from

struct bfd_link_hash_table.

Sometimes the _bfd_link_add_symbols function must store some information in the hash table entry to be used by the _bfd_final_link function. In such a case the creator field of the hash table must be checked to make sure that the hash table was created by an object file of the same format. The _bfd_final_link routine must be prepared to handle a hash entry without any extra information added by the _bfd_link_add_symbols function. A hash entry without extra information will also occur when the linker script directs the linker to create a symbol. Note that, regardless of how a hash table entry is added, all the fields will be initialized to some sort of null value by the hash table entry initialization function. See ecoff_link_add_externals for an example of how to check the creator field before saving information (in this case, the ECOFF external symbol debugging information) in a hash table entry.

2.18.2.2 Adding symbols from an object file

When the _bfd_link_add_symbols routine is passed an object file, it must add all externally visible symbols in that object file to the hash table. The actual work of adding the symbol to the hash table is normally handled by the function _bfd_generic_link_ add_one_symbol. The _bfd_link_add_symbols routine is responsible for reading all the symbols from the object file and passing the correct information to _bfd_generic_link_

add_one_symbol.
The _bfd_link_add_symbols routine should not use bfd_canonicalize_symtab to read

the symbols. The point of providing this routine is to avoid the overhead of converting the symbols into generic asymbol structures. _bfd_generic_link_add_one_symbol handles the details of combining common symbols, warning about multiple definitions, and so forth. It takes arguments which describe the symbol to add, notably symbol flags, a section, and an offset. The symbol flags include such things as BSF_WEAK or BSF_INDIRECT. The section is a section in the object file, or

something like bfd_und_section_ptr for an undefined symbol or bfd_com_section_ptr

for a common symbol. If the _bfd_final_link routine is also going to need to read the symbol information, the _bfd_link_add_symbols routine should save it somewhere attached to the object file BFD. However, the information should only be saved if the keep_memory field of the info argument is true, so that the -no-keep-memory linker switch is effective. The a.out function which adds symbols from an object file is aout_link_add_object_ symbols, and most of the interesting work is in aout_link_add_symbols. The latter saves pointers to the hash tables entries created by _bfd_generic_link_add_one_symbol indexed by symbol number, so that the _bfd_final_link routine does not have to call the hash table lookup routine to locate the entry.

2.18.2.3 Adding symbols from an archive

When the _bfd_link_add_symbols routine is passed an archive, it must look through the symbols defined by the archive and decide which elements of the archive should be included in the link. For each such element it must call the add_archive_element linker callback, and it must add the symbols from the object file to the linker hash table. In most cases the work of looking through the symbols in the archive should be done

by the _bfd_generic_link_add_archive_symbols function. This function builds a hash

table from the archive symbol table and looks through the list of undefined symbols to see which elements should be included. _bfd_generic_link_add_archive_symbols is passed a function to call to make the final decision about adding an archive element to the link and to do the actual work of adding the symbols to the linker hash table.

The function passed to _bfd_generic_link_add_archive_symbols must read the symbols

of the archive element and decide whether the archive element should be included in the link. If the element is to be included, the add_archive_element linker callback routine must be called with the element as an argument, and the elements symbols must be added to the linker hash table just as though the element had itself been passed to the _bfd_

link_add_symbols function.

When the a.out _bfd_link_add_symbols function receives an archive, it calls _bfd_

generic_link_add_archive_symbols passing aout_link_check_archive_element as
the function argument. aout_link_check_archive_element calls aout_link_check_

ar_symbols. If the latter decides to add the element (an element is only added if it provides a real, non-common, definition for a previously undefined or common symbol) it

calls the add_archive_element callback and then aout_link_check_archive_element

calls aout_link_add_symbols to actually add the symbols to the linker hash table. The ECOFF back end is unusual in that it does not normally call _bfd_generic_link_ add_archive_symbols, because ECOFF archives already contain a hash table of symbols. The ECOFF back end searches the archive itself to avoid the overhead of creating a new hash table.

2.18.3 Performing the final link

When all the input files have been processed, the linker calls the _bfd_final_link entry point of the output BFD. This routine is responsible for producing the final output file, which has several aspects. It must relocate the contents of the input sections and copy the data into the output sections. It must build an output symbol table including any local symbols from the input files and the global symbols from the hash table. When producing relocateable output, it must modify the input relocs and write them into the output file. There may also be object format dependent work to be done. The linker will also call the write_object_contents entry point when the BFD is closed. The two entry points must work together in order to produce the correct output file. The details of how this works are inevitably dependent upon the specific object file format.

The a.out _bfd_final_link routine is NAME(aout,final_link).

2.18.3.1 Information provided by the linker

Before the linker calls the _bfd_final_link entry point, it sets up some data structures for the function to use. The input_bfds field of the bfd_link_info structure will point to a list of all the input files included in the link. These files are linked through the link_next field of the bfd structure. Each section in the output file will have a list of link_order structures attached to the link_ order_head field (the link_order structure is defined in bfdlink.h). These structures describe how to create the contents of the output section in terms of the contents of various input sections, fill constants, and, eventually, other types of information. They also describe relocs that must be created by the BFD backend, but do not correspond to any input file; this is used to support -Ur, which builds constructors while generating a relocateable object file.

2.18.3.2 Relocating the section contents

The _bfd_final_link function should look through the link_order structures attached to each section of the output file. Each link_order structure should either be handled specially, or it should be passed to the function _bfd_default_link_order which will do

the right thing (_bfd_default_link_order is defined in linker.c).

For efficiency, a link_order of type bfd_indirect_link_order whose associated section belongs to a BFD of the same format as the output BFD must be handled specially. This type of link_order describes part of an output section in terms of a section belonging to one of the input files. The _bfd_final_link function should read the contents of the section and any associated relocs, apply the relocs to the section contents, and write out the modified section contents. If performing a relocateable link, the relocs themselves must also be modified and written out.

The functions _bfd_relocate_contents and _bfd_final_link_relocate provide some

general support for performing the actual relocations, notably overflow checking. Their arguments include information about the symbol the relocation is against and a reloc_howto_ type argument which describes the relocation to perform. These functions are defined in

reloc.c.

The a.out function which handles reading, relocating, and writing section contents is aout_

link_input_section. The actual relocation is done in aout_link_input_section_std
and aout_link_input_section_ext.

2.18.3.3 Writing the symbol table

The _bfd_final_link function must gather all the symbols in the input files and write them out. It must also write out all the symbols in the global hash table. This must be controlled by the strip and discard fields of the bfd_link_info structure. The local symbols of the input files will not have been entered into the linker hash table. The _bfd_final_link routine must consider each input file and include the symbols in the output file. It may be convenient to do this when looking through the link_order structures, or it may be done by stepping through the input_bfds list. The _bfd_final_link routine must also traverse the global hash table to gather all the externally visible symbols. It is possible that most of the externally visible symbols may be written out when considering the symbols of each input file, but it is still necessary to traverse the hash table since the linker script may have defined some symbols that are not in any of the input files. The strip field of the bfd_link_info structure controls which symbols are written out. The possible values are listed in bfdlink.h. If the value is strip_some, then the keep_hash field of the bfd_link_info structure is a hash table of symbols to keep; each symbol should be looked up in this hash table, and only symbols which are present should be included in the output file.

If the strip field of the bfd_link_info structure permits local symbols to be written out, the discard field is used to further controls which local symbols are included in the output file. If the value is discard_l, then all local symbols which begin with a certain prefix are discarded; this is controlled by the bfd_is_local_label_name entry point. The a.out backend handles symbols by calling aout_link_write_symbols on each input BFD and then traversing the global hash table with the function aout_link_write_other_ symbol. It builds a string table while writing out the symbols, which is written to the output

file at the end of NAME(aout,final_link).
2.18.3.4 bfd_link_split_section

Synopsis

     boolean bfd_link_split_section(bfd *abfd, asection *sec);

Description Return nonzero if sec should be split during a reloceatable or final link.

     #define bfd_link_split_section(abfd, sec) \
            BFD_SEND (abfd, _bfd_link_split_section, (abfd, sec))

2.19 Hash Tables

BFD provides a simple set of hash table functions. Routines are provided to initialize a hash table, to free a hash table, to look up a string in a hash table and optionally create an entry for it, and to traverse a hash table. There is currently no routine to delete an string from a hash table. The basic hash table does not permit any data to be stored with a string. However, a hash table is designed to present a base class from which other types of hash tables may be derived. These derived types may store additional information with the string. Hash tables were implemented in this way, rather than simply providing a data pointer in a hash table entry, because they were designed for use by the linker back ends. The linker may create thousands of hash table entries, and the overhead of allocating private data and storing and following pointers becomes noticeable. The basic hash table code is in hash.c.

2.19.1 Creating and freeing a hash table

To create a hash table, create an instance of a struct bfd_hash_table (defined in bfd.h) and call bfd_hash_table_init (if you know approximately how many entries you will need, the function bfd_hash_table_init_n, which takes a size argument, may be used). bfd_hash_table_init returns false if some sort of error occurs. The function bfd_hash_table_init take as an argument a function to use to create new entries. For a basic hash table, use the function bfd_hash_newfunc. See Section 2.19.4 [Deriving a New Hash Table Type], page 90, for why you would want to use a different value for this argument. bfd_hash_table_init will create an objalloc which will be used to allocate new entries. You may allocate memory on this objalloc using bfd_hash_allocate.

Use bfd_hash_table_free to free up all the memory that has been allocated for a hash table. This will not free up the struct bfd_hash_table itself, which you must provide.

2.19.2 Looking up or entering a string

The function bfd_hash_lookup is used both to look up a string in the hash table and to create a new entry.

If the create argument is false, bfd_hash_lookup will look up a string. If the string is found, it will returns a pointer to a struct bfd_hash_entry. If the string is not found in the table bfd_hash_lookup will return NULL. You should not modify any of the fields in

the returns struct bfd_hash_entry.

If the create argument is true, the string will be entered into the hash table if it is not already there. Either way a pointer to a struct bfd_hash_entry will be returned, either to the existing structure or to a newly created one. In this case, a NULL return means that an error occurred.

If the create argument is true, and a new entry is created, the copy argument is used to decide whether to copy the string onto the hash table objalloc or not. If copy is passed as false, you must be careful not to deallocate or modify the string as long as the hash table exists.

2.19.3 Traversing a hash table

The function bfd_hash_traverse may be used to traverse a hash table, calling a function on each element. The traversal is done in a random order.

bfd_hash_traverse takes as arguments a function and a generic void * pointer. The function is called with a hash table entry (a struct bfd_hash_entry *) and the generic pointer passed to bfd_hash_traverse. The function must return a boolean value, which indicates whether to continue traversing the hash table. If the function returns false, bfd_hash_traverse will stop the traversal and return immediately.

2.19.4 Deriving a new hash table type

Many uses of hash tables want to store additional information which each entry in the hash table. Some also find it convenient to store additional information with the hash table itself. This may be done using a derived hash table.

Since C is not an object oriented language, creating a derived hash table requires sticking together some boilerplate routines with a few differences specific to the type of hash table you want to create.

An example of a derived hash table is the linker hash table. The structures for this are defined in bfdlink.h. The functions are in linker.c.

You may also derive a hash table from an already derived hash table. For example, the a.out linker backend code uses a hash table derived from the linker hash table.

2.19.4.1 Define the derived structures

You must define a structure for an entry in the hash table, and a structure for the hash table itself. The first field in the structure for an entry in the hash table must be of the type used for an entry in the hash table you are deriving from. If you are deriving from a basic hash table this is struct bfd_hash_entry, which is defined in bfd.h. The first field in the structure for the hash table itself must be of the type of the hash table you are deriving from itself. If you are deriving from a basic hash table, this is struct bfd_hash_table. For example, the linker hash table defines struct bfd_link_hash_entry (in bfdlink.h). The first field, root, is of type struct bfd_hash_entry. Similarly, the first field in struct

bfd_link_hash_table, table, is of type struct bfd_hash_table.

2.19.4.2 Write the derived creation routine

You must write a routine which will create and initialize an entry in the hash table. This routine is passed as the function argument to bfd_hash_table_init. In order to permit other hash tables to be derived from the hash table you are creating, this routine must be written in a standard way. The first argument to the creation routine is a pointer to a hash table entry. This may be NULL, in which case the routine should allocate the right amount of space. Otherwise the space has already been allocated by a hash table type derived from this one. After allocating space, the creation routine must call the creation routine of the hash table type it is derived from, passing in a pointer to the space it just allocated. This will initialize any fields used by the base hash table. Finally the creation routine must initialize any local fields for the new hash table type. Here is a boilerplate example of a creation routine. function name is the name of the routine. entry type is the type of an entry in the hash table you are creating. base newfunc is the name of the creation routine of the hash table type your hash table is derived from.

     struct bfd_hash_entry *
     function name (entry, table, string)
          struct bfd_hash_entry *entry;
          struct bfd_hash_table *table;
          const char *string;
     {
       struct entry type *ret = (entry type *) entry;
      /* Allocate the structure if it has not already been allocated by a
         derived class.  */
       if (ret == (entry type *) NULL)
         {

ret = ((entry type *)

                  bfd_hash_allocate (table, sizeof (entry type)));
           if (ret == (entry type *) NULL)
             return NULL;
         }

      /* Call the allocation method of the base class.  */

ret = ((entry type *)

             base newfunc ((struct bfd_hash_entry *) ret, table, string));
      /* Initialize the local fields here.  */
       return (struct bfd_hash_entry *) ret;
     }

Description The creation routine for the linker hash table, which is in linker.c, looks just like this

example. function name is _bfd_link_hash_newfunc. entry type is struct bfd_link_

hash_entry. base newfunc is bfd_hash_newfunc, the creation routine for a basic hash table. _bfd_link_hash_newfunc also initializes the local fields in a linker hash table entry: type,

written and next.

2.19.4.3 Write other derived routines

You will want to write other routines for your new hash table, as well. You will want an initialization routine which calls the initialization routine of the hash table you are deriving from and initializes any other local fields. For the linker hash table, this

is _bfd_link_hash_table_init in linker.c.

You will want a lookup routine which calls the lookup routine of the hash table you are deriving from and casts the result. The linker hash table uses bfd_link_hash_lookup in linker.c (this actually takes an additional argument which it uses to decide how to return the looked up value). You may want a traversal routine. This should just call the traversal routine of the hash table you are deriving from with appropriate casts. The linker hash table uses bfd_link_

hash_traverse in linker.c.

These routines may simply be defined as macros. For example, the a.out backend linker hash table, which is derived from the linker hash table, uses macros for the lookup and

traversal routines. These are aout_link_hash_lookup and aout_link_hash_traverse in

aoutx.h.

3 BFD back ends

All of BFD lives in one directory.

3.1 a.out backends

Description BFD supports a number of different flavours of a.out format, though the major differences are only the sizes of the structures on disk, and the shape of the relocation information. The support is split into a basic support file ‘aoutx.h’ and other files which derive functions from the base. One derivation file is ‘aoutf1.h’ (for a.out flavour 1), and adds to the basic a.out functions support for sun3, sun4, 386 and 29k a.out files, to create a target jump vector for a specific target. This information is further split out into more specific files for each machine, including ‘sunos.c’ for sun3 and sun4, ‘newsos3.c’ for the Sony NEWS, and ‘demo64.c’ for a demonstration of a 64 bit a.out format. The base file ‘aoutx.h’ defines general mechanisms for reading and writing records to and from disk and various other methods which BFD requires. It is included by ‘aout32.c’

and ‘aout64.c’ to form the names aout_32_swap_exec_header_in, aout_64_swap_exec_
header_in, etc.

As an example, this is what goes on to make the back end for a sun4, from ‘aout32.c’:

            #define ARCH_SIZE 32
            #include "aoutx.h"

Which exports names:

            ...
            aout_32_canonicalize_reloc
            aout_32_find_nearest_line
            aout_32_get_lineno
            aout_32_get_reloc_upper_bound
            ...
from ‘sunos.c’:
            #define TARGET_NAME "a.out-sunos-big"
            #define VECNAME    sunos_big_vec
            #include "aoutf1.h"

requires all the names from ‘aout32.c’, and produces the jump vector

            sunos_big_vec

The file ‘host-aout.c’ is a special case. It is for a large set of hosts that use “more or less standard” a.out files, and for which cross-debugging is not interesting. It uses the standard 32-bit a.out support routines, but determines the file offsets and addresses of the text, data, and BSS sections, the machine architecture and machine type, and the entry point address, in a host-dependent manner. Once these values have been determined, generic code is used to handle the object file. When porting it to run on a new system, you must supply:

             HOST_PAGE_SIZE
             HOST_SEGMENT_SIZE
             HOST_MACHINE_ARCH       (optional)

             HOST_MACHINE_MACHINE    (optional)
             HOST_TEXT_START_ADDR
             HOST_STACK_END_ADDR

in the file ‘../include/sys/h-XXX.h’ (for your host). These values, plus the structures and macros defined in ‘a.out.h’ on your host system, will produce a BFD target that will access ordinary a.out files on your host. To configure a new machine to use ‘host-aout.c’, specify:

            TDEFAULTS = -DDEFAULT_VECTOR=host_aout_big_vec
            TDEPFILES= host-aout.o trad-core.o

in the ‘config/XXX.mt’ file, and modify ‘configure.in’ to use the ‘XXX.mt’ file (by setting "bfd_target=XXX") when your configuration is selected.

3.1.1 Relocations

Description The file ‘aoutx.h’ provides for both the standard and extended forms of a.out relocation records.

The standard records contain only an address, a symbol index, and a type field. The extended records (used on 29ks and sparcs) also have a full integer for an addend.

3.1.2 Internal entry points

Description ‘aoutx.h’ exports several routines for accessing the contents of an a.out file, which are gathered and exported in turn by various format specific files (eg sunos.c).

3.1.2.1 aout_size_swap_exec_header_in

Synopsis

     void aout_size_swap_exec_header_in,
        (bfd *abfd,
         struct external_exec *raw_bytes,
         struct internal_exec *execp);

Description Swap the information in an executable header raw bytes taken from a raw byte stream memory image into the internal exec header structure execp.

3.1.2.2 aout_size_swap_exec_header_out

Synopsis

     void aout_size_swap_exec_header_out
        (bfd *abfd,
         struct internal_exec *execp,
         struct external_exec *raw_bytes);

Description Swap the information in an internal exec header structure execp into the buffer raw bytes ready for writing to disk.

3.1.2.3 aout_size_some_aout_object_p

Synopsis

     const bfd_target *aout_size_some_aout_object_p
        (bfd *abfd,
         const bfd_target *(*callback_to_real_object_p)());

Description Some a.out variant thinks that the file open in abfd checking is an a.out file. Do some more checking, and set up for access if it really is. Call back to the calling environment’s "finish up" function just before returning, to handle any last-minute setup.

3.1.2.4 aout_size_mkobject

Synopsis

     boolean aout_size_mkobject, (bfd *abfd);

Description Initialize BFD abfd for use with a.out files.

3.1.2.5 aout_size_machine_type

Synopsis

     enum machine_type  aout_size_machine_type
        (enum bfd_architecture arch,
         unsigned long machine));

Description Keep track of machine architecture and machine type for a.out’s. Return the machine_ type for a particular architecture and machine, or M_UNKNOWN if that exact architecture and machine can’t be represented in a.out format. If the architecture is understood, machine type 0 (default) is always understood.

3.1.2.6 aout_size_set_arch_mach

Synopsis

     boolean aout_size_set_arch_mach,
        (bfd *,
         enum bfd_architecture arch,
         unsigned long machine));

Description Set the architecture and the machine of the BFD abfd to the values arch and machine. Verify that abfd’s format can support the architecture required.

3.1.2.7 aout_size_new_section_hook

Synopsis

     boolean aout_size_new_section_hook,
        (bfd *abfd,
         asection *newsect));

Description Called by the BFD in response to a bfd_make_section request.

3.2 coff backends

BFD supports a number of different flavours of coff format. The major differences between formats are the sizes and alignments of fields in structures on disk, and the occasional extra field. Coff in all its varieties is implemented with a few common files and a number of implementation specific files. For example, The 88k bcs coff format is implemented in the file ‘coff-m88k.c’. This file #includes ‘coff/m88k.h’ which defines the external structure of the coff format for the 88k, and ‘coff/internal.h’ which defines the internal structure. ‘coff-m88k.c’ also defines the relocations used by the 88k format See Section 2.10 [Relocations], page 37. The Intel i960 processor version of coff is implemented in ‘coff-i960.c’. This file has the same structure as ‘coff-m88k.c’, except that it includes ‘coff/i960.h’ rather than

‘coff-m88k.h’.

3.2.1 Porting to a new version of coff

The recommended method is to select from the existing implementations the version of coff which is most like the one you want to use. For example, we’ll say that i386 coff is the one you select, and that your coff flavour is called foo. Copy ‘i386coff.c’ to ‘foocoff.c’, copy

‘../include/coff/i386.h’ to ‘../include/coff/foo.h’, and add the lines to ‘targets.c’

and ‘Makefile.in’ so that your new back end is used. Alter the shapes of the structures in ‘../include/coff/foo.h’ so that they match what you need. You will probably also have to add #ifdefs to the code in ‘coff/internal.h’ and ‘coffcode.h’ if your version of coff is too wild. You can verify that your new BFD backend works quite simply by building ‘objdump’ from the ‘binutils’ directory, and making sure that its version of what’s going on and your host system’s idea (assuming it has the pretty standard coff dump utility, usually called att-dump or just dump) are the same. Then clean up your code, and send what you’ve done to Cygnus. Then your stuff will be in the next release, and you won’t have to keep integrating it.

3.2.2 How the coff backend works

3.2.2.1 File layout

The Coff backend is split into generic routines that are applicable to any Coff target and routines that are specific to a particular target. The target-specific routines are further split into ones which are basically the same for all Coff targets except that they use the external symbol format or use different values for certain constants. The generic routines are in ‘coffgen.c’. These routines work for any Coff target. They use some hooks into the target specific code; the hooks are in a bfd_coff_backend_data structure, one of which exists for each target. The essentially similar target-specific routines are in ‘coffcode.h’. This header file includes executable C code. The various Coff targets first include the appropriate Coff header file, make any special defines that are needed, and then include ‘coffcode.h’.

Some of the Coff targets then also have additional routines in the target source file itself. For example, ‘coff-i960.c’ includes ‘coff/internal.h’ and ‘coff/i960.h’. It then defines a few constants, such as I960, and includes ‘coffcode.h’. Since the i960 has complex relocation types, ‘coff-i960.c’ also includes some code to manipulate the i960 relocs. This code is not in ‘coffcode.h’ because it would not be used by any other target.

3.2.2.2 Bit twiddling

Each flavour of coff supported in BFD has its own header file describing the external layout of the structures. There is also an internal description of the coff layout, in ‘coff/internal.h’. A major function of the coff backend is swapping the bytes and twiddling the bits to translate the external form of the structures into the normal internal form. This is all performed in the bfd_swap thing direction routines. Some elements are different sizes between different versions of coff; it is the duty of the coff version specific include file to override the definitions of various packing routines in ‘coffcode.h’. E.g., the size of line number entry in coff is sometimes 16 bits, and sometimes 32 bits. #defineing PUT_LNSZ_LNNO and GET_LNSZ_ LNNO will select the correct one. No doubt, some day someone will find a version of coff which has a varying field size not catered to at the moment. To port BFD, that person will have to add more #defines. Three of the bit twiddling routines are exported to gdb;

coff_swap_aux_in, coff_swap_sym_in and coff_swap_lineno_in. GDB reads the symbol

table on its own, but uses BFD to fix things up. More of the bit twiddlers are exported

for gas; coff_swap_aux_out, coff_swap_sym_out, coff_swap_lineno_out, coff_swap_
reloc_out, coff_swap_filehdr_out, coff_swap_aouthdr_out, coff_swap_scnhdr_out.

Gas currently keeps track of all the symbol table and reloc drudgery itself, thereby saving the internal BFD overhead, but uses BFD to swap things on the way out, making cross ports much safer. Doing so also allows BFD (and thus the linker) to use the same header files as gas, which makes one avenue to disaster disappear.

3.2.2.3 Symbol reading

The simple canonical form for symbols used by BFD is not rich enough to keep all the information available in a coff symbol table. The back end gets around this problem by keeping the original symbol table around, "behind the scenes". When a symbol table is requested (through a call to bfd_canonicalize_symtab), a request gets through to coff_get_normalized_symtab. This reads the symbol table from the coff file and swaps all the structures inside into the internal form. It also fixes up all the pointers in the table (represented in the file by offsets from the first symbol in the table) into physical pointers to elements in the new internal table. This involves some work since the meanings of fields change depending upon context: a field that is a pointer to another structure in the symbol table at one moment may be the size in bytes of a structure at the next. Another pass is made over the table. All symbols which mark file names (C_FILE symbols) are modified so that the internal string points to the value in the auxent (the real filename) rather than the normal text associated with the symbol (".file"). At this time the symbol names are moved around. Coff stores all symbols less than nine characters long physically within the symbol table; longer strings are kept at the end of the file in the string table. This pass moves all strings into memory and replaces them with pointers to the strings.

The symbol table is massaged once again, this time to create the canonical table used by the BFD application. Each symbol is inspected in turn, and a decision made (using the sclass field) about the various flags to set in the asymbol. See Section 2.7 [Symbols], page 27. The generated canonical table shares strings with the hidden internal symbol table. Any linenumbers are read from the coff file too, and attached to the symbols which own the functions the linenumbers belong to.

3.2.2.4 Symbol writing

Writing a symbol to a coff file which didn’t come from a coff file will lose any debugging information. The asymbol structure remembers the BFD from which the symbol was taken, and on output the back end makes sure that the same destination target as source target is present. When the symbols have come from a coff file then all the debugging information is preserved. Symbol tables are provided for writing to the back end in a vector of pointers to pointers. This allows applications like the linker to accumulate and output large symbol tables without having to do too much byte copying. This function runs through the provided symbol table and patches each symbol marked as a file place holder (C_FILE) to point to the next file place holder in the list. It also marks each offset field in the list with the offset from the first symbol of the current symbol. Another function of this procedure is to turn the canonical value form of BFD into the form used by coff. Internally, BFD expects symbol values to be offsets from a section base; so a symbol physically at 0x120, but in a section starting at 0x100, would have the value 0x20. Coff expects symbols to contain their final value, so symbols have their values changed at this point to reflect their sum with their owning section. This transformation uses the output_section field of the asymbol’s asection See Section 2.6 [Sections], page 15.

  • coff_mangle_symbols

This routine runs though the provided symbol table and uses the offsets generated by the previous pass and the pointers generated when the symbol table was read in to create the structured hierachy required by coff. It changes each pointer to a symbol into the index into the symbol table of the asymbol.

  • coff_write_symbols

This routine runs through the symbol table and patches up the symbols from their internal form into the coff way, calls the bit twiddlers, and writes out the table to the file.

3.2.2.5 coff_symbol_type

Description The hidden information for an asymbol is described in a combined_entry_type:

     typedef struct coff_ptr_struct
     {
            /* Remembers the offset from the first symbol in the file for
               this symbol. Generated by coff_renumber_symbols. */

unsigned int offset;
       /* Should the value of this symbol be renumbered.  Used for
          XCOFF C_BSTAT symbols.  Set by coff_slurp_symbol_table.  */
unsigned int fix_value : 1;
       /* Should the tag field of this symbol be renumbered.
          Created by coff_pointerize_aux. */
unsigned int fix_tag : 1;
       /* Should the endidx field of this symbol be renumbered.
          Created by coff_pointerize_aux. */
unsigned int fix_end : 1;
       /* Should the x_csect.x_scnlen field be renumbered.
          Created by coff_pointerize_aux. */
unsigned int fix_scnlen : 1;
       /* Fix up an XCOFF C_BINCL/C_EINCL symbol.  The value is the
          index into the line number entries.  Set by
          coff_slurp_symbol_table.  */
unsigned int fix_line : 1;
       /* The container for the symbol structure as read and translated
           from the file. */
union {
   union internal_auxent auxent;
   struct internal_syment syment;
 } u;
} combined_entry_type;

/* Each canonical asymbol really looks like this: */
typedef struct coff_symbol_struct
{
   /* The actual symbol which the rest of BFD works with */
asymbol symbol;
   /* A pointer to the hidden information for this symbol */
combined_entry_type *native;
   /* A pointer to the linenumber information for this symbol */
struct lineno_cache_entry *lineno;
   /* Have the line numbers been relocated yet ? */
boolean done_lineno;
} coff_symbol_type;

3.2.2.6 bfd_coff_backend_data
/* COFF symbol classifications.  */
enum coff_symbol_classification
{
/* Global symbol.  */
COFF_SYMBOL_GLOBAL,
/* Common symbol.  */
COFF_SYMBOL_COMMON,
/* Undefined symbol.  */
COFF_SYMBOL_UNDEFINED,
/* Local symbol.  */
COFF_SYMBOL_LOCAL,
/* PE section symbol.  */
COFF_SYMBOL_PE_SECTION
};

Special entry points for gdb to swap in coff symbol table parts:

typedef struct
{
void (*_bfd_coff_swap_aux_in) PARAMS ((
bfd            *abfd,
PTR             ext,
int             type,
int             class,
int             indaux,
int             numaux,
PTR             in));
void (*_bfd_coff_swap_sym_in) PARAMS ((
bfd            *abfd ,
PTR             ext,
PTR             in));
void (*_bfd_coff_swap_lineno_in) PARAMS ((
bfd            *abfd,
PTR            ext,
PTR             in));

Special entry points for gas to swap out coff parts:

unsigned int (*_bfd_coff_swap_aux_out) PARAMS ((
bfd     *abfd,
PTR     in,
int     type,
int     class,
int     indaux,
int     numaux,
PTR     ext));

unsigned int (*_bfd_coff_swap_sym_out) PARAMS ((
     bfd      *abfd,
     PTR      in,
     PTR      ext));
unsigned int (*_bfd_coff_swap_lineno_out) PARAMS ((
      bfd     *abfd,
      PTR     in,
      PTR     ext));
unsigned int (*_bfd_coff_swap_reloc_out) PARAMS ((
      bfd     *abfd,
      PTR     src,
      PTR     dst));
unsigned int (*_bfd_coff_swap_filehdr_out) PARAMS ((
      bfd     *abfd,
      PTR     in,
      PTR     out));
unsigned int (*_bfd_coff_swap_aouthdr_out) PARAMS ((
      bfd     *abfd,
      PTR     in,
      PTR     out));
unsigned int (*_bfd_coff_swap_scnhdr_out) PARAMS ((
      bfd     *abfd,
      PTR     in,
      PTR     out));

Special entry points for generic COFF routines to call target dependent COFF routines:

unsigned int _bfd_filhsz;
unsigned int _bfd_aoutsz;
unsigned int _bfd_scnhsz;
unsigned int _bfd_symesz;
unsigned int _bfd_auxesz;
unsigned int _bfd_relsz;
unsigned int _bfd_linesz;
unsigned int _bfd_filnmlen;
boolean _bfd_coff_long_filenames;
boolean _bfd_coff_long_section_names;
unsigned int _bfd_coff_default_section_alignment_power;
void (*_bfd_coff_swap_filehdr_in) PARAMS ((
      bfd     *abfd,
      PTR     ext,
      PTR     in));
void (*_bfd_coff_swap_aouthdr_in) PARAMS ((
      bfd     *abfd,

PTR     ext,
PTR     in));
void (*_bfd_coff_swap_scnhdr_in) PARAMS ((
bfd     *abfd,
PTR     ext,
PTR     in));
void (*_bfd_coff_swap_reloc_in) PARAMS ((
bfd     *abfd,
PTR     ext,
PTR     in));
boolean (*_bfd_coff_bad_format_hook) PARAMS ((
bfd     *abfd,
PTR     internal_filehdr));
boolean (*_bfd_coff_set_arch_mach_hook) PARAMS ((
bfd     *abfd,
PTR     internal_filehdr));
PTR (*_bfd_coff_mkobject_hook) PARAMS ((
bfd     *abfd,
PTR     internal_filehdr,
PTR     internal_aouthdr));
flagword (*_bfd_styp_to_sec_flags_hook) PARAMS ((
bfd     *abfd,
PTR     internal_scnhdr,
const char *name,
asection *section));
void (*_bfd_set_alignment_hook) PARAMS ((
bfd     *abfd,
asection *sec,
PTR     internal_scnhdr));
boolean (*_bfd_coff_slurp_symbol_table) PARAMS ((
bfd     *abfd));
boolean (*_bfd_coff_symname_in_debug) PARAMS ((
bfd     *abfd,
struct internal_syment *sym));
boolean (*_bfd_coff_pointerize_aux_hook) PARAMS ((
bfd *abfd,
combined_entry_type *table_base,
combined_entry_type *symbol,
unsigned int indaux,
combined_entry_type *aux));
boolean (*_bfd_coff_print_aux) PARAMS ((
bfd *abfd,
FILE *file,
combined_entry_type *table_base,
combined_entry_type *symbol,
combined_entry_type *aux,
unsigned int indaux));
void (*_bfd_coff_reloc16_extra_cases) PARAMS ((
bfd     *abfd,

struct bfd_link_info *link_info,
struct bfd_link_order *link_order,
arelent *reloc,
bfd_byte *data,
unsigned int *src_ptr,
unsigned int *dst_ptr));
int (*_bfd_coff_reloc16_estimate) PARAMS ((
bfd *abfd,
asection *input_section,
arelent *r,
unsigned int shrink,
struct bfd_link_info *link_info));
enum coff_symbol_classification (*_bfd_coff_classify_symbol) PARAMS ((
bfd *abfd,
struct internal_syment *));
boolean (*_bfd_coff_compute_section_file_positions) PARAMS ((
bfd *abfd));
boolean (*_bfd_coff_start_final_link) PARAMS ((
bfd *output_bfd,
struct bfd_link_info *info));
boolean (*_bfd_coff_relocate_section) PARAMS ((
bfd *output_bfd,
struct bfd_link_info *info,
bfd *input_bfd,
asection *input_section,
bfd_byte *contents,
struct internal_reloc *relocs,
struct internal_syment *syms,
asection **sections));
reloc_howto_type *(*_bfd_coff_rtype_to_howto) PARAMS ((
bfd *abfd,
asection *sec,
struct internal_reloc *rel,
struct coff_link_hash_entry *h,
struct internal_syment *sym,
bfd_vma *addendp));
boolean (*_bfd_coff_adjust_symndx) PARAMS ((
bfd *obfd,
struct bfd_link_info *info,
bfd *ibfd,
asection *sec,
struct internal_reloc *reloc,
boolean *adjustedp));
boolean (*_bfd_coff_link_add_one_symbol) PARAMS ((
struct bfd_link_info *info,
bfd *abfd,
const char *name,
flagword flags,
asection *section,

       bfd_vma value,
       const char *string,
       boolean copy,
       boolean collect,
       struct bfd_link_hash_entry **hashp));
 boolean (*_bfd_coff_link_output_has_begun) PARAMS ((
       bfd * abfd,
       struct coff_final_link_info * pfinfo));
 boolean (*_bfd_coff_final_link_postscript) PARAMS ((
       bfd * abfd,
       struct coff_final_link_info * pfinfo));
} bfd_coff_backend_data;
#define coff_backend_info(abfd) ((bfd_coff_backend_data *) (abfd)->xvec->backend_data)
#define bfd_coff_swap_aux_in(a,e,t,c,ind,num,i) \
        ((coff_backend_info (a)->_bfd_coff_swap_aux_in) (a,e,t,c,ind,num,i))
#define bfd_coff_swap_sym_in(a,e,i) \
        ((coff_backend_info (a)->_bfd_coff_swap_sym_in) (a,e,i))
#define bfd_coff_swap_lineno_in(a,e,i) \
        ((coff_backend_info ( a)->_bfd_coff_swap_lineno_in) (a,e,i))
#define bfd_coff_swap_reloc_out(abfd, i, o) \
        ((coff_backend_info (abfd)->_bfd_coff_swap_reloc_out) (abfd, i, o))
#define bfd_coff_swap_lineno_out(abfd, i, o) \
        ((coff_backend_info (abfd)->_bfd_coff_swap_lineno_out) (abfd, i, o))
#define bfd_coff_swap_aux_out(a,i,t,c,ind,num,o) \
        ((coff_backend_info (a)->_bfd_coff_swap_aux_out) (a,i,t,c,ind,num,o))
#define bfd_coff_swap_sym_out(abfd, i,o) \
        ((coff_backend_info (abfd)->_bfd_coff_swap_sym_out) (abfd, i, o))
#define bfd_coff_swap_scnhdr_out(abfd, i,o) \
        ((coff_backend_info (abfd)->_bfd_coff_swap_scnhdr_out) (abfd, i, o))
#define bfd_coff_swap_filehdr_out(abfd, i,o) \
        ((coff_backend_info (abfd)->_bfd_coff_swap_filehdr_out) (abfd, i, o))
#define bfd_coff_swap_aouthdr_out(abfd, i,o) \
        ((coff_backend_info (abfd)->_bfd_coff_swap_aouthdr_out) (abfd, i, o))
#define bfd_coff_filhsz(abfd) (coff_backend_info (abfd)->_bfd_filhsz)
#define bfd_coff_aoutsz(abfd) (coff_backend_info (abfd)->_bfd_aoutsz)

#define bfd_coff_scnhsz(abfd) (coff_backend_info (abfd)->_bfd_scnhsz)
#define bfd_coff_symesz(abfd) (coff_backend_info (abfd)->_bfd_symesz)
#define bfd_coff_auxesz(abfd) (coff_backend_info (abfd)->_bfd_auxesz)
#define bfd_coff_relsz(abfd)  (coff_backend_info (abfd)->_bfd_relsz)
#define bfd_coff_linesz(abfd) (coff_backend_info (abfd)->_bfd_linesz)
#define bfd_coff_filnmlen(abfd) (coff_backend_info (abfd)->_bfd_filnmlen)
#define bfd_coff_long_filenames(abfd) (coff_backend_info (abfd)->_bfd_coff_long_filenames)
#define bfd_coff_long_section_names(abfd) \
        (coff_backend_info (abfd)->_bfd_coff_long_section_names)
#define bfd_coff_default_section_alignment_power(abfd) \
        (coff_backend_info (abfd)->_bfd_coff_default_section_alignment_power)
#define bfd_coff_swap_filehdr_in(abfd, i,o) \
        ((coff_backend_info (abfd)->_bfd_coff_swap_filehdr_in) (abfd, i, o))
#define bfd_coff_swap_aouthdr_in(abfd, i,o) \
        ((coff_backend_info (abfd)->_bfd_coff_swap_aouthdr_in) (abfd, i, o))
#define bfd_coff_swap_scnhdr_in(abfd, i,o) \
        ((coff_backend_info (abfd)->_bfd_coff_swap_scnhdr_in) (abfd, i, o))
#define bfd_coff_swap_reloc_in(abfd, i, o) \
        ((coff_backend_info (abfd)->_bfd_coff_swap_reloc_in) (abfd, i, o))
#define bfd_coff_bad_format_hook(abfd, filehdr) \
        ((coff_backend_info (abfd)->_bfd_coff_bad_format_hook) (abfd, filehdr))
#define bfd_coff_set_arch_mach_hook(abfd, filehdr)\
        ((coff_backend_info (abfd)->_bfd_coff_set_arch_mach_hook) (abfd, filehdr))
#define bfd_coff_mkobject_hook(abfd, filehdr, aouthdr)\
        ((coff_backend_info (abfd)->_bfd_coff_mkobject_hook) (abfd, filehdr, aouthdr))
#define bfd_coff_styp_to_sec_flags_hook(abfd, scnhdr, name, section)\
        ((coff_backend_info (abfd)->_bfd_styp_to_sec_flags_hook)\
         (abfd, scnhdr, name, section))
#define bfd_coff_set_alignment_hook(abfd, sec, scnhdr)\
        ((coff_backend_info (abfd)->_bfd_set_alignment_hook) (abfd, sec, scnhdr))
#define bfd_coff_slurp_symbol_table(abfd)\
        ((coff_backend_info (abfd)->_bfd_coff_slurp_symbol_table) (abfd))
#define bfd_coff_symname_in_debug(abfd, sym)\
        ((coff_backend_info (abfd)->_bfd_coff_symname_in_debug) (abfd, sym))
#define bfd_coff_print_aux(abfd, file, base, symbol, aux, indaux)\
        ((coff_backend_info (abfd)->_bfd_coff_print_aux)\
         (abfd, file, base, symbol, aux, indaux))
#define bfd_coff_reloc16_extra_cases(abfd, link_info, link_order, reloc, data, src_ptr, dst_ptr)\

((coff_backend_info (abfd)->_bfd_coff_reloc16_extra_cases)\
 (abfd, link_info, link_order, reloc, data, src_ptr, dst_ptr))
#define bfd_coff_reloc16_estimate(abfd, section, reloc, shrink, link_info)\
((coff_backend_info (abfd)->_bfd_coff_reloc16_estimate)\
 (abfd, section, reloc, shrink, link_info))
#define bfd_coff_classify_symbol(abfd, sym)\
((coff_backend_info (abfd)->_bfd_coff_classify_symbol)\
 (abfd, sym))
#define bfd_coff_compute_section_file_positions(abfd)\
((coff_backend_info (abfd)->_bfd_coff_compute_section_file_positions)\
 (abfd))
#define bfd_coff_start_final_link(obfd, info)\
((coff_backend_info (obfd)->_bfd_coff_start_final_link)\
 (obfd, info))
#define bfd_coff_relocate_section(obfd,info,ibfd,o,con,rel,isyms,secs)\
((coff_backend_info (ibfd)->_bfd_coff_relocate_section)\
 (obfd, info, ibfd, o, con, rel, isyms, secs))
#define bfd_coff_rtype_to_howto(abfd, sec, rel, h, sym, addendp)\
((coff_backend_info (abfd)->_bfd_coff_rtype_to_howto)\
 (abfd, sec, rel, h, sym, addendp))
#define bfd_coff_adjust_symndx(obfd, info, ibfd, sec, rel, adjustedp)\
((coff_backend_info (abfd)->_bfd_coff_adjust_symndx)\
 (obfd, info, ibfd, sec, rel, adjustedp))
#define bfd_coff_link_add_one_symbol(info,abfd,name,flags,section,value,string,cp,coll,hashp)\
((coff_backend_info (abfd)->_bfd_coff_link_add_one_symbol)\
 (info, abfd, name, flags, section, value, string, cp, coll, hashp))
#define bfd_coff_link_output_has_begun(a,p) \
((coff_backend_info (a)->_bfd_coff_link_output_has_begun) (a,p))
#define bfd_coff_final_link_postscript(a,p) \
((coff_backend_info (a)->_bfd_coff_final_link_postscript) (a,p))

3.2.2.7 Writing relocations

To write relocations, the back end steps though the canonical relocation table and create an internal_reloc. The symbol index to use is removed from the offset field in the symbol table supplied. The address comes directly from the sum of the section base address and the relocation offset; the type is dug directly from the howto field. Then the internal_reloc is swapped into the shape of an external_reloc and written out to disk.

3.2.2.8 Reading linenumbers

Creating the linenumber table is done by reading in the entire coff linenumber table, and creating another table for internal use.

A coff linenumber table is structured so that each function is marked as having a line number of 0. Each line within the function is an offset from the first line in the function. The base of the line number information for the table is stored in the symbol associated with the function. Note: The PE format uses line number 0 for a flag indicating a new source file. The information is copied from the external to the internal table, and each symbol which marks a function is marked by pointing its... How does this work ?

3.2.2.9 Reading relocations

Coff relocations are easily transformed into the internal BFD form (arelent). Reading a coff relocation table is done in the following stages:

• Read the entire coff relocation table into memory.

• Process each relocation in turn; first swap it from the external to the internal form.

• Turn the symbol referenced in the relocation’s symbol index into a pointer into the canonical symbol table. This table is the same as the one returned by a call to bfd_ canonicalize_symtab. The back end will call that routine and save the result if a canonicalization hasn’t been done.

• The reloc index is turned into a pointer to a howto structure, in a back end specific way. For instance, the 386 and 960 use the r_type to directly produce an index into a howto table vector; the 88k subtracts a number from the r_type field and creates an addend field.

3.3 ELF backends

BFD support for ELF formats is being worked on. Currently, the best supported back ends are for sparc and i386 (running svr4 or Solaris 2). Documentation of the internals of the support code still needs to be written. The code is changing quickly enough that we haven’t bothered yet.

3.3.0.1 bfd_elf_find_section

Synopsis

     struct elf_internal_shdr *bfd_elf_find_section (bfd *abfd, char *name);

Description Helper functions for GDB to locate the string tables. Since BFD hides string tables from callers, GDB needs to use an internal hook to find them. Sun’s .stabstr, in particular, isn’t even pointed to by the .stab section, so ordinary mechanisms wouldn’t work to find it, even if we had some.

Index 108

Index

bfd final link relocate . . . . . . . . . . . . . . . . . . . . . . . 88 bfd generic link add archive symbols . . . . . . . . . 86 bfd generic link add one symbol. . . . . . . . . . . . . . 86 bfd link add symbols in target vector . . . . . . . . . 85 bfd link final link in target vector . . . . . . . . . . . . 87 bfd link hash table create in target vector. . . . . 85 bfd relocate contents . . . . . . . . . . . . . . . . . . . . . . . . 88 bfd strip section from output . . . . . . . . . . . . . . . . 27

A

aout size machine type . . . . . . . . . . . . . . . . . . . . . . . 95 aout size mkobject. . . . . . . . . . . . . . . . . . . . . . . . . . . 95 aout size new section hook . . . . . . . . . . . . . . . . . . . 95 aout size set arch mach . . . . . . . . . . . . . . . . . . . . . . 95 aout size some aout object p . . . . . . . . . . . . . . . . . 94 aout size swap exec header in . . . . . . . . . . . . . . . . 94 aout size swap exec header out . . . . . . . . . . . . . . . 94 arelent chain . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

B

BFD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 BFD canonical format. . . . . . . . . . . . . . . . . . . . . . . . . 3 bfd alloc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 bfd arch bits per address . . . . . . . . . . . . . . . . . . . . . 76 bfd arch bits per byte . . . . . . . . . . . . . . . . . . . . . . . . 76 bfd arch get compatible . . . . . . . . . . . . . . . . . . . . . . 75 bfd arch list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 bfd arch mach octets per byte . . . . . . . . . . . . . . . . 77 bfd cache close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 bfd cache init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 bfd cache lookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 bfd cache lookup worker. . . . . . . . . . . . . . . . . . . . . . 84 BFD CACHE MAX OPEN macro . . . . . . . . . . . . 83 bfd canonicalize reloc . . . . . . . . . . . . . . . . . . . . . . . . 10 bfd canonicalize symtab . . . . . . . . . . . . . . . . . . . . . . 32 bfd check format . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 bfd check format matches . . . . . . . . . . . . . . . . . . . . 36 bfd check overflow . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 bfd close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 bfd close all done . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 bfd coff backend data . . . . . . . . . . . . . . . . . . . . . . . . 99 bfd copy private bfd data . . . . . . . . . . . . . . . . . . . . 12 bfd copy private section data . . . . . . . . . . . . . . . . . 27 bfd copy private symbol data . . . . . . . . . . . . . . . . . 34 bfd core file failing command . . . . . . . . . . . . . . . . . 61 bfd core file failing signal. . . . . . . . . . . . . . . . . . . . . 62 bfd create . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 bfd decode symclass . . . . . . . . . . . . . . . . . . . . . . . . . 33 bfd default arch struct . . . . . . . . . . . . . . . . . . . . . . . 75 bfd default compatible . . . . . . . . . . . . . . . . . . . . . . . 76 bfd default reloc type lookup . . . . . . . . . . . . . . . . . 60 bfd default scan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 bfd default set arch mach . . . . . . . . . . . . . . . . . . . . 75 bfd elf find section . . . . . . . . . . . . . . . . . . . . . . . . . . 107 bfd errmsg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 bfd fdopenr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 bfd find target. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 bfd format string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 bfd generic gc sections . . . . . . . . . . . . . . . . . . . . . . . 61 bfd generic get relocated section contents . . . . . . 61 bfd generic relax section. . . . . . . . . . . . . . . . . . . . . . 61 bfd get arch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 bfd get arch info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 bfd get error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 bfd get error handler. . . . . . . . . . . . . . . . . . . . . . . . . 10 bfd get gp size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 bfd get mach . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 bfd get mtime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 bfd get next mapent . . . . . . . . . . . . . . . . . . . . . . . . . 35 bfd get reloc code name . . . . . . . . . . . . . . . . . . . . . . 60 bfd get reloc size . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 bfd get reloc upper bound . . . . . . . . . . . . . . . . . . . . 10 bfd get section by name . . . . . . . . . . . . . . . . . . . . . . 24 bfd get section contents . . . . . . . . . . . . . . . . . . . . . . 26 bfd get size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11, 80 bfd get symtab upper bound . . . . . . . . . . . . . . . . . 32 bfd h put size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 bfd hash allocate . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 bfd hash lookup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 bfd hash newfunc . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 bfd hash table free . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 bfd hash table init . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 bfd hash table init n . . . . . . . . . . . . . . . . . . . . . . . . . 89 bfd hash traverse . . . . . . . . . . . . . . . . . . . . . . . . . . . . 90 bfd init . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 bfd install relocation . . . . . . . . . . . . . . . . . . . . . . . . . 44 bfd is local label . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 bfd is local label name . . . . . . . . . . . . . . . . . . . . . . . 32 bfd is undefined symclass. . . . . . . . . . . . . . . . . . . . . 33 bfd last cache . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 bfd link split section . . . . . . . . . . . . . . . . . . . . . . . . . 89 bfd log2. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 bfd lookup arch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 bfd make debug symbol . . . . . . . . . . . . . . . . . . . . . . 33 bfd make empty symbol . . . . . . . . . . . . . . . . . . . . . . 33 bfd make readable . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 bfd make section . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 bfd make section anyway . . . . . . . . . . . . . . . . . . . . . 25 bfd make section old way. . . . . . . . . . . . . . . . . . . . . 25 bfd make writable . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 bfd map over sections . . . . . . . . . . . . . . . . . . . . . . . . 25 bfd merge private bfd data . . . . . . . . . . . . . . . . . . . 13 bfd octets per byte . . . . . . . . . . . . . . . . . . . . . . . . . . 77 bfd open file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 bfd openr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 bfd openr next archived file . . . . . . . . . . . . . . . . . . 35 bfd openstreamr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 bfd openw. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 bfd perform relocation . . . . . . . . . . . . . . . . . . . . . . . 44 bfd perror . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 bfd print symbol vandf . . . . . . . . . . . . . . . . . . . . . . . 33 bfd printable arch mach . . . . . . . . . . . . . . . . . . . . . . 77 bfd printable name . . . . . . . . . . . . . . . . . . . . . . . . . . 74 bfd put size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 BFD_RELOC_12_PCREL . . . . . . . . . . . . . . . . . . . . . . . . . 45 BFD_RELOC_14 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 BFD_RELOC_16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 BFD_RELOC_16_BASEREL. . . . . . . . . . . . . . . . . . . . . . . 46

BFD_RELOC_16_GOT_PCREL . . . . . . . . . . . . . . . . . . . . 46

BFD_RELOC_16_GOTOFF. . . . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_16_PCREL . . . . . . . . . . . . . . . . . . . . . . . . . 45 BFD_RELOC_16_PCREL_S2 . . . . . . . . . . . . . . . . . . . . . 46

BFD_RELOC_16_PLT_PCREL . . . . . . . . . . . . . . . . . . . . 46

BFD_RELOC_16_PLTOFF. . . . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_23_PCREL_S2 . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_24 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 BFD_RELOC_24_PCREL . . . . . . . . . . . . . . . . . . . . . . . . . 45

BFD_RELOC_24_PLT_PCREL . . . . . . . . . . . . . . . . . . . . 46

BFD_RELOC_26 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 BFD_RELOC_32 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 BFD_RELOC_32_BASEREL. . . . . . . . . . . . . . . . . . . . . . . 46

BFD_RELOC_32_GOT_PCREL . . . . . . . . . . . . . . . . . . . . 45

BFD_RELOC_32_GOTOFF. . . . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_32_PCREL . . . . . . . . . . . . . . . . . . . . . . . . . 45 BFD_RELOC_32_PCREL_S2 . . . . . . . . . . . . . . . . . . . . . 46

BFD_RELOC_32_PLT_PCREL . . . . . . . . . . . . . . . . . . . . 46

BFD_RELOC_32_PLTOFF. . . . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_386_COPY . . . . . . . . . . . . . . . . . . . . . . . . . 51

BFD_RELOC_386_GLOB_DAT . . . . . . . . . . . . . . . . . . . . 51

BFD_RELOC_386_GOT32. . . . . . . . . . . . . . . . . . . . . . . . 51 BFD_RELOC_386_GOTOFF. . . . . . . . . . . . . . . . . . . . . . . 51 BFD_RELOC_386_GOTPC. . . . . . . . . . . . . . . . . . . . . . . . 51

                                                BFD_RELOC_386_JUMP_SLOT . . . . . . . . . . . . . . . . . . . 51

BFD_RELOC_386_PLT32. . . . . . . . . . . . . . . . . . . . . . . . 51

                                                BFD_RELOC_386_RELATIVE . . . . . . . . . . . . . . . . . . . . 51

BFD_RELOC_64 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 BFD_RELOC_64_PCREL . . . . . . . . . . . . . . . . . . . . . . . . . 45

                                                BFD_RELOC_68K_GLOB_DAT . . . . . . . . . . . . . . . . . . . . 46
                                                BFD_RELOC_68K_JMP_SLOT . . . . . . . . . . . . . . . . . . . . 46
                                                BFD_RELOC_68K_RELATIVE . . . . . . . . . . . . . . . . . . . . 46

BFD_RELOC_8 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 BFD_RELOC_8_BASEREL. . . . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_8_FFnn . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_8_GOT_PCREL . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_8_GOTOFF . . . . . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_8_PCREL . . . . . . . . . . . . . . . . . . . . . . . . . . 45 BFD_RELOC_8_PLT_PCREL . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_8_PLTOFF . . . . . . . . . . . . . . . . . . . . . . . . . 46

                                                BFD_RELOC_ALPHA_CODEADDR . . . . . . . . . . . . . . . . . . 49
                                                BFD_RELOC_ALPHA_ELF_LITERAL . . . . . . . . . . . . . . . 49
                                                BFD_RELOC_ALPHA_GPDISP . . . . . . . . . . . . . . . . . . . . 48
                                                BFD_RELOC_ALPHA_GPDISP_HI16 . . . . . . . . . . . . . . . 48
                                                BFD_RELOC_ALPHA_GPDISP_LO16 . . . . . . . . . . . . . . . 48

BFD_RELOC_ALPHA_HINT. . . . . . . . . . . . . . . . . . . . . . . 49

                                                BFD_RELOC_ALPHA_LINKAGE . . . . . . . . . . . . . . . . . . . 49
                                                BFD_RELOC_ALPHA_LITERAL . . . . . . . . . . . . . . . . . . . 48
                                                BFD_RELOC_ALPHA_LITUSE . . . . . . . . . . . . . . . . . . . . 49
                                                BFD_RELOC_ALPHA_USER_GPDISP . . . . . . . . . . . . . . . 49
                                                BFD_RELOC_ALPHA_USER_GPRELHIGH . . . . . . . . . . . . 49
                                                BFD_RELOC_ALPHA_USER_GPRELLOW . . . . . . . . . . . . . 49
                                                BFD_RELOC_ALPHA_USER_LITERAL . . . . . . . . . . . . . . 49
                                                BFD_RELOC_ALPHA_USER_LITUSE_BASE. . . . . . . . . . 49
                                                BFD_RELOC_ALPHA_USER_LITUSE_BYTOFF . . . . . . . 49
                                                BFD_RELOC_ALPHA_USER_LITUSE_JSR . . . . . . . . . . . 49
                                                BFD_RELOC_ARC_B22_PCREL . . . . . . . . . . . . . . . . . . . 54

BFD_RELOC_ARC_B26 . . . . . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_ARM_ADR_IMM . . . . . . . . . . . . . . . . . . . . . 53

                                                BFD_RELOC_ARM_ADRL_IMMEDIATE . . . . . . . . . . . . . . 53

BFD_RELOC_ARM_COPY . . . . . . . . . . . . . . . . . . . . . . . . . 53

                                                BFD_RELOC_ARM_CP_OFF_IMM . . . . . . . . . . . . . . . . . . 53
                                                BFD_RELOC_ARM_GLOB_DAT . . . . . . . . . . . . . . . . . . . . 53

BFD_RELOC_ARM_GOT12. . . . . . . . . . . . . . . . . . . . . . . . 53 BFD_RELOC_ARM_GOT32. . . . . . . . . . . . . . . . . . . . . . . . 53 BFD_RELOC_ARM_GOTOFF. . . . . . . . . . . . . . . . . . . . . . . 53 BFD_RELOC_ARM_GOTPC. . . . . . . . . . . . . . . . . . . . . . . . 53

                                                BFD_RELOC_ARM_HWLITERAL . . . . . . . . . . . . . . . . . . . 53
                                                BFD_RELOC_ARM_IMMEDIATE . . . . . . . . . . . . . . . . . . . 53

BFD_RELOC_ARM_IN_POOL . . . . . . . . . . . . . . . . . . . . . 53

                                                BFD_RELOC_ARM_JUMP_SLOT . . . . . . . . . . . . . . . . . . . 53

BFD_RELOC_ARM_LDR_IMM . . . . . . . . . . . . . . . . . . . . . 53 BFD_RELOC_ARM_LITERAL . . . . . . . . . . . . . . . . . . . . . 53 BFD_RELOC_ARM_MULTI. . . . . . . . . . . . . . . . . . . . . . . . 53

BFD_RELOC_ARM_OFFSET_IMM . . . . . . . . . . . . . . . . . . 53
BFD_RELOC_ARM_OFFSET_IMM8 . . . . . . . . . . . . . . . . . 53
BFD_RELOC_ARM_PCREL_BLX . . . . . . . . . . . . . . . . . . . 52
BFD_RELOC_ARM_PCREL_BRANCH . . . . . . . . . . . . . . . . 52

BFD_RELOC_ARM_PLT32. . . . . . . . . . . . . . . . . . . . . . . . 53

BFD_RELOC_ARM_RELATIVE . . . . . . . . . . . . . . . . . . . . 53
BFD_RELOC_ARM_SHIFT_IMM . . . . . . . . . . . . . . . . . . . 53

BFD_RELOC_ARM_SWI . . . . . . . . . . . . . . . . . . . . . . . . . . 53

BFD_RELOC_ARM_THUMB_ADD . . . . . . . . . . . . . . . . . . . 53
BFD_RELOC_ARM_THUMB_IMM . . . . . . . . . . . . . . . . . . . 53
BFD_RELOC_ARM_THUMB_OFFSET . . . . . . . . . . . . . . . . 53
BFD_RELOC_ARM_THUMB_SHIFT . . . . . . . . . . . . . . . . . 53
BFD_RELOC_AVR_13_PCREL . . . . . . . . . . . . . . . . . . . . 59

BFD_RELOC_AVR_16_PM. . . . . . . . . . . . . . . . . . . . . . . . 59 BFD_RELOC_AVR_7_PCREL . . . . . . . . . . . . . . . . . . . . . 58 BFD_RELOC_AVR_CALL . . . . . . . . . . . . . . . . . . . . . . . . . 60 BFD_RELOC_AVR_HH8_LDI . . . . . . . . . . . . . . . . . . . . . 59

BFD_RELOC_AVR_HH8_LDI_NEG . . . . . . . . . . . . . . . . . 59
BFD_RELOC_AVR_HH8_LDI_PM . . . . . . . . . . . . . . . . . . 59
BFD_RELOC_AVR_HH8_LDI_PM_NEG . . . . . . . . . . . . . . 60

BFD_RELOC_AVR_HI8_LDI . . . . . . . . . . . . . . . . . . . . . 59

BFD_RELOC_AVR_HI8_LDI_NEG . . . . . . . . . . . . . . . . . 59
BFD_RELOC_AVR_HI8_LDI_PM . . . . . . . . . . . . . . . . . . 59
BFD_RELOC_AVR_HI8_LDI_PM_NEG . . . . . . . . . . . . . . 60

BFD_RELOC_AVR_LO8_LDI . . . . . . . . . . . . . . . . . . . . . 59

BFD_RELOC_AVR_LO8_LDI_NEG . . . . . . . . . . . . . . . . . 59
BFD_RELOC_AVR_LO8_LDI_PM . . . . . . . . . . . . . . . . . . 59
BFD_RELOC_AVR_LO8_LDI_PM_NEG . . . . . . . . . . . . . . 59

bfd reloc code type . . . . . . . . . . . . . . . . . . . . . . . . . . 44 BFD_RELOC_CTOR . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

BFD_RELOC_D10V_10_PCREL_L . . . . . . . . . . . . . . . . . 54
BFD_RELOC_D10V_10_PCREL_R . . . . . . . . . . . . . . . . . 54

BFD_RELOC_D10V_18 . . . . . . . . . . . . . . . . . . . . . . . . . . 54

BFD_RELOC_D10V_18_PCREL . . . . . . . . . . . . . . . . . . . 55

BFD_RELOC_D30V_15 . . . . . . . . . . . . . . . . . . . . . . . . . . 55

BFD_RELOC_D30V_15_PCREL . . . . . . . . . . . . . . . . . . . 55
BFD_RELOC_D30V_15_PCREL_R . . . . . . . . . . . . . . . . . 55

BFD_RELOC_D30V_21 . . . . . . . . . . . . . . . . . . . . . . . . . . 55

BFD_RELOC_D30V_21_PCREL . . . . . . . . . . . . . . . . . . . 55
BFD_RELOC_D30V_21_PCREL_R . . . . . . . . . . . . . . . . . 55

BFD_RELOC_D30V_32 . . . . . . . . . . . . . . . . . . . . . . . . . . 55

BFD_RELOC_D30V_32_PCREL . . . . . . . . . . . . . . . . . . . 55

BFD_RELOC_D30V_6 . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

BFD_RELOC_D30V_9_PCREL . . . . . . . . . . . . . . . . . . . . 55
BFD_RELOC_D30V_9_PCREL_R . . . . . . . . . . . . . . . . . . 55

                                                BFD_RELOC_FR30_10_IN_8 . . . . . . . . . . . . . . . . . . . . 58
                                                BFD_RELOC_FR30_12_PCREL . . . . . . . . . . . . . . . . . . . 58

BFD_RELOC_FR30_20 . . . . . . . . . . . . . . . . . . . . . . . . . . 58 BFD_RELOC_FR30_48 . . . . . . . . . . . . . . . . . . . . . . . . . . 58 BFD_RELOC_FR30_6_IN_4 . . . . . . . . . . . . . . . . . . . . . 58 BFD_RELOC_FR30_8_IN_8 . . . . . . . . . . . . . . . . . . . . . 58 BFD_RELOC_FR30_9_IN_8 . . . . . . . . . . . . . . . . . . . . . 58

                                                BFD_RELOC_FR30_9_PCREL . . . . . . . . . . . . . . . . . . . . 58

BFD_RELOC_GPREL16 . . . . . . . . . . . . . . . . . . . . . . . . . . 47 BFD_RELOC_GPREL32 . . . . . . . . . . . . . . . . . . . . . . . . . . 47 BFD_RELOC_HI16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

                                                BFD_RELOC_HI16_BASEREL . . . . . . . . . . . . . . . . . . . . 46

BFD_RELOC_HI16_GOTOFF . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_HI16_PLTOFF . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_HI16_S . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

                                                BFD_RELOC_HI16_S_BASEREL . . . . . . . . . . . . . . . . . . 46
                                                BFD_RELOC_HI16_S_GOTOFF . . . . . . . . . . . . . . . . . . . 46
                                                BFD_RELOC_HI16_S_PLTOFF . . . . . . . . . . . . . . . . . . . 46

BFD_RELOC_HI22 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 BFD_RELOC_I370_D12 . . . . . . . . . . . . . . . . . . . . . . . . . 52 BFD_RELOC_I960_CALLJ. . . . . . . . . . . . . . . . . . . . . . . 47 BFD_RELOC_LO10 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 BFD_RELOC_LO16 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

                                                BFD_RELOC_LO16_BASEREL . . . . . . . . . . . . . . . . . . . . 46

BFD_RELOC_LO16_GOTOFF . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_LO16_PLTOFF . . . . . . . . . . . . . . . . . . . . . 46

                                                BFD_RELOC_M32R_10_PCREL . . . . . . . . . . . . . . . . . . . 55
                                                BFD_RELOC_M32R_18_PCREL . . . . . . . . . . . . . . . . . . . 56

BFD_RELOC_M32R_24 . . . . . . . . . . . . . . . . . . . . . . . . . . 55

                                                BFD_RELOC_M32R_26_PCREL . . . . . . . . . . . . . . . . . . . 56
                                                BFD_RELOC_M32R_HI16_SLO . . . . . . . . . . . . . . . . . . . 56
                                                BFD_RELOC_M32R_HI16_ULO . . . . . . . . . . . . . . . . . . . 56

BFD_RELOC_M32R_LO16. . . . . . . . . . . . . . . . . . . . . . . . 56 BFD_RELOC_M32R_SDA16. . . . . . . . . . . . . . . . . . . . . . . 56

                                                BFD_RELOC_MCORE_PCREL_32 . . . . . . . . . . . . . . . . . . 58
                                                BFD_RELOC_MCORE_PCREL_IMM11BY2 . . . . . . . . . . . . 58
                                                BFD_RELOC_MCORE_PCREL_IMM4BY2 . . . . . . . . . . . . . 58
                                                BFD_RELOC_MCORE_PCREL_IMM8BY4 . . . . . . . . . . . . . 58
                                                BFD_RELOC_MCORE_PCREL_JSR_IMM11BY2 . . . . . . . 58

BFD_RELOC_MCORE_RVA. . . . . . . . . . . . . . . . . . . . . . . . 58

                                                BFD_RELOC_MIPS_CALL_HI16 . . . . . . . . . . . . . . . . . . 50
                                                BFD_RELOC_MIPS_CALL_LO16 . . . . . . . . . . . . . . . . . . 50

BFD_RELOC_MIPS_CALL16 . . . . . . . . . . . . . . . . . . . . . 50

                                                BFD_RELOC_MIPS_GOT_DISP . . . . . . . . . . . . . . . . . . . 50
                                                BFD_RELOC_MIPS_GOT_HI16 . . . . . . . . . . . . . . . . . . . 50
                                                BFD_RELOC_MIPS_GOT_LO16 . . . . . . . . . . . . . . . . . . . 50
                                                BFD_RELOC_MIPS_GOT_OFST . . . . . . . . . . . . . . . . . . . 50
                                                BFD_RELOC_MIPS_GOT_PAGE . . . . . . . . . . . . . . . . . . . 50

BFD_RELOC_MIPS_GOT16. . . . . . . . . . . . . . . . . . . . . . . 50 BFD_RELOC_MIPS_GPREL. . . . . . . . . . . . . . . . . . . . . . . 50

BFD_RELOC_MIPS_GPREL32 . . . . . . . . . . . . . . . . . . . . 50

BFD_RELOC_MIPS_JMP . . . . . . . . . . . . . . . . . . . . . . . . . 49

BFD_RELOC_MIPS_LITERAL . . . . . . . . . . . . . . . . . . . . 50

BFD_RELOC_MIPS_SUB . . . . . . . . . . . . . . . . . . . . . . . . . 50

BFD_RELOC_MIPS16_GPREL . . . . . . . . . . . . . . . . . . . . 50

BFD_RELOC_MIPS16_JMP. . . . . . . . . . . . . . . . . . . . . . . 50

BFD_RELOC_MN10300_16_PCREL . . . . . . . . . . . . . . . . 57
BFD_RELOC_MN10300_32_PCREL . . . . . . . . . . . . . . . . 57

BFD_RELOC_NONE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

BFD_RELOC_NS32K_DISP_16 . . . . . . . . . . . . . . . . . . . 51
BFD_RELOC_NS32K_DISP_16_PCREL . . . . . . . . . . . . . 51
BFD_RELOC_NS32K_DISP_32 . . . . . . . . . . . . . . . . . . . 51
BFD_RELOC_NS32K_DISP_32_PCREL . . . . . . . . . . . . . 51
BFD_RELOC_NS32K_DISP_8 . . . . . . . . . . . . . . . . . . . . 51
BFD_RELOC_NS32K_DISP_8_PCREL . . . . . . . . . . . . . . 51
BFD_RELOC_NS32K_IMM_16 . . . . . . . . . . . . . . . . . . . . 51
BFD_RELOC_NS32K_IMM_16_PCREL . . . . . . . . . . . . . . 51
BFD_RELOC_NS32K_IMM_32 . . . . . . . . . . . . . . . . . . . . 51
BFD_RELOC_NS32K_IMM_32_PCREL . . . . . . . . . . . . . . 51

BFD_RELOC_NS32K_IMM_8 . . . . . . . . . . . . . . . . . . . . . 51

BFD_RELOC_NS32K_IMM_8_PCREL . . . . . . . . . . . . . . . 51
BFD_RELOC_PCREL_HI16_S . . . . . . . . . . . . . . . . . . . . 50

BFD_RELOC_PCREL_LO16. . . . . . . . . . . . . . . . . . . . . . . 50

BFD_RELOC_PJ_CODE_DIR16 . . . . . . . . . . . . . . . . . . . 51
BFD_RELOC_PJ_CODE_DIR32 . . . . . . . . . . . . . . . . . . . 51
BFD_RELOC_PJ_CODE_HI16 . . . . . . . . . . . . . . . . . . . . 51
BFD_RELOC_PJ_CODE_LO16 . . . . . . . . . . . . . . . . . . . . 51
BFD_RELOC_PJ_CODE_REL16 . . . . . . . . . . . . . . . . . . . 51
BFD_RELOC_PJ_CODE_REL32 . . . . . . . . . . . . . . . . . . . 51

BFD_RELOC_PPC_B16 . . . . . . . . . . . . . . . . . . . . . . . . . . 52

BFD_RELOC_PPC_B16_BRNTAKEN . . . . . . . . . . . . . . . . 52
BFD_RELOC_PPC_B16_BRTAKEN . . . . . . . . . . . . . . . . . 52

BFD_RELOC_PPC_B26 . . . . . . . . . . . . . . . . . . . . . . . . . . 51 BFD_RELOC_PPC_BA16 . . . . . . . . . . . . . . . . . . . . . . . . . 52

BFD_RELOC_PPC_BA16_BRNTAKEN . . . . . . . . . . . . . . . 52
BFD_RELOC_PPC_BA16_BRTAKEN . . . . . . . . . . . . . . . . 52

BFD_RELOC_PPC_BA26 . . . . . . . . . . . . . . . . . . . . . . . . . 52 BFD_RELOC_PPC_COPY . . . . . . . . . . . . . . . . . . . . . . . . . 52

BFD_RELOC_PPC_EMB_BIT_FLD . . . . . . . . . . . . . . . . . 52
BFD_RELOC_PPC_EMB_MRKREF . . . . . . . . . . . . . . . . . . 52
BFD_RELOC_PPC_EMB_NADDR16 . . . . . . . . . . . . . . . . . 52
BFD_RELOC_PPC_EMB_NADDR16_HA . . . . . . . . . . . . . . 52
BFD_RELOC_PPC_EMB_NADDR16_HI . . . . . . . . . . . . . . 52
BFD_RELOC_PPC_EMB_NADDR16_LO . . . . . . . . . . . . . . 52
BFD_RELOC_PPC_EMB_NADDR32 . . . . . . . . . . . . . . . . . 52
BFD_RELOC_PPC_EMB_RELSDA . . . . . . . . . . . . . . . . . . 52

                                                BFD_RELOC_PPC_EMB_RELSEC16 . . . . . . . . . . . . . . . . 52
                                                BFD_RELOC_PPC_EMB_RELST_HA . . . . . . . . . . . . . . . . 52
                                                BFD_RELOC_PPC_EMB_RELST_HI . . . . . . . . . . . . . . . . 52
                                                BFD_RELOC_PPC_EMB_RELST_LO . . . . . . . . . . . . . . . . 52
                                                BFD_RELOC_PPC_EMB_SDA21 . . . . . . . . . . . . . . . . . . . 52
                                                BFD_RELOC_PPC_EMB_SDA2I16 . . . . . . . . . . . . . . . . . 52
                                                BFD_RELOC_PPC_EMB_SDA2REL . . . . . . . . . . . . . . . . . 52
                                                BFD_RELOC_PPC_EMB_SDAI16 . . . . . . . . . . . . . . . . . . 52
                                                BFD_RELOC_PPC_GLOB_DAT . . . . . . . . . . . . . . . . . . . . 52
                                                BFD_RELOC_PPC_JMP_SLOT . . . . . . . . . . . . . . . . . . . . 52
                                                BFD_RELOC_PPC_LOCAL24PC . . . . . . . . . . . . . . . . . . . 52
                                                BFD_RELOC_PPC_RELATIVE . . . . . . . . . . . . . . . . . . . . 52

BFD_RELOC_PPC_TOC16. . . . . . . . . . . . . . . . . . . . . . . . 52 BFD_RELOC_RVA . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 BFD_RELOC_SH_ALIGN . . . . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_CODE . . . . . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_COUNT . . . . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_DATA . . . . . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_IMM4 . . . . . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_IMM4BY2. . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_IMM4BY4. . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_IMM8 . . . . . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_IMM8BY2. . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_IMM8BY4. . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_LABEL . . . . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_LOOP_END . . . . . . . . . . . . . . . . . . . . . 54

                                                BFD_RELOC_SH_LOOP_START . . . . . . . . . . . . . . . . . . . 54
                                                BFD_RELOC_SH_PCDISP12BY2 . . . . . . . . . . . . . . . . . . 54
                                                BFD_RELOC_SH_PCDISP8BY2 . . . . . . . . . . . . . . . . . . . 53
                                                BFD_RELOC_SH_PCRELIMM8BY2 . . . . . . . . . . . . . . . . . 54
                                                BFD_RELOC_SH_PCRELIMM8BY4 . . . . . . . . . . . . . . . . . 54

BFD_RELOC_SH_SWITCH16 . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_SWITCH32 . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SH_USES . . . . . . . . . . . . . . . . . . . . . . . . . . 54 BFD_RELOC_SPARC_10 . . . . . . . . . . . . . . . . . . . . . . . . . 48 BFD_RELOC_SPARC_11 . . . . . . . . . . . . . . . . . . . . . . . . . 48 BFD_RELOC_SPARC_5 . . . . . . . . . . . . . . . . . . . . . . . . . . 48 BFD_RELOC_SPARC_6 . . . . . . . . . . . . . . . . . . . . . . . . . . 48 BFD_RELOC_SPARC_64 . . . . . . . . . . . . . . . . . . . . . . . . . 47 BFD_RELOC_SPARC_7 . . . . . . . . . . . . . . . . . . . . . . . . . . 48

                                                BFD_RELOC_SPARC_BASE13 . . . . . . . . . . . . . . . . . . . . 47
                                                BFD_RELOC_SPARC_BASE22 . . . . . . . . . . . . . . . . . . . . 47

BFD_RELOC_SPARC_COPY. . . . . . . . . . . . . . . . . . . . . . . 47

                                                BFD_RELOC_SPARC_DISP64 . . . . . . . . . . . . . . . . . . . . 48
                                                BFD_RELOC_SPARC_GLOB_DAT . . . . . . . . . . . . . . . . . . 47

BFD_RELOC_SPARC_GOT10 . . . . . . . . . . . . . . . . . . . . . 47 BFD_RELOC_SPARC_GOT13 . . . . . . . . . . . . . . . . . . . . . 47 BFD_RELOC_SPARC_GOT22 . . . . . . . . . . . . . . . . . . . . . 47 BFD_RELOC_SPARC_H44. . . . . . . . . . . . . . . . . . . . . . . . 48 BFD_RELOC_SPARC_HH22. . . . . . . . . . . . . . . . . . . . . . . 48 BFD_RELOC_SPARC_HIX22 . . . . . . . . . . . . . . . . . . . . . 48 BFD_RELOC_SPARC_HM10. . . . . . . . . . . . . . . . . . . . . . . 48

BFD_RELOC_SPARC_JMP_SLOT . . . . . . . . . . . . . . . . . . 47

BFD_RELOC_SPARC_L44. . . . . . . . . . . . . . . . . . . . . . . . 48 BFD_RELOC_SPARC_LM22. . . . . . . . . . . . . . . . . . . . . . . 48 BFD_RELOC_SPARC_LOX10 . . . . . . . . . . . . . . . . . . . . . 48 BFD_RELOC_SPARC_M44. . . . . . . . . . . . . . . . . . . . . . . . 48 BFD_RELOC_SPARC_OLO10 . . . . . . . . . . . . . . . . . . . . . 48

BFD_RELOC_SPARC_PC_HH22 . . . . . . . . . . . . . . . . . . . 48
BFD_RELOC_SPARC_PC_HM10 . . . . . . . . . . . . . . . . . . . 48
BFD_RELOC_SPARC_PC_LM22 . . . . . . . . . . . . . . . . . . . 48

BFD_RELOC_SPARC_PC10. . . . . . . . . . . . . . . . . . . . . . . 47 BFD_RELOC_SPARC_PC22. . . . . . . . . . . . . . . . . . . . . . . 47 BFD_RELOC_SPARC_PLT64 . . . . . . . . . . . . . . . . . . . . . 48

BFD_RELOC_SPARC_REGISTER . . . . . . . . . . . . . . . . . . 48
BFD_RELOC_SPARC_RELATIVE . . . . . . . . . . . . . . . . . . 47

BFD_RELOC_SPARC_REV32 . . . . . . . . . . . . . . . . . . . . . 48 BFD_RELOC_SPARC_UA32. . . . . . . . . . . . . . . . . . . . . . . 47

BFD_RELOC_SPARC_WDISP16 . . . . . . . . . . . . . . . . . . . 48
BFD_RELOC_SPARC_WDISP19 . . . . . . . . . . . . . . . . . . . 48
BFD_RELOC_SPARC_WDISP22 . . . . . . . . . . . . . . . . . . . 47
BFD_RELOC_SPARC_WPLT30 . . . . . . . . . . . . . . . . . . . . 47

BFD_RELOC_SPARC13 . . . . . . . . . . . . . . . . . . . . . . . . . . 47 BFD_RELOC_SPARC22 . . . . . . . . . . . . . . . . . . . . . . . . . . 47

BFD_RELOC_THUMB_PCREL_BLX . . . . . . . . . . . . . . . . . 53
BFD_RELOC_THUMB_PCREL_BRANCH12 . . . . . . . . . . . . 54
BFD_RELOC_THUMB_PCREL_BRANCH23 . . . . . . . . . . . . 54
BFD_RELOC_THUMB_PCREL_BRANCH9 . . . . . . . . . . . . . 54

BFD_RELOC_TIC30_LDP. . . . . . . . . . . . . . . . . . . . . . . . 57

BFD_RELOC_TIC54X_16_OF_23 . . . . . . . . . . . . . . . . . 58

BFD_RELOC_TIC54X_23. . . . . . . . . . . . . . . . . . . . . . . . 57

BFD_RELOC_TIC54X_MS7_OF_23 . . . . . . . . . . . . . . . . 58
BFD_RELOC_TIC54X_PARTLS7 . . . . . . . . . . . . . . . . . . 57
BFD_RELOC_TIC54X_PARTMS9 . . . . . . . . . . . . . . . . . . 57

bfd reloc type lookup . . . . . . . . . . . . . . . . . . . . . . . . 60

BFD_RELOC_V850_22_PCREL . . . . . . . . . . . . . . . . . . . 56
BFD_RELOC_V850_9_PCREL . . . . . . . . . . . . . . . . . . . . 56
BFD_RELOC_V850_CALLT_16_16_OFFSET. . . . . . . . . 57
BFD_RELOC_V850_CALLT_6_7_OFFSET . . . . . . . . . . . 57
BFD_RELOC_V850_SDA_15_16_OFFSET . . . . . . . . . . . 56
BFD_RELOC_V850_SDA_16_16_OFFSET . . . . . . . . . . . 56
BFD_RELOC_V850_SDA_16_16_SPLIT_OFFSET . . . . 57
BFD_RELOC_V850_TDA_16_16_OFFSET . . . . . . . . . . . 57
BFD_RELOC_V850_TDA_4_4_OFFSET . . . . . . . . . . . . . 57
BFD_RELOC_V850_TDA_4_5_OFFSET . . . . . . . . . . . . . 57
BFD_RELOC_V850_TDA_6_8_OFFSET . . . . . . . . . . . . . 56

                                                BFD_RELOC_V850_TDA_7_7_OFFSET . . . . . . . . . . . . . 57
                                                BFD_RELOC_V850_TDA_7_8_OFFSET . . . . . . . . . . . . . 56
                                                BFD_RELOC_V850_ZDA_15_16_OFFSET . . . . . . . . . . . 56
                                                BFD_RELOC_V850_ZDA_16_16_OFFSET . . . . . . . . . . . 56
                                                BFD_RELOC_V850_ZDA_16_16_SPLIT_OFFSET . . . . 57
                                                BFD_RELOC_VTABLE_ENTRY . . . . . . . . . . . . . . . . . . . . 60
                                                BFD_RELOC_VTABLE_INHERIT . . . . . . . . . . . . . . . . . . 60

bfd scan arch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 bfd scan vma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 bfd seach for target . . . . . . . . . . . . . . . . . . . . . . . . . . 70 bfd set arch info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 bfd set archive head . . . . . . . . . . . . . . . . . . . . . . . . . 35 bfd set default target . . . . . . . . . . . . . . . . . . . . . . . . 70 bfd set error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 bfd set error handler . . . . . . . . . . . . . . . . . . . . . . . . . . 9 bfd set error program name . . . . . . . . . . . . . . . . . . 10 bfd set file flags . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 bfd set format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 bfd set gp size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 bfd set private flags . . . . . . . . . . . . . . . . . . . . . . . . . . 13 bfd set reloc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 bfd set section contents . . . . . . . . . . . . . . . . . . . . . . 26 bfd set section flags . . . . . . . . . . . . . . . . . . . . . . . . . . 25 bfd set section size. . . . . . . . . . . . . . . . . . . . . . . . . . . 26 bfd set start address . . . . . . . . . . . . . . . . . . . . . . . . . 11 bfd set symtab . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 bfd symbol info. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 bfd target list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 bfd write bigendian 4byte int . . . . . . . . . . . . . . . . . 80

C

coff symbol type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 core file matches executable p . . . . . . . . . . . . . . . . 62

H

Hash tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89

I

internal object-file format . . . . . . . . . . . . . . . . . . . . . 3

L

Linker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84

S

stuff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

T

target vector ( bfd final link) . . . . . . . . . . . . . . . . . 87 target vector ( bfd link add symbols). . . . . . . . . . 85 target vector ( bfd link hash table create) . . . . . 85 The HOWTO Macro . . . . . . . . . . . . . . . . . . . . . . . . . 42

W

what is it? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

The body of this manual is set in

cmr10 at 10.95pt, with headings in cmb10 at 10.95pt

and examples in cmtt10 at 10.95pt.

cmti10 at 10.95pt and

cmsl10 at 10.95pt are used for emphasis.

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1.1 History . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 How To Use BFD . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 What BFD Version 2 Can Do . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

1.3.1 Information Loss . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3.2 The BFD canonical object-file format . . . . . . . . . . . . 3

2 BFD front end. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

2.1 typedef bfd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.2 Error reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

2.2.1 Type bfd_error_type . . . . . . . . . . . . . . . . . . . . . . . . . . 8 2.2.1.1 bfd_get_error . . . . . . . . . . . . . . . . . . . . . . . . 9 2.2.1.2 bfd_set_error . . . . . . . . . . . . . . . . . . . . . . . . 9 2.2.1.3 bfd_errmsg . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.2.1.4 bfd_perror . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2.2.2 BFD error handler . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.2.2.1 bfd_set_error_handler . . . . . . . . . . . . . . 10

2.2.2.2 bfd_set_error_program_name. . . . . . . . . 10

2.2.2.3 bfd_get_error_handler . . . . . . . . . . . . . . 10

2.3 Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

2.3.0.1 bfd_get_reloc_upper_bound . . . . . . . . . . 10
2.3.0.2 bfd_canonicalize_reloc . . . . . . . . . . . . . 10

2.3.0.3 bfd_set_reloc . . . . . . . . . . . . . . . . . . . . . . . 11 2.3.0.4 bfd_set_file_flags. . . . . . . . . . . . . . . . . . 11 2.3.0.5 bfd_set_start_address . . . . . . . . . . . . . . 11 2.3.0.6 bfd_get_mtime . . . . . . . . . . . . . . . . . . . . . . . 11 2.3.0.7 bfd_get_size . . . . . . . . . . . . . . . . . . . . . . . . 11 2.3.0.8 bfd_get_gp_size . . . . . . . . . . . . . . . . . . . . . 12 2.3.0.9 bfd_set_gp_size . . . . . . . . . . . . . . . . . . . . . 12 2.3.0.10 bfd_scan_vma . . . . . . . . . . . . . . . . . . . . . . . 12

2.3.0.11 bfd_copy_private_bfd_data. . . . . . . . . 12
2.3.0.12 bfd_merge_private_bfd_data. . . . . . . . 13

2.3.0.13 bfd_set_private_flags . . . . . . . . . . . . . 13 2.3.0.14 stuff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

2.4 Memory usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.5 Initialization . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2.5.0.1 bfd_init. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

2.6 Sections . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.6.1 Section input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 2.6.2 Section output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.6.3 Link orders . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 2.6.4 typedef asection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.6.5 Section prototypes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

2.6.5.1 bfd_get_section_by_name . . . . . . . . . . . . 24
2.6.5.2 bfd_make_section_old_way . . . . . . . . . . . 25
2.6.5.3 bfd_make_section_anyway . . . . . . . . . . . . 25

2.6.5.4 bfd_make_section . . . . . . . . . . . . . . . . . . . . 25 2.6.5.5 bfd_set_section_flags . . . . . . . . . . . . . . 25 2.6.5.6 bfd_map_over_sections . . . . . . . . . . . . . . 26 2.6.5.7 bfd_set_section_size . . . . . . . . . . . . . . . 26

2.6.5.8 bfd_set_section_contents . . . . . . . . . . . 26
2.6.5.9 bfd_get_section_contents . . . . . . . . . . . 27
2.6.5.10 bfd_copy_private_section_data . . . . 27
2.6.5.11 _bfd_strip_section_from_output . . . 27

2.7 Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 2.7.1 Reading symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 2.7.2 Writing symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 2.7.3 Mini Symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 2.7.4 typedef asymbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 2.7.5 Symbol handling functions . . . . . . . . . . . . . . . . . . . . . 32

2.7.5.1 bfd_get_symtab_upper_bound. . . . . . . . . 32

2.7.5.2 bfd_is_local_label. . . . . . . . . . . . . . . . . . 32

2.7.5.3 bfd_is_local_label_name . . . . . . . . . . . . 32
2.7.5.4 bfd_canonicalize_symtab . . . . . . . . . . . . 33

2.7.5.5 bfd_set_symtab . . . . . . . . . . . . . . . . . . . . . . 33

2.7.5.6 bfd_print_symbol_vandf . . . . . . . . . . . . . 33

2.7.5.7 bfd_make_empty_symbol . . . . . . . . . . . . . . 33 2.7.5.8 bfd_make_debug_symbol . . . . . . . . . . . . . . 33 2.7.5.9 bfd_decode_symclass . . . . . . . . . . . . . . . . 33

2.7.5.10 bfd_is_undefined_symclass . . . . . . . . 34

2.7.5.11 bfd_symbol_info . . . . . . . . . . . . . . . . . . . . 34

2.7.5.12 bfd_copy_private_symbol_data . . . . . 34

2.8 Archives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 2.8.0.1 bfd_get_next_mapent . . . . . . . . . . . . . . . . 35 2.8.0.2 bfd_set_archive_head . . . . . . . . . . . . . . . 35

2.8.0.3 bfd_openr_next_archived_file . . . . . . 35

2.9 File formats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 2.9.0.1 bfd_check_format . . . . . . . . . . . . . . . . . . . . 36

2.9.0.2 bfd_check_format_matches . . . . . . . . . . . 36

2.9.0.3 bfd_set_format . . . . . . . . . . . . . . . . . . . . . . 36 2.9.0.4 bfd_format_string. . . . . . . . . . . . . . . . . . . 37

2.10 Relocations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

2.10.1 typedef arelent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 2.10.1.1 enum complain_overflow . . . . . . . . . . . . 40 2.10.1.2 reloc_howto_type. . . . . . . . . . . . . . . . . . . 40 2.10.1.3 The HOWTO Macro. . . . . . . . . . . . . . . . . . . . . 42 2.10.1.4 bfd_get_reloc_size . . . . . . . . . . . . . . . . 43 2.10.1.5 arelent_chain . . . . . . . . . . . . . . . . . . . . . . 43 2.10.1.6 bfd_check_overflow . . . . . . . . . . . . . . . . 43

2.10.1.7 bfd_perform_relocation . . . . . . . . . . . . 44
2.10.1.8 bfd_install_relocation . . . . . . . . . . . . 44

2.11 The howto manager . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 2.11.0.1 bfd_reloc_code_type . . . . . . . . . . . . . . . 45 2.11.0.2 bfd_reloc_type_lookup . . . . . . . . . . . . . 60

2.11.0.3 bfd_default_reloc_type_lookup . . . . 60
2.11.0.4 bfd_get_reloc_code_name . . . . . . . . . . . 61
2.11.0.5 bfd_generic_relax_section. . . . . . . . . 61
2.11.0.6 bfd_generic_gc_sections . . . . . . . . . . . 61
2.11.0.7 bfd_generic_get_relocated_section_

contents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

2.12 Core files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

2.12.0.1 bfd_core_file_failing_command . . . . 62
2.12.0.2 bfd_core_file_failing_signal . . . . . 62
2.12.0.3 core_file_matches_executable_p . . . 62

2.13 Targets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62

2.13.1 bfd target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

2.13.1.1 bfd_set_default_target . . . . . . . . . . . . 70

2.13.1.2 bfd_find_target . . . . . . . . . . . . . . . . . . . . 70 2.13.1.3 bfd_target_list . . . . . . . . . . . . . . . . . . . . 70 2.13.1.4 bfd_seach_for_target . . . . . . . . . . . . . . 70

2.14 Architectures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 2.14.1 bfd architecture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 2.14.2 bfd arch info . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

2.14.2.1 bfd_printable_name . . . . . . . . . . . . . . . . 74 2.14.2.2 bfd_scan_arch . . . . . . . . . . . . . . . . . . . . . . 75 2.14.2.3 bfd_arch_list . . . . . . . . . . . . . . . . . . . . . . 75

2.14.2.4 bfd_arch_get_compatible . . . . . . . . . . . 75
2.14.2.5 bfd_default_arch_struct . . . . . . . . . . . 75

2.14.2.6 bfd_set_arch_info. . . . . . . . . . . . . . . . . . 75

2.14.2.7 bfd_default_set_arch_mach. . . . . . . . . 75

2.14.2.8 bfd_get_arch . . . . . . . . . . . . . . . . . . . . . . . 76 2.14.2.9 bfd_get_mach . . . . . . . . . . . . . . . . . . . . . . . 76

2.14.2.10 bfd_arch_bits_per_byte . . . . . . . . . . . 76
2.14.2.11 bfd_arch_bits_per_address. . . . . . . . 76
2.14.2.12 bfd_default_compatible . . . . . . . . . . . 76

2.14.2.13 bfd_default_scan. . . . . . . . . . . . . . . . . . 76 2.14.2.14 bfd_get_arch_info . . . . . . . . . . . . . . . . 77 2.14.2.15 bfd_lookup_arch . . . . . . . . . . . . . . . . . . . 77

2.14.2.16 bfd_printable_arch_mach . . . . . . . . . . 77

2.14.2.17 bfd_octets_per_byte . . . . . . . . . . . . . . 77

2.14.2.18 bfd_arch_mach_octets_per_byte . . . 77

2.15 Opening and closing BFDs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 2.15.0.1 bfd_openr . . . . . . . . . . . . . . . . . . . . . . . . . . 78 2.15.0.2 bfd_fdopenr . . . . . . . . . . . . . . . . . . . . . . . . 78 2.15.0.3 bfd_openstreamr . . . . . . . . . . . . . . . . . . . . 78 2.15.0.4 bfd_openw . . . . . . . . . . . . . . . . . . . . . . . . . . 78 2.15.0.5 bfd_close . . . . . . . . . . . . . . . . . . . . . . . . . . 79 2.15.0.6 bfd_close_all_done . . . . . . . . . . . . . . . . 79 2.15.0.7 bfd_create . . . . . . . . . . . . . . . . . . . . . . . . . 79 2.15.0.8 bfd_make_writable. . . . . . . . . . . . . . . . . . 79 2.15.0.9 bfd_make_readable. . . . . . . . . . . . . . . . . . 80 2.15.0.10 bfd_alloc . . . . . . . . . . . . . . . . . . . . . . . . . 80

2.16 Internal functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80

2.16.0.1 bfd_write_bigendian_4byte_int . . . . 80

2.16.0.2 bfd_put_size . . . . . . . . . . . . . . . . . . . . . . . 80 2.16.0.3 bfd_get_size . . . . . . . . . . . . . . . . . . . . . . . 80 2.16.0.4 bfd_h_put_size . . . . . . . . . . . . . . . . . . . . . 82 2.16.0.5 bfd_log2. . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

2.17 File caching. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83

2.17.0.1 BFD_CACHE_MAX_OPEN macro . . . . . . . . . . 83

2.17.0.2 bfd_last_cache . . . . . . . . . . . . . . . . . . . . . 83 2.17.0.3 bfd_cache_lookup. . . . . . . . . . . . . . . . . . . 83 2.17.0.4 bfd_cache_init . . . . . . . . . . . . . . . . . . . . . 83 2.17.0.5 bfd_cache_close . . . . . . . . . . . . . . . . . . . . 84 2.17.0.6 bfd_open_file . . . . . . . . . . . . . . . . . . . . . . 84

2.17.0.7 bfd_cache_lookup_worker . . . . . . . . . . . 84

2.18 Linker Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 2.18.1 Creating a linker hash table . . . . . . . . . . . . . . . . . . . 85 2.18.2 Adding symbols to the hash table . . . . . . . . . . . . . . 85

2.18.2.1 Differing file formats . . . . . . . . . . . . . . . . . 85 2.18.2.2 Adding symbols from an object file …. 86 2.18.2.3 Adding symbols from an archive . . . . . . 86

2.18.3 Performing the final link . . . . . . . . . . . . . . . . . . . . . . 87 2.18.3.1 Information provided by the linker …. 87 2.18.3.2 Relocating the section contents. . . . . . . . 88 2.18.3.3 Writing the symbol table . . . . . . . . . . . . . 88

2.18.3.4 bfd_link_split_section . . . . . . . . . . . . 89

2.19 Hash Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 89 2.19.1 Creating and freeing a hash table . . . . . . . . . . . . . . 89 2.19.2 Looking up or entering a string . . . . . . . . . . . . . . . . 90 2.19.3 Traversing a hash table . . . . . . . . . . . . . . . . . . . . . . . 90 2.19.4 Deriving a new hash table type . . . . . . . . . . . . . . . . 90

2.19.4.1 Define the derived structures. . . . . . . . . . 91 2.19.4.2 Write the derived creation routine . . . . . 91 2.19.4.3 Write other derived routines . . . . . . . . . . 92

3 BFD back ends . . . . . . . . . . . . . . . . . . . . . . . . . . . 93

3.1 a.out backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 3.1.1 Relocations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 94 3.1.2 Internal entry points . . . . . . . . . . . . . . . . . . . . . . . . . . . 94

3.1.2.1 aout_size_swap_exec_header_in . . . . . . 94
3.1.2.2 aout_size_swap_exec_header_out . . . . . 94
3.1.2.3 aout_size_some_aout_object_p . . . . . . . 95

3.1.2.4 aout_size_mkobject. . . . . . . . . . . . . . . . . . . 95 3.1.2.5 aout_size_machine_type . . . . . . . . . . . . . . 95 3.1.2.6 aout_size_set_arch_mach . . . . . . . . . . . . . 95 3.1.2.7 aout_size_new_section_hook. . . . . . . . . . 95

3.2 coff backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 3.2.1 Porting to a new version of coff . . . . . . . . . . . . . . . . . 96 3.2.2 How the coff backend works . . . . . . . . . . . . . . . . . . . . 96

3.2.2.1 File layout. . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 3.2.2.2 Bit twiddling . . . . . . . . . . . . . . . . . . . . . . . . . 97 3.2.2.3 Symbol reading . . . . . . . . . . . . . . . . . . . . . . . 97 3.2.2.4 Symbol writing . . . . . . . . . . . . . . . . . . . . . . . 98 3.2.2.5 coff_symbol_type . . . . . . . . . . . . . . . . . . . . 98 3.2.2.6 bfd_coff_backend_data . . . . . . . . . . . . . 100 3.2.2.7 Writing relocations . . . . . . . . . . . . . . . . . . . 106 3.2.2.8 Reading linenumbers . . . . . . . . . . . . . . . . . 106 3.2.2.9 Reading relocations . . . . . . . . . . . . . . . . . . 107

3.3 ELF backends . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 3.3.0.1 bfd_elf_find_section . . . . . . . . . . . . . . 107

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108